### v1

# 00footnotetext: Shared first author; Shared last author

Abstract
We introduce DataComp for Language Models (DCLM), a testbed for controlled dataset experiments with the goal of improving language models. As part of DCLM, we provide a standardized corpus of 240T tokens extracted from Common Crawl, effective pretraining recipes based on the OpenLM framework, and a broad suite of 53 downstream evaluations. Participants in the DCLM benchmark can experiment with data curation strategies such as deduplication, filtering, and data mixing at model scales ranging from 412M to 7B parameters. As a baseline for DCLM, we conduct extensive experiments and find that model-based filtering is key to assembling a high-quality training set. The resulting dataset, DCLM-baseline, enables training a 7B parameter language model from scratch to 6464% 5-shot accuracy on MMLU with 2.6T training tokens. Compared to MAP-Neo, the previous state-of-the-art in open-data language models, DCLM-baseline represents a 6.6 percentage point improvement on MMLU while being trained with 40% less compute. Our baseline model is also comparable to Mistral-7B-v0.3 and Llama 3 8B on MMLU (63% & 66%), and performs similarly on an average of 53 natural language understanding tasks while being trained with $6.6\times$ less compute than Llama 3 8B. Our results highlight the importance of dataset design for training language models and offer a starting point for further research on data curation.
Introduction
Large training datasets are an important driver of progress in the recent language modeling (LM) revolution [82, 128, 147, 57, 120, 152, 168, 62]. As the cost of training state-of-the-art language models continues to grow, researchers increasingly focus not only on scaling but also on improving training datasets that enable efficient generalization on a wide range of downstream tasks. Indeed, there is a growing number of proposals for filtering data, removing (near-) duplicates, finding new data sources, weighting data points, generating synthetic data, and so on [110, 86, 8, 173, 94, 67, 89, 2].
Figure 1: Improving training sets leads to better models that are cheaper to train. Using DataComp-LM, we develop a high-quality dataset, DCLM-baseline, which we use to train models with state-of-the-art trade-off between compute and performance. We compare on both (left) a Core set of tasks and on (right) MMLU 5-shot. Specifically DCLM-baseline (orange) shows favorable performance relative to both close-source models (crosses) and other open-source datasets and models (circles).
A key challenge in this emerging research area is a lack of controlled comparisons. While the aforementioned proposals generally use the same evaluation datasets, researchers often compare models that are trained with different architectures, compute, or hyperparameters. Hence, it is often unclear what data curation strategies work best: Are the results of training set A better than training set B because training set A is truly better, or because the model trained on A was combined with a better architecture, learning rate schedule, or more compute? Disentangling the many factors influencing the quality of a language model is crucial to understanding which data curation strategies work best and ultimately building better language models.
Beyond the lack of standardized benchmarks, another challenge for research on training data is that details about training sets are becoming increasingly rare, even for open weight models such as the Llama, Mistral, or Gemma models [157, 75, 151]. For all of these models, the training sets are not publicly available, and the corresponding model documentation only provides a coarse description of the respective training data, if any at all. As a result, it is currently unclear what ingredients constitute a state-of-the-art training set for language models.
To address these challenges, we introduce DataComp for Language Models (DCLM), the first benchmark for language model training data curation. In DCLM, researchers propose new training sets and data curation algorithms and then evaluate their datasets by training language models with a fixed training recipe on their data. By measuring the performance of the resulting model on downstream tasks, researchers can quantify the strengths and weaknesses of the corresponding training set.
Figure 2: The DCLM workflow. (A) A participant first chooses a scale, where larger scales reflect more training tokens or model parameters. (B) A participant then filters a pool of data (filtering track) or mixes data of their own (mixing track) to create a dataset. (C) Using the curated dataset, a participant trains a language model, with standardized training code and scale-specific hyperparameters, which is then (D) evaluated on 53 downstream tasks to judge dataset quality.
To enable DCLM, we contribute a comprehensive experimental testbed. A key component is DCLM-Pool, a corpus of 240 trillion tokens derived from Common Crawl [40]. DCLM-Pool is the largest public corpus for language model training and forms the cornerstone of the DCLM filtering track, where participants aim to curate the best possible training set out of DCLM-Pool. In addition, we provide open-source software for processing large datasets with several filtering approaches.
The high cost of training language models makes it necessary to understand the performance of training recipes across different compute and data scales. Hence, our third contribution is an investigation of scaling trends for dataset design. We find that models as small as 400M parameters can still provide signal on which training sets perform better at larger scales. Based on our experiments, we organize DCLM into five compute scales spanning a range of about $600\times$ in compute from 400M parameter models to over-trained 7B models. This multi-scale design makes DCLM accessible to researchers with varying compute budgets.
As a starting point for DCLM, we conduct 416 baseline experiments with different training sets and compute scales. Our experiments identify model-based filtering as a key component in an effective data curation pipeline. We also show that details of the filtering model can have a large impact on performance, ranging from 35% to 44% accuracy on MMLU 5-shot [69] at the 7B parameter scale (280B training tokens). Interestingly, a simple bigram classifier, combined with a carefully selected set of positive and negative examples, performs best among the classifiers we experimented with. In addition, we find that human quality judgments have only limited value in identifying high-quality training data.
Finally, we combine our results into DCLM-baseline, a new state-of-the-art public training set for language models. When training a 7B parameter language model on 2.6 trillion tokens using DCLM-baseline, the resulting model achieves 6464% on MMLU, which is state-of-the-art among open-data models and close to models such as Mistral-7B-v0.3 (63%) or Llama 3 8B (66%) that are trained with up to $6.6\times$ more compute (Llama 3 8B). Compared to Llama 2 7B, training a 7B parameter model on 280B tokens from DCLM-baseline achieves 5 pp higher MMLU while being trained with $7\times$ less compute. As our 7B model uses a standard decoder-only Transformer [161, 124, 157], our results also highlight that a systematic approach to data curation is key to training performant language models.
We publicly release our DCLM infrastructure, models, and training sets at https://datacomp.ai/dclm to enable other researchers to participate in DCLM and to strengthen the empirical foundations for data-centric research on language models.
Related work
We summarize closely related work in this section and provide additional related work in Appendix˜B.
Data curation for language models.
To collect large datasets for training LMs [28], researchers typically resort to web crawls, which can contain undesirable content that can be improved via curation. Most data curation efforts focus on methods for improving model performance [128, 28, 166, 125, 118, 147], including filtering by language [42, 128, 176, 84], heuristic-based filtering [57, 125, 118, 32, 147], quality filtering [166, 47, 96, 174, 136], data deduplication [3, 86] and mixing [173, 145, 6]. While prior work examines a limited set of filters, we conduct the largest public investigation of data curation, resulting in a strong DCLM-baseline dataset.
Open-source datasets.
As the scale of LMs has increased over the past years [157, 158, 4, 34, 71, 112, 125, 151], the community has curated larger datasets to match. Early works include the C4 dataset with 160 billion (B) tokens and The Pile [57] with 300B tokens. More recently, RefinedWeb [118] contains 600B tokens, Dolma [147] 3 trillion (T) tokens, FineWeb 15T tokens [119], and RedPajama-v2 30T tokens [41]. There are also large domain-specific datasets, such as the code-focused StackV2 with 900B tokens [98], as well as high-quality filtered subsets such as FineWeb-Edu [97] with 1.3T tokens. We include performance comparisons with various datasets in Figure˜1 and examine FineWeb’s LightEval evaluation framework more closely in Appendix˜G. We release the largest pool of raw text data to date with 240T web-crawled tokens. We also release DCLM-baseline, a high-quality dataset from our pool that yields better models than prior datasets.
Data-centric benchmarks.
Past work on benchmarking data improvements includes dataset distillation [44], curriculum learning [134], and transfer learning [5, 29]. In DataComp [55] and DataPerf [103], participants iterate on a dataset with a fixed model and training recipe for vision, vision-language, and speech tasks. The BabyLM challenge Loose track [163] focuses on efficient development of LMs with 125M to 220M parameters trained on 10M to 100M tokens. With a 200T token pool and 7B models, DCLM is the first large-scale data-centric benchmark for language models.
The DataComp for language models (DCLM) benchmark
Table 1: DCLM competition scales. DCLM contains five competition scales, enabling research in varying compute regimes. Each scale specifies the model size (‘Model parameters’, NN), the number of tokens seen during training (‘Train tokens’, $D$), and the size of the original pool that can be used for filtering (‘Pool size’). We provide an estimate of the compute required for training (‘Train FLOPs’=6ND=6ND) and GPU hours (‘Train H100 hours’) using the OpenLM training framework [68].
Scale Model parameters Train tokens Train FLOPs Train H100 hours Pool size
400M-1x 412M 8.2B 2.0e19 26 469B
1B-1x 1.4B 28.8B 2.4e20 240 1.64T
1B-5x 1.4B 144B 1.2e21 1,200 8.20T
7B-1x 6.9B 138B 5.7e21 3,700 7.85T
7B-2x 6.9B 276B 1.1e22 7,300 15.7T
This section describes the main components of DCLM. We start with DCLM-Pool, the raw text corpus underlying our benchmark (Section˜3.1). We then develop the DCLM workflow, visualized in Figure˜2: selecting a competition scale (Section˜3.2), curating a dataset by filtering DCLM-Pool and potentially mixing in other sources (Section˜3.3), training a model with fixed hyperparameters (Section˜3.4), and evaluating the model to score the dataset (Section˜3.5).
DCLM-Pool
DCLM-Pool is an unfiltered web-text corpus comprised of all Common Crawl [40] data prior to 2023. Based on Section˜4.2, we re-extract text from HTML using resiliparse [19, 20] instead of using Common Crawl’s pre-extracted text. DCLM-Pool contains 200B documents (370TB after gzip compression), resulting in 240T GPT-NeoX [22] tokens. See Appendix˜E for additional details.
Decontamination.
Test set samples often contaminate language model training sets [46, 177, 49]; however, the effect of such samples on downstream performance remains largely unclear [86, 147, 112]. To allow researchers to better understand contamination, we release decontamination tooling instead of decontaminating DCLM-Pool directly. Our tools are based on Lee et al. [86] and allow participants to examine their datasets for overlap with our test sets. We ask all submissions to disclose a decontamination report and avoid using highly-contaminated data. For the highest scoring submissions, we plan to specifically evaluate them for contamination. In Section˜4.6, we apply our tools to DCLM-Pool and evaluate whether contamination affects our models.
Competition scales: Supporting participants with different compute constraints
Figure 3: Datasets rank consistently across competition scales in DCLM. This makes it possible to iterate on data curation at small scales.
To ensure DCLM is accessible to researchers with different compute constraints and to facilitate the study of scaling trends, we create different competition scales spanning three orders of compute magnitude (Table˜1).Each scale (i.e., 400M-1x, 1B-1x, 1B-5x, 7B-1x, and 7B-2x) specifies the number of model parameters (e.g., 7B) and a Chinchilla multiplier (e.g., 1x). The number of training tokens for each scale is 2020 $\times$ number of parameters $\times$ Chinchilla multiplier so that a multiplier of 1x corresponds to a compute allocation that Hoffmann et al. [71] found near-optimal.
A potential pitfall in our multi-scale design is that the ranking of data curation methods may change when increasing the compute scale. To better understand this concern, in Figure˜3, we plot the performance of 10 methods at the 7B-1x scale as a function of their 400M-1x and 1B-1x performance. We find high rank correlation between the smaller 400M-1x, 1B-1x results and larger 7B-1x results (Pearson’s $r=0.885$ and $r=0.919$, respectively), suggesting better curation strategies at smaller scales transfer to larger scales. For more competition scale ablations, including experiments suggesting dataset improvements are largely orthogonal to training hyperparameters, see Appendix˜H.
Benchmark tracks: Filtering and mixing
After choosing a scale, participants choose one of two tracks. (i) In the filtering track, participants propose algorithms to select training data from a candidate pool. We start with five pools, one for each scale in Table˜1, which are random document subsets of DCLM-Pool. We restrict initial pool sizes by scale to encourage scalable filtering strategies and reflect realistic data download and storage constraints. (ii) In the mixing track, a submission combines documents from potentially many sources. For instance, participants can synthesize documents from DCLM-Pool, a custom crawl, Stack Overflow, and Wikipedia. Appendix˜C provides detailed rules for each track, and Appendix˜D describes our open-source, extensible tooling for executing filtering and mixing operations.
Training
To isolate the effect of dataset interventions, we fix a training recipe at each scale. Based on prior ablations on model architectures and training [124, 170, 56, 85, 71, 28, 34, 157, 158, 4], we adopt a decoder-only Transformer (e.g., GPT-2, Llama) [161, 157, 124], implemented in OpenLM [68]. We also provide unified data processing utilities. Appendix˜F contains additional training details.
Evaluation
Our full evaluation suite, based on LLM-Foundry [106], contains 53 downstream tasks suitable for base model evaluation (i.e., without fine-tuning): from question answering to open-ended generation formats, considering varied domains like coding, text-book knowledge, and common-sense reasoning. To evaluate data curation algorithms, we focus on three main performance metrics. First, we consider MMLU 5-shot accuracy [70], which is widely used to compare state-of-the-art models like GPT-4 [112] and Llama 3 70B [4]. Second, we propose Core centered accuracy, computed over a subset of 22 tasks (e.g., HellaSwag [182] and ARC-E [38]) that provide a low-variance signal even at small scales, linearly rescaling the accuracy per task so that 0 corresponds to random guessing and 1 corresponds to perfect accuracy. Finally, we report Extended centered accuracy, which averages the centered performance for all of our 53 tasks. For more metric details, see Appendix˜G.
Building high-quality training datasets with DCLM
We now show how the DCLM workflow can lead to high-quality datasets and quantify the effect of data curation methods. This section describes the process of converting Common Crawl into our dataset, DCLM-baseline, as shown in Figure˜4. We provide ablation experiments for each step along the way. We first evaluate open-source datasets as a starting point (Section˜4.1). Next, we experiment with alternatives for several key phases of dataset construction: text extraction (Section˜4.2), deduplication (Section˜4.3), and model-based filtering (Section˜4.4). We then experiment with mixing in high-quality sources (Section˜4.5) and provide a contamination analysis (Section˜4.6). In Section˜5, we scale up this approach to train a 7B model for 2T tokens.
Figure 4: Construction of DCLM-baseline from DCLM-Pool. Before this pipeline, we extracted DCLM-Pool from Common Crawl with resiliparse. Percentages are based on the total number of original documents.
Refer to caption Figure 4: Construction of DCLM-baseline from DCLM-Pool. Before this pipeline, we extracted DCLM-Pool from Common Crawl with resiliparse. Percentages are based on the total number of original documents.
We begin by evaluating several well-known open-source datasets (C4 [127, 46], RefinedWeb [118], RedPajama [156], and Dolma-V1 [147]) in Table˜2. While all four datasets use various heuristic filters and data cleaning steps, we find that RefinedWeb performs the best on our Core and Extended metrics at the 7B-1x scale. RefinedWeb applies the following filtering pipeline: Common Crawl text extraction, heuristic selection rules (e.g., to remove spam), and deduplication of repeated content. Interestingly, RefinedWeb is solely filtered from Common Crawl, unlike RedPajama and Dolma-V1, which additionally mix in curated, “high-quality” sources like Wikipedia. The comparison suggests the relative strength of filtering, which we explore later in our experiments.
Table 2: Comparison to existing datasets (7B-1x scale). Despite not mixing high-quality sources, RefinedWeb performs best.
Dataset Core Extended
C4 34.2 18.0
Dolma-V1 35.0 18.4
RedPajama 35.3 18.2
RefinedWeb 36.9 19.8
Evaluating existing training datasets
Takeaway: For DCLM-baseline and other experiments, we adopt RefinedWeb’s heuristic filters.
Text extraction
Text extraction is a common early processing step that pulls content from raw HTML. To understand the effect of this step, we compare three text extraction approaches: resiliparse, trafilatura (used by RefinedWeb), and the Common Crawl-provided WET files that contain pre-extracted text. We then apply RefinedWeb’s heuristic quality filters to each of the text extractions. In Table˜3, we find both resiliparse and trafilatura improve Core by at least 2.5 points over the WET extraction. This is significant because most open source datasets, including C4, RedPajama, and Dolma-V1, use the WET extraction, which could partially explain their worse performance in Table˜2. While resiliparse and trafilatura have similar downstream performance, resiliparse is $8\times$ faster to run and hence more practical for large-scale processing. For more analysis, see Appendix˜J.
Takeaway: For DCLM-Pool and the remaining experiments, we use resiliparse to extract text.
Table 3: Comparison of text extractors (1B-1x scale). We apply three approaches for text extraction from HTML, process their output using the RefinedWeb heuristic quality filters, and evaluate the quality of models trained on the resulting datasets. We find stricter extractors such as resiliparse and trafilatura are superior to WET files provided by Common Crawl.
Text Extraction Core Extended
resiliparse 24.1 13.4
trafilatura 24.5 12.5
WET files 20.7 12.2
Deduplication
Web-crawled datasets often contain many duplicate or near-duplicate data strings. Removing these duplicates from a training set serves the dual purpose of improving performance by reducing memorization [31, 86] and increasing data diversity. For deduplication, we explore MinHash [26], as part of a suffix array pipeline [86, 118], and near-duplicate Bloom filtering, which modifies an exact document and paragraph deduplication scheme [147]. We find that both approaches provide comparable downstream performance: within 0.2 Core percentage points at the 7B-2x scale. However, our modified Bloom filter approach scales more easily to datasets surpassing 10TB. We provide additional analysis in Appendix˜K.
Takeaway: We use a Bloom filter for DCLM-baseline and MinHash for other experiments.
Model-based quality filtering
Recent literature [53, 147, 25] indicates that using learnable models as quality filters leads to downstream improvements. In this section, we investigate model-based filtering.
Table 4: Quality filtering comparison (1B-1x scale). We evaluate various choices for model-based quality filters. Training a fastText classifier for filtering performs best.
Filter Core Extended
RefinedWeb reproduction 27.5 14.6
Top 20% by Pagerank 26.1 12.9
SemDedup [1] 27.1 13.8
Classifier on BGE features [172] 27.2 14.0
AskLLM [136] 28.6 14.3
Perplexity filtering 29.0 15.0
Top-k average logits 29.2 14.7
fastText [79] OH-2.5 +ELI5 30.2 15.4
Comparing model-based filtering approaches.
We compare many strategies: 1) PageRank score filtering to retain documents based on how likely they are to be linked to other documents, 2) Semantic Deduplication (SemDedup) to remove documents with similar informational content [1], 3) linear classifiers fit on pre-trained BGE text embeddings [172], 4) AskLLM that prompts an LM to see if a document is helpful [136], 5) Perplexity filtering where we retain low perplexity sequences following CCNet [166], 6) Top-k average logits where we average the top-$k$ model logits over all words in a document to score how confident a model is that the correct words are within $k$ reasonable choices, and 7) fastText [79] binary classifiers to distinguish data quality. For training classifiers, we train on $\sim 400$k documents split equally between positive and negative classes. We experiment with different options for positive data and fix negative data as a sample from RefinedWeb. For the perplexity filtering and the top-k average logits strategies, we utilize a 154M parameter causal Transformer trained on a mix of English Wikipedia, the books subset of RedPajama v1, and peS2o [146, 156]. We compare the aforementioned approaches in Table˜4 and find that fastText-based filtering outperforms all other approaches. We next aim to understand how fastText training recipes affect its effectiveness as a data filtering network [53].
Text classifier ablations.
To better understand the limits of fastText, we train several variants, exploring different choices for the reference data (i.e., the examples given positive labels), feature space, and filtering threshold, as shown in Table˜5. For reference positive data, we considered commonly used sources like Wikipedia [57], OpenWebText2 [57], and RedPajama-books [156], following the reference data used for GPT-3 [28]. We also try a novel approach, using instruction-formatted data, drawing examples from OpenHermes 2.5 [153] (OH-2.5) and high-scoring posts from the r/ExplainLikeImFive (ELI5) subreddit. Overall, we find, when controlling for other hyperparameters, the fastText OH-2.5 +ELI5 approach gives a 3.5 percentage point lift on Core compared to the conventional choices. It is natural to ask whether using OH-2.5 data for filtering could preclude additional gains from instruction-tuning. In Appendix˜P, we show this is not the case, further suggesting the strength and compatibility of this approach with modern fine-tuning paradigms. Finally, we observe that using a fairly strict threshold, which keeps the top-10% of examples, helps over more permissive top-15% and top-20% thresholds. We further study the unintuitive behavior of dataset filtering and its connection to human judgment in Appendix˜M.
Table 5: fastText ablations (7B-1x scale). We ablate choices for the positive data (top) and threshold (bottom). ‘Dataset’ is positive set for fastText, while the negatives are randomly sampled from RefinedWeb. ‘Threshold’ is the percentile used for filtering based on fastText scores. “GPT-3 Approx” refers to a uniform mix of Wikipedia, OpenWebText2, and RPJ Books, as in [28].
Dataset Threshold Core MMLU Extended
OH-2.5 + ELI5 10% 41.0 29.2 21.4
Wikipedia 10% 35.7 27.0 19.1
OpenWebText2 10% 34.7 25.0 18.7
GPT-3 Approx 10% 37.5 24.4 20.0
OH-2.5 + ELI5 15% 39.8 27.2 21.5
OH-2.5 + ELI5 20% 38.7 24.2 20.3
Takeaway: For DCLM-baseline and the remaining experiments, we use fastText OH-2.5 + ELI5 classifier score to keep the top 10% of documents. The result of this filtering is DCLM-baseline.
Dataset mixing
Researchers often combine Common Crawl (CC) with other data sources that are considered high-quality [156, 57, 63, 158] (e.g., Wikipedia, arXiv, Stack exchange, and peS2o [146]). Since DCLM participants can include additional data sources in the bring your own data track, we examined the potential benefits of adding high-quality sources to training sets derived from Common Crawl only. We compare a model trained on 100% filtered CC data to models trained with the mixing proportion from Llama 1 and RedPajama: 67% CC, and 33% from Wikipedia, Books, Stack exchange, arXiv, and Github. For the CC component, we consider different variants: a subset of our DCLM-baseline, RedPajama’s CC portion, RefinedWeb, and C4. The results in Table˜6 show that mixing improves performance for the lower-performing CC subsets (C4, RedPajama-CC, and RefinedWeb). In the case of DCLM-baseline however, mixing actually hurts performance on average, which suggests it can be counterproductive given performant filtering. For additional mixing results, see Appendix˜L.
Table 6: Mixing high-quality sources with subsets of CommonCrawl (1B-1x scale). We evaluate the impact of mixing high-quality source (‘RPJ extras’) to various datasets derived from CommonCrawl, using the mixing ratios from Llama and RedPajama. Numbers in parentheses indicate the improvement/degradation due to mixing, compared to using only the base dataset.
Core Extended
Dataset Base w/ RPJ extras Base w/ RPJ extras
C4 23.7 25.9 (+2.2+2.2) 12.5 13.3 (+0.8+0.8)
RPJ CC only 24.0 25.7 (+1.7+1.7) 12.1 13.5 (+1.4+1.4)
RefinedWeb 25.1 26.5 (+1.4+1.4) 12.9 13.1 (+0.2+0.2)
DCLM-baseline  [^1] The DCLM-baseline subset used here relied on an earlier version of our classifier and our pool. 31.1 29.5 (1.6-1.6) 16.0 15.3 (0.7-0.7)
Decontamination
Here, we perform analysis to examine whether contamination of our pretraining data with our evaluation is an issue that influences our results. We focus on MMLU as our evaluation set of choice, given its popularity as a metric for language model performance at this scale.
As an experiment, we also attempt to detect and remove questions from MMLU that exist in DCLM-baseline. Our strategy is to flag training documents that contain the last sentence of a question from MMLU along with one of the corresponding options. For these flagged examples, we then remove all matched question and option strings. In order to improve recall, we opt to detect only the last sentence from each question, reducing the chance of missing questions due to formatting differences. Based on inspection, this also incurs many false positives. We then train a 7B-2x model with our DCLM-baseline without the detected MMLU overlap.
The results of this analysis can be seen in Table˜7. We see that this removal of contaminated samples does not lead to a decrease in performance for our model. As such, we can see that our performance gains in MMLU are not caused by increased presence of MMLU in our dataset.
Table 7: MMLU overlap removal results. We remove overlaps detected with MMLU, in cases where a question and one of its options are detected in the text. We compare the performance between a model trained with and without this data removed, and see that there is no gain from increased contamination. This experiment is done at the 7B-2x scale.
Dataset MMLU
DCLM-baseline 51.8
DCLM-baseline (MMLU removed) 52.7
We also apply the above removal strategy on Dolma-V1.7 [147] and FineWeb-Edu [97], to measure what the contamination differences are between DCLM-baseline and those datasets. The results can be seen in Table˜8. We see that our DCLM-baseline has roughly similar contamination stats as other high performing datasets according to this analysis. We provide further contamination analysis that extends to our entire benchmark suite in Appendix˜N.
Table 8: MMLU overlap removal comparison. We remove overlaps detected with MMLU, in cases where a question and one of its options are detected in the text. For Dolma-V1.7 [147], we sample 1/10th of the dataset for this analysis (roughly 230B tokens). For FineWeb-Edu [97], we use the 10B token subset released by the authors. Note that because our flagging rule prioritizes recall over precision, these numbers are likely to be overestimates of the true contamination rates.
Dataset Percentage of samples flagged
DCLM-baseline 0.007%
Dolma-V1.7 0.001%
FineWeb-Edu 0.009%
Scaling up DCLM-baseline to the trillion token scale
Here, we test if datasets that perform well on the DCLM benchmark also maintain their strength with an order of magnitude more compute. To ensure our trained model is broadly useful, including for math and coding tasks, we combine our 3.8T DCLM-baseline with the StarCoder [88] and ProofPile2 [13] data to arrive at a 4.1T token dataset. We train a 7B model for 2.5T tokens on this dataset with the same hyperparameters as our largest competition scale except for an annealing phase for the 200B tokens on a modified distributed that was 70% DCLM-baseline with a tighter fastText threshold, and 30% math datasets (see Appendix˜P). We then adopt the continual pre-training methodology from Pouransari et al. [123] for 100B tokens on the same distribution to increase the context length from 2048 to 8192, we provide more details on this procedure in Section˜P.2.
In Table˜9, we show that our model outperforms all 7B models trained on public training sets and approaches closed-data models trained for more tokens such as Llama-8B, Mistral-7B, and Gemma-7B. Additionally, in Appendix˜O, we show that our model achieves strong instruction-tuning performance. After instruction tuning on publicly available IT datasets, our model maintains most of its benchmark performance and achieves an AlpacaEval2.0 LC Win-rate of 16.6, which outperforms Gemma-Instruct (10.4), while approaching the strong performance of Mistral-v0.2-7B (17.1) and Llama3-Instruct (22.9).
Table 9: State-of-the-art comparison (beyond 7B-2x scale). We compare our final model with others in the 7–8B parameter regime. Our DCLM-baseline dataset yields a model that outperforms models trained on open datasets and is competitive with models trained on private datasets.
Model Params Tokens Open dataset? Core MMLU Extended
Open weights, closed datasets
Llama2 7B 2T 49.2 45.8 34.1
DeepSeek 7B 2T 50.7 48.5 35.3
Mistral-0.3 7B ? 57.0 62.7 45.1
QWEN-2 7B ? 57.5 71.9 50.5
Llama3 8B 15T 57.6 66.2 46.3
Gemma 8B 6T 57.8 64.3 44.6
Phi-3 7B ? 61.0 69.9 57.9
Open weights, open datasets
Falcon 7B 1T 44.1 27.4 25.1
OLMo-1.7 7B 2.1T 47.0 54.0 34.2
MAP-Neo 7B 4.5T 50.2 57.1 40.4
Models we trained
FineWeb edu 7B 0.14T 38.7 26.3 22.1
FineWeb edu 7B 0.28T 41.9 37.3 24.5
DCLM-baseline 7B 0.14T 44.1 38.3 25.0
DCLM-baseline 7B 0.28T 48.9 50.8 31.8
DCLM-baseline 7B 2.6T 57.1 63.7 45.4
Conclusion & limitations
We introduced the DCLM testbed and demonstrated how it leads to new state-of-the-art training sets. Our exploration of the dataset design space is only the beginning and has clear limitations. Due to compute constraints, we could only ablate design dimensions individually and could not test all approaches on larger scales. Moreover, there are many variations of DCLM-baseline that we did not explore. For instance, understanding the impact of sharded deduplication in more detail is important, and there are many more ways of training filtering models, both in terms of their architecture and training data. We also conducted most of our experiments with only one tokenizer (GPT-NeoX), and other tokenizers may perform better on multilingual tasks or math. Another limitation is that we could not sufficiently explore the run-to-run variation from different random seeds. Still, we hope that this paper is a starting point for further research on data curation that pushes the state-of-the-art beyond DCLM-baseline.
While our models trained on DCLM-baseline are competitive on common language understanding evaluations, they currently do not perform as well on code and math. We view this as a consequence of our focus on language understanding in the first version of DCLM, and not an inherent limitation of our benchmark or the DCLM-baseline training set. Indeed, prior work has shown that adding specific training data and post training methods for code and math can substantially improve performance on those domains [188, 184, 13, 88, 165]. Combining DCLM-baseline with these domain-specific training sets and extending DCLM to cover code and math are intresting avenues for future work.
There are other important performance dimensions our evaluation suite currently does not incorporate such as fairness, multilinguality, and safety. Similarly, studying toxicity or privacy filtering in the context of DCLM would be fruitful. Expanding DCLM along these dimensions is a fruitful direction for future work, and we hope that our open and accessible testbed can strengthen the foundations of data-centric research in these directions as well.
Lastly, we have only trained 7B parameter models as part of DCLM. In contrast, state-of-the-art language models are now substantially larger. While we are optimistic that our gains will also extend to larger model scales, future work still needs to test this experimentally. One possible limitation of the approach behind DCLM-baseline may be its stringent filtering ratio. After an exact global deduplication at the document level, DCLM-baseline contains approximately 2T tokens, and after removing all near-duplicates globally, about 1T tokens remain. Understanding the interaction between data quality, filtering ratio, deduplication, and multi-epoch training will be key for building larger training sets in the future.
Acknowledgements.
We would like to thank Lilith Bat-Leah, Loubna Ben Allal, Samy Bengio, Mia Chiquier, Adrien Gaidon, Lizzy Grant, Tom Gunter, Awni Hannun, Jonathan Hayase, Mike Lewis, Percy Liang, Ian Magnusson, Yifan Mai, Sewon Min, David Mizrahi, Praveen Paritosh, Guilherme Penedo, Kyle Richardson, Weijia Shi, Karanjeet Singh, Joshua Susskind, Oyvind Tafjord, Carl Vondrick, and Elle Wohlmuth, for helpful feedback at various stages of the project. We would like to thank Mike Garrison and Romil Shah for help with compute and infrastructure.
This research was supported by Allen Institute for AI, Open Philanthropy, Institute for Foundations of Machine Learning (IFML), AFOSR MURI grant FA9550-22-1-0380, Israeli Science Foundation (ISF) grant no. 2486/21, Alon Fellowship, Adelis foundation, Israeli Council for Higher Education, Onassis Foundation - Scholarship ID: F ZS 056-1/2022-2023, NSF Grants AF 1901292, CNS 2148141, Tripods CCF 1934932, IFML CCF 2019844,, research gifts by Western Digital, Amazon, WNCG IAP, UT Austin Machine Learning Lab (MLL), Cisco and the Stanly P. Finch Centennial Professorship in Engineering, NSF Graduate Research Fellowship. MN acknowledges funding by the Federal Ministry of Education and Research of Germany under grant no. 01IS22094B WestAI - AI Service Center West.
We gratefully acknowledge compute budget granted by Gauss Centre for Supercomputing e.V. and by the John von Neumann Institute for Computing (NIC) on the supercomputers JUWELS Booster and JURECA at Jülich Supercomputing Centre (JSC)
Appendix
Appendix A Contributions
All authors are listed alphabetically by last name.
Data processing
DCLM-baseline
Khyathi Chandu, Alex Fang, Saurabh Garg, Thao Nguyen, Vaishaal Shankar
Decontamination
Achal Dave, Saurabh Garg, Jeffrey Li, Vaishaal Shankar, Georgios Smyrnis
Deduplication
Achal Dave, Alex Fang, Jeffrey Li, Matt Jordan, Vaishaal Shankar
Extra data sources
Yonatan Bitton, Mayee Chen, Giannis Daras, Achal Dave, Alex Fang, Joshua Gardner, Maciej Kilian, Jeffrey Li, Niklas Muennighoff, Marianna Nezhurina, Vaishaal Shankar, Hanlin Zhang
Baseline datasets
Pre-processing existing datasets
Achal Dave, Alex Fang, Samir Gadre, Reinhard Heckel, Sedrick Keh, Vaishaal Shankar, Georgios Smyrnis
Reproductions
Amro Abbas, Hritik Bansal, Yonatan Bitton, Yair Carmon, Khyathi Chandu, Alex Fang, Dhruba Ghosh, Cheng-Yu Hsieh, Maor Ivgi, Matt Jordan, Sedrick Keh, Jeffrey Li, Kyle Lo, Luca Soldaini, Hanlin Zhang, Jieyu Zhang
Data curation
Filtering
Hritik Bansal, Alex Fang, Saurabh Garg, Maor Ivgi, Matt Jordan, Jeffrey Li, Vaishaal Shankar
Rewriting
Saurabh Garg, Maor Ivgi, Niklas Muennighoff
Mixing
Achal Dave, Jeffrey Li, Georgios Smyrnis
Evaluations
Evaluation framework
Alon Albalak, Kushal Arora, Hritik Bansal, Achal Dave, Maor Ivgi, Sedrick Keh, Vaishaal Shankar, Rulin Shao, Rui Xin
Evaluation runs
Kushal Arora, Achal Dave, Alex Fang, Jeffrey Li, Sedrick Keh, Vaishaal Shankar
Evaluation metrics
Achal Dave, Alex Fang, Jeffrey Li, Vaishaal Shankar
Training
Training code
Achal Dave, Samir Gadre, Suchin Gururangan, Kalyani Marathe, Jean Mercat, Hadi Pouransari, Sunny Sanyal, Georgios Smyrnis, Igor Vasiljevic, Mitchell Wortsman
Data preprocessing
Alex Fang, Matt Jordan, Vaishaal Shankar, Georgios Smyrnis
Training runs
Kushal Arora, Achal Dave, Alex Fang, Samir Gadre, Sedrick Keh, Jeffrey Li, Vaishaal Shankar, Georgios Smyrnis
Trillion-token models
Alex Fang, Jeffrey Li, Hadi Pouransari, Vaishaal Shankar
Instruction tuning
Kushal Arora, Hritik Bansal, Aaron Gokaslan, Etash Guha, Niklas Muennighoff
Competition setup
Competition design
Yair Carmon, Achal Dave, Alex Fang, Samir Gadre, Reinhard Heckel, Jeffrey Li, Ludwig Schmidt, Vaishaal Shankar
Competition tooling
Gabriel Ilharco, Maor Ivgi, Jeffrey Li, Sarah Pratt
Paper writing
Alon Albalak, Hritik Bansal, Yair Carmon, Achal Dave, Alex Dimakis, Alex Fang, Samir Gadre, Etash Guha, Reinhard Heckel, Maor Ivgi, Sedrick Keh, Jeffrey Li, Niklas Muennighoff, Sarah Pratt, Ludwig Schmidt, Vaishaal Shankar, Georgios Smyrnis
Leadership and advising
Advising
Alaa El-Nouby, Fartash Faghri, Dirk Groeneveld, Reinhard Heckel, Jenia Jitsev, Sham Kakade, Pang Wei Koh, Thomas Kollar, Kyle Lo, Niklas Muennighoff, Sewoong Oh, Sujay Sanghavi, Luca Soldaini, Shuran Song, Alexander Toshev, Stephanie Wang, Luke Zettlemoyer
Leadership
Yair Carmon, Achal Dave, Alexandros G. Dimakis, Ludwig Schmidt, Vaishaal Shankar
Project coordination
Achal Dave, Ludwig Schmidt, Vaishaal Shankar
Appendix B Additional related work
Data curation methods have been proposed that can be grouped into two categories: methods that aim to enhance performance, and those with non-performance related goals. Performance-oriented methods include language detection, heuristics-based filtering, quality filtering, data deduplication, data mixing, and synthetic data. Non-performance-oriented filters include the removal of copyrighted text, toxic content, personally identifiable information (PII), opt-out, and evaluation data.
Language detection.
Language detection methods most often rely on a fastText classifier that has been trained to identify 157 languages [64, 118, 147], but past methods have also utilized other classifiers including a naive Bayes classifier [128]. When collecting multilingual datasets, another curation option is to filter web pages based on country domains or by selecting URLs that are correlated with data from certain languages [100, 144, 121].
Heuristics.
It is widely known that web-scraped text contains high quantities of boilerplate HTML, error messages, stock tickers, and other undesirable data for training language models, much of which can be detected and removed by heuristic-based systems. The exact heuristics used by each data curation method vary but can be grouped into five categories: item count, repetition count, existence, ratio, and statistics. For example, Rae et al. [125] remove any lines that contain at least two of the following stop words: the, be, to, of, and, that, have, with, defining an item count heuristic. An example of a statistic might be removing documents that have a mean line length greater than 100 characters [32].
Quality filtering.
Filtering for “high-quality” data (data which was written by humans and has likely gone through an editing process [8]) is a common step for data curation pipelines. The most commonly used method for quality filtering is to train a binary classifier on data from a perceived high-quality dataset (e.g. Wikipedia) and a perceived low-quality dataset (e.g. unfiltered web text) and filter out data where the classifier assigns sufficiently low scores [57, 28, 47]. A less commonly used method is train a language model on the high-quality dataset and calculate perplexity on the data to be filtered, where high perplexity scores suggest that the data is lower quality [166, 110].
Recently, works have proposed the use of pretrained language models to identify and curate high-quality data through prompting for various dimensions of perceived quality [167, 136, 97]. Ankner et al. [11] even find that it’s possible to use a small pretrained model (125M parameters) to prune training data for models as large as 3B. MiniPile [80] demonstrated that a 1M document subset of the pile, selected by clustering and removing low-quality clusters, can lead to small LMs that maintain performance on GLUE, while significantly reducing the scale of training data. RHO-1 [90] has a similar goal to quality filtering, but rather than filtering data out of the dataset, they propose Selective Language Modeling, an objective function that selectively masks the loss of tokens that are predicted to be low quality.
Deduplication.
Deduplication has proven to be a beneficial step in almost every data curation pipeline. The methods used vary in complexity, including deduplication based on URLs, hashing, string metrics, and using model-based representations. URL deduplication has long been in use for deduplicating web snapshots [3]. Commonly used hash-based deduplication methods include Bloom filters [23, 147], suffix array-based methods [86, 118], and MinHash-based methods [27] such as MinHashLSH [28]. Model-based methods include SemDeDup [1] which embeds each point in a dataset, clusters data points together, and removes data points within clusters that are too similar, and D4 [155] which further applies the SSL prototypes method from Sorscher et al. [148] and removes the most prototypical example from each cluster.
Data mixing.
When the training dataset is composed of data from multiple domains or sources (e.g. web text, Wikipedia, and books), then an additional challenge for data curation is to determine what percent of the final dataset comes from each source, known as data mixing. Methods for data mixing include using heuristics (such as human judgment) [57, 157], or empirically determining the best domain weights according to some downstream evaluation [47, 125]. More principled approaches have been proposed that are based on Group DRO [173], multi-armed bandits [7], and information theory [6]. Further methods have been proposed building off of these principled approaches, including DoGE [52], Skill-it [33], and ShearedLlama [171], each bringing some improvements. Thudi & Maddison [154] develop MixMax, a provably optimal method under a concave objective, which improves upon Group DRO-based alternatives but has not been proven at scales typical of language modeling. Ge et al. [58] propose BiMix, a unified scaling law that simultaneously models the behaviors of data quantity and mixing weights, using only small models as proxies to calculate the scaling laws.
Synthetic data.
With the improvements in the ability of language models to accurately model text distributions, the generation of synthetic data has become an additional avenue for data curation. Notable methods for pretraining include the Phi models [67, 2], which generate synthetic textbook data from the GPT series of models, as well as WRAP [101] which uses a similar method to the Phi models, but demonstrates that the synthetic data generation pipeline is feasible with much smaller models (1.8B and 7B parameters). Beyond generating synthetic data for pretraining, Singh et al. [143] propose ReSTEM, a method for generating synthetic data for math and coding benchmarks, which uses binary feedback (eg. whether the code gives the correct output) to repeatedly filter self-generated data. Similarly, Zelikman et al. [181] propose STaR, which bootstraps a dataset of rationales for commonsense question answering and mathematics datasets.
Non-performance related methods have been designed for a variety of purposes, including to remove copyrighted content [95, 141, 142], toxic speech [128, 160], private information [9, 98], opt-out data [98, 109] or to decontaminate data to avoid benchmark leakage [178]. While these methods are less relevant to this work, they are nonetheless important in real-world data curation pipelines.
Appendix C Benchmark rules
This section provides detailed guidelines for submissions in the two DCLM tracks.
General rules
The following applies to both the filtering and mixing tracks.
  1. 1.
    Submissions should include documentation detailing their key components.
  2. 2.
    The dataset underlying a submission to the leaderboard, or fully working code to reproduce it, should be freely available to encourage reproducibility of submissions. Submissions that do not satisfy this requirements may still be accepted, but we will mark them as such in the leaderboard.
  3. 3.
    Tokenization must be performed with our provided script that tokenizes the data and performs a global shuffle.
  4. 4.
    Submissions cannot make any changes to the training or evaluation code.
  5. 5.
    Use of evaluation data (test data from our evaluation tasks) for purposes other than evaluation and decontamination is forbidden.
Filtering track
The defining characteristic of entries in the filtering track is that they form the dataset by applying a processing pipeline on the subset of DCLM-Pool corresponding to the chosen compute scale (see Table˜1) without including any external data. The rationale behind this requirement is twofold. First, the size and quality of initial data for filtering affects both the processing cost and the quality of the processed dataset. By fixing the initial dataset we level the playing field and allow comparison to focus on core curation techniques. Second, we wish to encourage the development of methods potentially relevant even at frontier-model scale. Using the 7B-2x pool (containing roughly 16T tokens) for the 400M-1x compute scale (requiring roughly 8B tokens for training) would allow filtering strategies that keep less than 0.1% of the data and cannot scale to generating a trillion-token dataset.
As we wish to encourage creative and performant submissions, our requirement for using only DCLM-Pool comes with the following qualifications:
  1. 1.
    Modifying HTML extraction. We create DCLM-Pool by extracting text from Common Crawl archives using resiliparse, which eases the computational burden on participants who may not have resources to extract text themselves. However, we additionally specify the Common Crawl archives for each pool to allow experimentation with text extraction. Participants may either start with our parsed DCLM-Pool data or work directly with the relevant Common Crawl WARC archives.
  2. 2.
    Using models trained on external data. We allow the DCLM-Pool processing pipeline to leverage models for quality filtering, paraphrasing, etc. These models may be trained on external data with the exception of evaluation data as per the general guidelines. We will not accept submissions abusing this allowance to introduce external data via a backdoor, e.g., by “paraphrasing” documents from DCLM-Pool into memorized data.
Mixing track
In the mixing track, participants are free to use any data source, provided it meets the general guidelines by being freely available and not including evaluation data. Submissions to the mixing track should clearly document their data sources, the weight given to each source, and the ratio of tokens used for training (fixed for every benchmark scale) to the overall custom pool size.
Appendix D Tooling
Download.
For the construction of our pool, we download WARC files from Common Crawl, and process them via resiliparse, we do this by streaming data directly from S3 to EC2 using the Ray data processing framework. This is the starting point for our data processing pipeline. For the dataset released to participants, we release various sizes of DCLM-Pool, that we make available for download. For details on the data, see Appendix˜E.
Processing.
Given raw pool of text, it is often useful to define a processing pipeline to clean, modify and filter it. We provide a robust framework to do that at scale, by sharding the pool and processing it in parallel. Namely, to process a pool one needs to define a sequence of Mappers, each taking a single document with its associated metadata as input, and output a list of documents. Our mappers include:
  1. 1.
    Filters which either retain or discard the input document according to some filtering criteria such as having a maximum or minimum length.
  2. 2.
    Enrichers which always return a list of with the page as is, adding additional information to the metadata, such as detected language or number of tokens.
  3. 3.
    Modifiers change the content of the text itself, and can also split the document to create several new documents. This is useful for example, as a participant may design a function to remove padding white-space.
In particular, we implement all mappers used in RefinedWeb (which includes those from Gopher as a subset) and C4 along with many new ones, and allow users to integrate custom mappers into their pipeline. Additionally, while mappers allow for document-level processing, in some cases it may also be necessary to execute corpora-level operations. For instance, a user may wish to deduplicate spans that appear in several documents. Our tooling also supports global functions that depend on all documents.
Contamination Analysis.
We use the tools provided by Lee et al. [86] as a base and adapt them to evaluate the contamination of our training set with the evaluation sets. As done in Touvron et al. [158], we measure the number of tokens that appear in the same consecutive sequence of at least 10 tokens, between a training sample and an evaluation sample. With this number, we calculate how many tokens on average per evaluation sample are “contaminated”, appearing both in the training and the evaluation data.
Tokenization and shuffling.
Once documents have been mapped, filtered, or globally processed, we provide standardized code to tokenize and shuffle data. The output of this code is a trainable dataset artifact. For tokenization, our code uses the GPT-NeoX [22] tokenizer. Our tokenization code adopts Ray  [^2] https://github.com/ray-project/ray for cluster management and scales from a single node setups for small datasets to multiple nodes for larger ones. After tokenizing, we perform a global shuffle of our dataset.
Training Setup.
We base our training code on OpenLM [68], and provide configuration files for each of our scales. We also provide scripts that train models using each configuration, and produce json files that describe a trained model in detail. For further training details, see Appendix˜F.
Evaluation.
We base our evaluation pipeline on the evaluation tasks provided by LLM-foundry [107]. Using one of the aforementioned model json files as input, our tools evaluate the associated checkpoint on all of our tasks. A new json file is then produced, including the evaluation results in each task, as well as aggregate metrics. This json file can then be submitted via a pull request to submit the results to our leaderboard.
Reproducibility.
All of our results, including data processing, model training, evaluations, and plots included in this paper, are reproducible using our open-source framework and the recipes in https://datacomp.ai/dclm. We list compute requirements for our code in Appendix˜Q.
Appendix E DCLM-Pool
DCLM-Pool was collected by taking all 5.1M Common Crawl WARC dumps from 2008 to 2022 (inclusive) and extracting text from the html using the resiliparse framework. We opted to omit 2023 and above to prevent large amounts of language model generated text from polluting our datasets and to provide a hold out for future use. We release DCLM-Pool on HuggingFace with CC-BY-4 license. We release DCLM-Pool as a set of .jsonl files similar to Dolma-V1 and RedPajama. We provide the fields that are in the .jsonl in Table˜10. The entire pool is 5.1M gzip compressed .jsonl\ files, and 340TB compressed on disk. The use of this dataset is also subject to CommonCrawl’s Terms of Use: https://commoncrawl.org/terms-of-use.
Common Crawl respects robots.txt, and thus our pool does so as well, giving content creators a mechanism to opt out of Common Crawl and DCLM-Pool. Since DCLM-Pool is a large subset of Common Crawl it will contain some PII data, however Common Crawl does honor deletion requests and periodically redacts dumps. We designed DCLM-Pool to maintain a one-to-one mapping between raw Common Crawl WARC files and DCLM-Pool .jsonl files, allowing us to update DCLM-Pool based on redactions.
We note that Common Crawl includes raw data as collected from the web without filtering. While some of our pools, such as DCLM-baseline, underwent some filtering of malicious URLs, none have had any special treatment for PII and sensitive content to preserve representativeness of the raw data. For a more complete discussion on PII and consent regarding our pools, see Appendix˜S.
Table 10: Metadata provided in DCLM-Pool data.
Label Additional notes
metadata.Content-Length Length of the content.
metadata.Content-Type Type of the content.
metadata.WARC-Block-Digest Digest for data integrity.
metadata.WARC-Concurrent-To Related WARC record.
metadata.WARC-Date Date of the WARC record.
metadata.WARC-IP-Address IP address of the source.
metadata.WARC-Identified-Payload-Type Identified payload type.
metadata.WARC-Payload-Digest Payload digest for integrity.
metadata.WARC-Record-ID Unique ID of the WARC record.
metadata.WARC-Target-URI Target URI of the record.
metadata.WARC-Type Type of WARC record.
metadata.WARC-Warcinfo-ID Related warcinfo record ID.
text Text content.
url URL of the source.
warcinfo Information about the WARC file.
Appendix F Training details
Overview.
Our training setup follows closely that of Wortsman et al. [170] and Gadre et al. [56]. Specifically, we build our training infrastructure using the OpenLM [68], which supports decoder-only, pre-normalization Transformers [161], following an architecture inspired by GPT-2 [124] and Llama [157]. OpenLM is a PyTorch [115, 12] code-base that targets FSDP modules for distributed training [186].
Architecture details.
We utilize LayerNorm [14] without bias parameters for all normalization, qk-LayerNorm [45] on queries and keys for training stability, SwiGLU [140] multilayer perceptrons (MLPs), and a depth-scaled initialization scheme following Zhang et al. [183]. Our sequence length, during pre-training is 2048. We pack multiple sequences into batches to fill the entire context, with an EOS token to split documents. We allow causal attention to attend across documents; we experimented with masking attention across documents but early experiments indicated little impact on downstream performance.
Training sets and tokenization.
Since the focus of our paper is dataset development, we train on over 270 data distributions, mostly filtered from Common Crawl. For the majority of our experiments we use GPT-NeoX [22] for tokenization, which yields a vocabulary size of 50k.
Optimization details.
As metioned in the main body, we train with a standard next-token prediction objective. Following Chowdhery et al. [34], we employ z-loss to encourage output logit magnitudes to remain in a numerically stable range.
Hyperparameters.
We detail the hyperparameters for our models in Table˜11. For the 400M-1x and 1B-1x, we follow hyperparameters from [56], which were tuned to optimize perplexity on a validation set containing tokens from recent arXiv papers, the OpenLM codebase itself, and news articles. For the 1B-1x scale, we also investigated alternative hyperparameters in Table˜13, and find the hyperparameters from [56] perform best. For the 7B-1x and 7B-2x, we used a higher learning rate, and a lower weight decay, guided by the hyperparameter sweep in Table˜12. We use a cooldown of 3ee-5 for all experiments. For Table˜2, we trained with a lower learning rate following [56] as these experiments were performed before our sweep. Specifically, we used a learning rate of 3ee-4 and weight decay of 0.33.
Table 11: Main models and hyperparameters used in our investigation. For each scale, we list the number of layers $n_{\text{layers}}$, number of attention heads $n_{\text{heads}}$, model width $d_{\text{model}}$, and width per attention head $d_{\text{head}}$. Batch sizes are global and in units of sequences. Each sequence has 2,048 tokens.
Scale $n_{\text{layers}}$ $n_{\text{heads}}$ $d_{\text{model}}$ $d_{\text{head}}$ Warmup Learning rate Weight decay z-loss Batch size
400M-1x 24 8 1,024 128 2,000 3ee-3 0.033 1ee-4 512
1B-1x 24 16 2,048 128 5,000 3ee-3 0.033 1ee-4 256
7B-1x, 7B-2x 32 32 4,096 128 5,000 2ee-3 0.05 5e-6 2,048
Table 12: Learning rate and weight decay sweep (7B-1x scale). We evaluated the impact of learning rate and weight decay on an earlier iteration of DCLM-baseline. Based on this sweep, we specify the settings for Table˜11 for the 7B-1x and 7B-2x scales.
LR WD Core
1ee-03 0.1 44.1
2ee-03 0.05 44.8
3ee-03 0.033 44.7
1ee-02 0.01 43.8
Appendix G Evaluation details
Below we outline the tasks we used to evaluate our models in LLM Foundry. We also examine the LightEval [54] evaluation pipeline used in the FineWeb-Edu [97] evaluations.
Evaluation Tasks
We divide our evaluations into two high-level categories: Core (22 tasks) and Extended (53 tasks). The set of Core tasks were selected due to their ability to provide a low variance signal of learning, even at small scales. We include a diverse range of tasks aimed at assessing a variety of model capabilities.
Core tasks.
  • The AGI Eval LSAT-AR dataset [189] (3-shot) tests for model knowledge in the legal domain and evaluates analytical reasoning capabilities.
  • The ARC easy and ARC challenge datasets [38] (10-shot) contain four-way multiple choice questions taken from grade 3-9 science exams, where questions in the easy dataset require knowledge of basic science, and the challenge questions require some procedural reasoning.
  • We use a series of 6 datasets from Big-Bench [17] (all 10-shot): (1) QA Wikidata which requires models to complete factual statements with the correct answer, (2) Dyck languages where the model needs to complete a partially balanced expression consisting of parentheses and braces, (3) Operators where the model is given some newly defined operators and asked to compute the output from some expression using those operators, (4) Repeat Copy Logic which requires the model to differentiate instructions from text-to-copy and to perform a sequence of operations, (5) CS Algorithms which requires the model to execute algorithms such as recursion and dynamic programming, and (6) Language Identification where the model is expected to identify the language of a sequence of natural language text.
  • BoolQ [36] (10-shot) is a binary question answering dataset where the model is expected to answer questions about relevant passages.
  • CommonsenseQA [150] (10-shot) is a 5-way multiple choice question answering dataset which evaluates the models ability to understand and apply commonsense knowledge on everyday scenarios.
  • COPA [133] (0-shot) consists of causal reasoning questions where the model is given two possible outcomes to a scenario and must use commonsense to select the outcome that is more likely.
  • CoQA [131] (0-shot) is a conversational question answering dataset where the model is given a passage and conversation between two participants and then expected to extract an answer from the passage to a question from one of the participants.
  • HellaSwag [182] (0-shot and 10-shot) is a 4-way multiple choice commonsense reasoning dataset, where the model is required to understand implicit context and common knowledge in order to correctly select the continuation to a context.
  • Jeopardy [81] (10-shot) is a dataset of questions posed in the format of the “Jeopardy!” quiz show, covering a wide variety of topics.
  • LAMBADA [113] (0-shot) is a collection of narratives where a human is able to guess the final word of the narrative, but is not able to if they are only given the final sentence. To perform well on this task requires the model to attend to context from the full narrative and cannot simply rely on the local context.
  • OpenBookQA [105] (0-shot) is a 4-way multiple choice question answering dataset that requires the model to use multi-step reasoning and commonsense knowledge.
  • PIQA [21] (10-shot) is a binary multiple choice question answering dataset that requires the model to use physical commonsense reasoning to answer correctly.
  • SQuAD [130] (10-shot) is a question answering dataset where the model is given a question and a passage containing the answer to that question.
  • The Winograd Schema Challenge [87] (0-shot) is binary multiple choice pronoun resolution task where the model is given a context and asked to determine which entity a pronoun refers to, requiring the model to exhibit commonsense knowledge and contextual understanding.
  • The Winogrande [137] (0-shot) dataset extends the Winograd Schema Challenge dataset by expanding the dataset to a wider variety of domains.
Extended tasks.
  • We use a series of 4 additional tasks from the AGI Eval suite of datasets [189] (all 3-shot): (1) LSAT-LR and (2) LSAT-RC test for model knowledge in the legal domain and evaluate logical reasoning and reading comprehension, respectively, (3) SAT-En evaluates the model’s capabilities in English, and (4) SAT-Math evaluates the model’s capability in math using chain-of-thought prompting.
  • AQuA [91] (3-shot) is a 4-way multiple choice question answering dataset that evaluates the model on algebra questions using chain-of-thought prompting.
  • BBQ [114] (3-shot) is a multiple choice question answering dataset designed to detect model’s biases along nine social dimensions.
  • We use a series of 9 additional datasets from Big-Bench [17] (all 10-shot): (1) Conceptual Combinations which evaluates the model’s capability to parse conceptual combinations by selecting sentences where these combinations are used correctly, (2) Conlang Translation where the model is expected to deduce a new translation from English to an obscure constructed language based on a limited number of translation examples, (3) Elementary Math QA which is a multiple choice question answering dataset of simple quantitative reasoning problems, (4) Logical Deduction which requires a model to parse, understand, and apply information about objects and relationships between objects to infer new information, (5) Misconceptions evaluates whether a model can discern popular misconceptions from truth, (6) Novel Concepts measures the models ability to creatively construct a necessary abstraction that is unlikely to have existed in training data, (7) Strange Stories measures a model’s capacity for Theory of Mind, (8) Strategy QA is a test that requires a model to answer questions requiring multi-step implicit reasoning, (9) Understanding Fables which evaluates the model’s capability to understand the moral of a short story.
  • Enterprise PII classification [117] (10-shot) is a binary classification task that evaluates whether a model can detect PII (e.g. usernames, emails) within text.
  • GPQA-main and GPQA-diamond [132] (5-shot) are 4-way multiple choice question answering datasets written by domain experts in biology, physics, and chemistry, which are intended to be very difficult for non-experts to answer (even with access to the web). The diamond set is a high-quality subset including only questions where two experts answer correctly, but most non-experts answer incorrectly.
  • GSM8K [39] (3-shot) is a dataset of grade school math word problems that requires between 2 to 8 steps to solve, where the model uses chain-of-thought prompting.
  • LogiQA [92] (10-shot) is a 4-way multiple choice question answering dataset that evaluates logical reasoning.
  • Math QA [10] (10-shot) is a 5-way multiple choice question answering dataset that evaluates math word problem solving capabilities, built on top of AQuA.
  • MMLU [70] (0-shot and 5-shot) is a 4-way multiple choice question answering dataset that covers 57 different domains and tasks, evaluating both world knowledge and problem solving capabilities.
  • PubMedQA [76] (10-shot) is a 3-way multiple choice question answering dataset which evaluates the model’s ability to answer biomedical research questions given context from a relevant research article.
  • Simple arithmetic with spaces and without spaces [107] (10-shot) are datasets consisting of simple arithmetic problems with up to 3 operations using numbers with up to 3 digits, evaluating a model’s ability to follow the correct order of operations and perform arithmetic.
  • Social Interaction QA [138] (10-shot) is a binary multiple choice question answering dataset that evaluates a model’s social commonsense intelligence.
  • SVAMP [116] (3-shot) is a set of challenging elementary-level math word problems that uses chain-of-thought prompting.
  • Trivia QA [78] (3-shot) is an open-ended question answering dataset that evaluates the world knowledge of a model.
  • The Winogender male and Winogender female datasets [135] (10-shot) are variants of the winograd schemas method that creates a minimal pair of sentences that differ only by the gender of one pronoun, designed to evaluate a model’s gender bias.
  • HumanEval [32] (0-shot) is a code completion dataset composed of hand-written problems where the model is given a function signature, and docstring and expected to correctly produce a function that will pass several unit tests.
LightEval
Given the multiple ways of evaluating accuracy [51], we conducted a miniature study using the LightEval evaluation framework [54]. Notably, under this framework, we are able to achieve scores above random (25%) for 0-shot MMLU for 1B models by considering the log-probabilities of entire answer passages as opposed to single letters. The 1B Hugging Face model trained on FineWeb-Edu [97] has shown to work well on this, so we wanted to more closely examine how LightEval evaluation scores correlate with evaluation scores from LLM Foundry. We present our findings in Figure˜5.
Figure 5: Comparisons Between LightEval MMLU scores (x-axis) and LLM Foundry MMLU scores (y-axis). LightEval is able to provide signal (i.e. score above random baseline) earlier for weaker models, but the LightEval scores at larger scales appear to be capped at a much lower threshold and are more closely clumped together.
The key difference between LightEval and LLM Foundry for multiple choice tasks like MMLU is that LightEval considers the log probabilities of entire answer sequences, whereas LLM Foundry only considers log probabilities of single letters. Nonetheless, Figure˜5 shows a positive correlation between the two evaluation frameworks on MMLU 0-shot accuracy.
In Figure˜5 we were able to reproduce the MMLU scores reported in the FineWeb-Edu blog [97]. Notably, we found that LightEval indeed gave MMLU scores above random for the 1B scales, whereas in LLM Foundry, all the 1B models have accuracies around 0.25. At larger scales, however, the LightEval scores for the models become quite cramped together, which may make it more difficult to compare models and may make the comparisons more susceptible to noise. For example, the models Gemma-7B, Llama3-8B, and Mistral-7B all have scores between 0.43 and 0.44 in LightEval, while their scores range from 0.56 to 0.62 for LLM Foundry. We also see that FineWeb-Edu 7B-2x and DCLM 7B-2x perform quite similarly in LightEval, but DCLM-7B is better by close to 10 points in LLM Foundry. In conclusion, we believe that LightEval can be potentially a good choice when evaluating smaller models, but other frameworks like LLM Foundry could give clearer signals when comparing larger models.
One limitation of this study is that we took MMLU as a representative task, and we did not evaluate on other tasks. In the future, it would be interesting to compare with additional tasks, as well as additional frameworks like Eleuther LLM Harness.
Appendix H Hyperparameter study
A potential concern is that the training recipe can change conclusions about which dataset is optimal for training, due to interaction between training hyperparameters and dataset distributions. To address this confounder, we show in Table˜13 that orderings between datasets are preserved for various combinations of weight decay and learning rate. Moreover we find that performance gains from optimal hyper-parameter choice and dataset design tend to be orthogonal and complement each other. We illustrate this effect in Table˜14.
Table 13: Rankings are stable across hyperparameters (1B-1x scale). We train models on 3 datasets with 5 hyperparameter settings, varying learning rate and weight decay settings. Across the hyperparameter settings, the dataset ranking remains largely stable, with DCLM-baseline outperforming RedPajama, which in turns outperforms C4. With improved hyperparameters, the gaps between the datasets grows: e.g., at ‘Default’ (the best hyperparameter setting), DCLM-baseline outperforms RedPajama by 4.5 points and RedPajama outperforms C4 by 2 points, while at ‘0.1x Learning Rate’ (the lowest performing setting), the gaps reduce to 3.3 points and 0.9 points respectively. Note: When changing learning rate, we also update weight decay so the product of the two remains the same.
Hyperparameters Dataset Learning Rate (LR) Weight Decay (WD) Core Extended
0.1x Learning Rate C4 3.0e-04 3.3e-01 22.1 11.6
RedPajama 3.0e-04 3.3e-01 23.0 11.8
DCLM-baseline 3.0e-04 3.3e-01 26.3 14.4
0.1x Weight Decay C4 3.0e-03 3.3e-03 21.8 11.6
RedPajama 3.0e-03 3.3e-03 23.0 11.8
DCLM-baseline 3.0e-03 3.3e-03 28.3 15.1
Default C4 3.0e-03 3.3e-02 23.7 12.5
RedPajama 3.0e-03 3.3e-02 25.7 13.5
DCLM-baseline 3.0e-03 3.3e-02 30.2 15.4
10x Weight Decay C4 3.0e-03 3.3e-01 21.8 12.0
RedPajama 3.0e-03 3.3e-01 22.5 11.9
DCLM-baseline 3.0e-03 3.3e-01 27.1 14.0
10x Learning Rate C4 3.0e-02 3.3e-03 22.7 12.3
RedPajama 3.0e-02 3.3e-03 26.0 13.2
DCLM-baseline 3.0e-02 3.3e-03 29.0 15.0
Table 14: Improvements from better hyperparameters stack with better datasets. (7B-1x scale). We evaluate the impact of the most influential step in our dataset design, model based filtering (‘fastText filtering’), stacked with a better hyperparameter setting. We see that for both MMLU and Core benchmarks, the two inteventions (better dataset and better hyperparmaeters) seem to be orthogonal and stack on top of each other.
Hyperparameters fastText filtering LR WD MMLU Core
Low LR 3.0e-04 0.33 25.4 38.3
3.0e-04 0.33 29.2 41.0
High LR 1.0e-03 0.1 25.5 39.7
1.0e-03 0.1 38.3 44.1
Appendix I Model-based quality filters
We presented results for several different model-based quality filters in Section˜4.4. In this section, we describe their implementations in further detail, focusing especially on fastText classifiers which were our method of choice for DCLM-baseline.
fastText classifiers
Training
We use the supervised fastText package from Joulin et al. [79] to train models to classify between chosen “high-quality” reference data which are given positive labels, and web-crawled data which are given negative labels. We then apply these classifiers to score each document from the pool we wish to filter, taking the predicted probability of the positive label as the score and computing a percentile-based threshold. In terms of training hyperparameters, we mostly used the default choices from the fastText package; the only hyperparameter change that we tried was to expand the feature space from unigrams only to both unigrams and bigrams (via setting the wordNgrams argument to be 2 instead of the default 1). This helped improve the quality of downstream filtered datasets, as shown in Table˜15 (which extends Table˜5 from Section˜4.4).
Table 15: fastText feature-space ablation (7B-1x scale). Adding bigrams to the feature space helps over the default setting of unigrams only.
Dataset Threshold Features Core MMLU Extended
OH-2.5 + ELI5 10% Unigrams + Bigrams 41.0 29.2 21.4
OH-2.5 + ELI5 10% Unigrams 40.0 28.3 22.1
Data preparation.
The bulk of our experimentation for training fastText models focused on constructing their underlying training sets, specifically the positively labeled reference data. For each experiment, we fixed the size of the training set to be 400K examples (i.e., 200K positive, 200K negative). The negatively labeled examples were sampled randomly from a set of documents that came from an earlier (smaller) version of our RefinedWeb reproduction. This version used trafilatura as the extractor instead of resiliparse, which we hypothesize might actually help for training the filtering model; as shown in Appendix˜J, trafilatura more aggressively removes boilerplate content that may appear in many pages (especiall from the same website). This type of content, if left in, may lead to the fastText models over-relying on these “spurious” features instead of the main contents of the page. For the positively labeled reference data, we tried several different sources, some of which involved further pre-processing:
  • Wikipedia. We use the processed version from RedPajama [156] and apply English filtering by only keeping pages from the en.wikipedia.org domain. To encourage the classifier to rely on the core content of each page, we remove occurrences of the section titles "See Also" and "References", at least one of which occurs in 90% of articles.
  • OpenWebText2. We use this dataset as is, taken from the version in The Pile [57].
  • GPT-3 Approx. We mix together Wikipedia and OpenWebText2 along with the books source from RedPajama [156]. Given the long length of individual books, we instead define examples by extracting chunks of text that are at most 2048 tokens long.
  • OH-2.5 + ELI5. Our goal for this mix was to source instruction and question-answer formatted data that is both high-quality and covers a wide range of potential topics. We sample 100K examples from OH-2.5, which we do not further pre-process. For ELI5, each raw page from the r/ExplainLikeImFive subreddit contains a post asking a specific question and then some number of comments aiming to answer said question. We curate examples for training fastText models by taking a post and combining it with the top-scoring answer (using the karma score derived from community up/down-votes). If there are ties, the longest answer is chosen. We also filter these examples by keeping only those where the post has score $\geq$ 3, the best comment has score $\geq$ 5, and there are at least 3 comments total.
Other quality filtering baselines
We also examined other quality filters, though found none as effective as the fastText methods described above, as shown in Table˜4. We now provide further details for some of these baselines.
PageRank.
An intuitively promising, but ultimately unfruitful approach was to consider page centrality metrics such as PageRank and Harmonic centrality metrics, with the idea that more "central" web text would yield higher quality data. We collected PageRank metrics from Common Crawl’s host level webgraph dataset [^3] https://commoncrawl.org/web-graphs and omitted any hosts that did not appear in the crawl. Next we partitioned our RefinedWeb reproduction into quintiles based on their PageRank score and trained several models at the 1B-1x scale. These results are collated in Table˜16, but unfortunately no quintile performed better than a pool sampled from the union of all quintiles.
Table 16: PageRank-based filtering (1B-1x scale). Using PageRank score to select data is not helpful for improving upon our RefinedWeb reproduction. Using any quintile based on this score performs worse than a random sample from the same initial pool.
Quintile All 1 2 3 4 5
Core 27.8 26.1 27.3 26.6 26.3 27.1
AskLLM.
A recent line of work studies using instruction-tuned models as annotators to determine the potential usefulness of a document. Sachdeva et al. [136] proposed AskLLM, in which the authors prompted Flan-T5 models [35] to evaluate whether the given document $\dots$ contain[s] informative signal for pre-training a large-language model? An informative datapoint should be well-formatted, contain some usable knowledge of the world, and strictly NOT have any harmful, racist, sexist, etc. content.”. We implemented this method ourselves, testing several models as annotators, different settings for maximal sequence length, and several prompts on a small scale. We found that the best configuration was using Mistral-7B-Instruct-v0.2 [75], clipping the document at 1024 tokens, and taking the cumulative probabilities of Yes and yes tokens as the model score. We used the following prompt template:
        Evaluate the following paragraph for LLM pre-training suitability:        - Is it well-structured or contains useful examples to natural texts?        - Does it offer insights, useful facts or relevant information?        {internallinenumbers*} - Does it teach how to comply with open-ended tasks such as writing letters, poems, emails etc.?        - Is it free from harmful content?              {internallinenumbers*} If most criteria are met based on the content below, indicate ’yes’ for suitable. Otherwise, indicate ’no’ for unsuitable.              ### <input> ###              Is it suitable for LLM pre-training? OPTIONS:        - yes        - no
where <input> is replaced with the document tokens, clipped if too long, and appended with ‘‘…[The rest of the paragraph is omitted]’’ in such cases. While this method worked slightly better than random sampling from our pool (see Table˜4), it significantly underperformed compared to our fastText experiments. Considering the high costs associated with applying it at scale, we did not perform this experiment on a larger scale.
Semantic deduplication.
Following the success of the different deduplication methods we used (Appendix˜K), we studied the effect of Semantic deduplication as proposed by Abbas et al. [1]. In this approach, the authors propose embedding the documents using pre-trained language models, clustering them using k-means, and removing all but one document from each group of closely related documents to encourage diversity in the dataset. We began by embedding each document in a pool of approximately 100 million documents (following Abbas et al. [1]’s best practices) with BGE-base [172]. We then used faiss-GPU [77] to perform spherical k-means clustering, with 20 iterations and $K=11000$. We sampled documents after discarding 25% of the data. As seen in Table˜4, this intervention only negatively impacted the trained model. We hypothesize that the model used for embedding has a significant impact on the outcomes of this method. However, due to the large computational overhead when scaled, making it infeasible, we opted to rely on the deduplication methods outlined in Appendix˜K and leave this line of research for future work.
Appendix J Text extraction comparison
Here, we share more detailed quantitative and qualitative comparisons between our chosen extractor, resiliparse [19], and the two alternatives previously used by other datasets: WET files, trafilatura [16].
Profiling
We compute basic summary statistics for each extractor based on a sample of 10 WARC files (corresponding to 900K individual pages), presenting the results in Table˜17. Notably, both resiliparse and trafilatura result at least 2x shorter documents on average compared to WET files. As shown in the examples in Section˜J.2, WET files indeed contain many additional lines with seemingly little value for pre-training (e.g. navigation bars, boilerplate notices, copyright statements). trafilatura and resiliparse trim most of these lines out, with the former being more strict about doing so. Between the two, resiliparse still keeps in about 10% more text; some of this additional text may provide useful content such as section titles and dates for articles. In terms of runtime, the two are much farther apart, with resiliparse being roughly 8x faster.
Table 17: Text extractor profiling. Characters and tokens are averaged over the number of resulting output pages (note that this may differ for each extractor due to due to the possibility of extraction failures). Throughput is measured in GBs of input WARCs processed per second for each CPU core.
Extractor Avg. Chars Avg. Tokens Throughput (GB / sec / core)
WET 6,580 2,824
resiliparse 3,227 1,329 $4.55\times 10^{-3}$
trafilatura 2,901 1,179 $0.56\times 10^{-3}$
Extraction examples
J.2.1 Example set 1
        [Trafilatura]              {internallinenumbers*} HERE is a sampling of some of the better antiques and flea markets around the United States.        Two or Three Times a Year        {internallinenumbers*} BRIMFIELD Route 20, Brimfield, Mass. 01010; 413-245-3436. Second weekend of May and July, and the second weekend after Labor Day.        {internallinenumbers*} RENNINGER’S OUTDOOR EXTRAVAGANZA Noble Street, Kutztown, Pa.; 717-385-0104. Thursday, Friday and Saturday of the last weekend of April, June, September.        {internallinenumbers*} FARMINGTON ANTIQUES WEEKEND Farmington Polo Grounds, Town Farm Road, Farmington, Conn. 06032; 508-839-9735. Starting Wednesday before shows open; 203-677-7862. June 9-10 and Sept. 1-2.        Monthly        {internallinenumbers*} ANN ARBOR ANTIQUES MARKET, P.O. Box 1512, Ann Arbor, Mich. 48106; 313-662-9453. May through October, third Sunday.        {internallinenumbers*} Continue reading the main storyKANE COUNTY FLEA MARKET, Kane County Fairgrounds, P.O. Box 549, St. Charles, Ill. 60174; 708-377-2252. Year-round, first weekend.        {internallinenumbers*} THE METROLINA EXPO, 7100 Statesville Road, Charlotte, N.C. 28213; 704-596-4643. Year-round, first weekend of every month.        {internallinenumbers*} SPRINGFIELD ANTIQUE SHOW AND FLEA MARKET, Clark County Fairgrounds, Route 41, Springfield, Ohio, 45501; 513-325-0053. Year-round, third weekend.        Weekly        {internallinenumbers*} BAKERSFIELD SWAP-O-RAMA, 4501 Wible Road, Bakersfield, Calif. 93313; 805-831-9342. Saturday and Sunday.        {internallinenumbers*} LAMBERTVILLE ANTIQUE MARKET, Route 29, Lambertville, N.J. 08530. Weekend number: 609-397-0456. Weekday: 215-752-4485, between 5 and 7 P.M. Market on Saturday and Sunday.        {internallinenumbers*} ATLANTA FLEA MARKET AND ANTIQUE CENTER, 5360 Peachtree Industrial Boulevard, Chamblee, Ga. 30341; 404-458-0456. Friday, Saturday and Sunday.        Continue reading the main story
        [Resiliparse]              {internallinenumbers*} This is a digitized version of an article from The Times’s print archive, before the start of online publication in 1996. To preserve these articles as they originally appeared, The Times does not alter, edit or update them.              {internallinenumbers*} Occasionally the digitization process introduces transcription errors or other problems. Please send reports of such problems to archive_feedback@nytimes.com.              May 10, 1990, Page 00006 The New York Times Archives              {internallinenumbers*} HERE is a sampling of some of the better antiques and flea markets around the United States.              Two or Three Times a Year              {internallinenumbers*} BRIMFIELD Route 20, Brimfield, Mass. 01010; 413-245-3436. Second weekend of May and July, and the second weekend after Labor Day.              {internallinenumbers*} RENNINGER’S OUTDOOR EXTRAVAGANZA Noble Street, Kutztown, Pa.; 717-385-0104. Thursday, Friday and Saturday of the last weekend of April, June, September.              {internallinenumbers*} FARMINGTON ANTIQUES WEEKEND Farmington Polo Grounds, Town Farm Road, Farmington, Conn. 06032; 508-839-9735. Starting Wednesday before shows open; 203-677-7862. June 9-10 and Sept. 1-2.              Monthly              {internallinenumbers*} ANN ARBOR ANTIQUES MARKET, P.O. Box 1512, Ann Arbor, Mich. 48106; 313-662-9453. May through October, third Sunday.              Continue reading the main story              {internallinenumbers*} KANE COUNTY FLEA MARKET, Kane County Fairgrounds, P.O. Box 549, St. Charles, Ill. 60174; 708-377-2252. Year-round, first weekend.              {internallinenumbers*} THE METROLINA EXPO, 7100 Statesville Road, Charlotte, N.C. 28213; 704-596-4643. Year-round, first weekend of every month.              {internallinenumbers*} SPRINGFIELD ANTIQUE SHOW AND FLEA MARKET, Clark County Fairgrounds, Route 41, Springfield, Ohio, 45501; 513-325-0053. Year-round, third weekend.              Weekly              {internallinenumbers*} BAKERSFIELD SWAP-O-RAMA, 4501 Wible Road, Bakersfield, Calif. 93313; 805-831-9342. Saturday and Sunday.              {internallinenumbers*} LAMBERTVILLE ANTIQUE MARKET, Route 29, Lambertville, N.J. 08530. Weekend number: 609-397-0456. Weekday: 215-752-4485, between 5 and 7 P.M. Market on Saturday and Sunday.              {internallinenumbers*} ATLANTA FLEA MARKET AND ANTIQUE CENTER, 5360 Peachtree Industrial Boulevard, Chamblee, Ga. 30341; 404-458-0456. Friday, Saturday and Sunday.              Continue reading the main story
        [WET file]              A Guide To Markets - The New York Times        {internallinenumbers*} NYTimes.com no longer supports Internet Explorer 9 or earlier. Please upgrade your browser. LEARN MORE »        Sections        Home        Search        Skip to content Skip to navigation View mobile version        The New York Times        Archives|A Guide To Markets        Search        Subscribe Now        Log In        0        Settings        Close search        Site Search Navigation        Search NYTimes.com        Clear this text input        Go        https://nyti.ms/29nVV3Q        Loading...        See next articles        See previous articles        Site Navigation        Site Mobile Navigation        Advertisement        Archives | 1990        A Guide To Markets        MAY 10, 1990        Continue reading the main story Share This Page        Continue reading the main story        About the Archive        {internallinenumbers*} This is a digitized version of an article from The Times’s print archive, before the start of online publication in 1996. To preserve these articles as they originally appeared, The Times does not alter, edit or update them.        {internallinenumbers*} Occasionally the digitization process introduces transcription errors or other problems. Please send reports of such problems to archive_feedback@nytimes.com.        May 10, 1990, Page 00006 The New York Times Archives        {internallinenumbers*} HERE is a sampling of some of the better antiques and flea markets around the United States.        Two or Three Times a Year        {internallinenumbers*} BRIMFIELD Route 20, Brimfield, Mass. 01010; 413-245-3436. Second weekend of May and July, and the second weekend after Labor Day.        {internallinenumbers*} RENNINGER’S OUTDOOR EXTRAVAGANZA Noble Street, Kutztown, Pa.; 717-385-0104. Thursday, Friday and Saturday of the last weekend of April, June, September.        {internallinenumbers*} FARMINGTON ANTIQUES WEEKEND Farmington Polo Grounds, Town Farm Road, Farmington, Conn. 06032; 508-839-9735. Starting Wednesday before shows open; 203-677-7862. June 9-10 and Sept. 1-2.        Advertisement        Continue reading the main story        Monthly        {internallinenumbers*} ANN ARBOR ANTIQUES MARKET, P.O. Box 1512, Ann Arbor, Mich. 48106; 313-662-9453. May through October, third Sunday.        Continue reading the main story        Advertisement        Continue reading the main story        {internallinenumbers*} KANE COUNTY FLEA MARKET, Kane County Fairgrounds, P.O. Box 549, St. Charles, Ill. 60174; 708-377-2252. Year-round, first weekend.        {internallinenumbers*} THE METROLINA EXPO, 7100 Statesville Road, Charlotte, N.C. 28213; 704-596-4643. Year-round, first weekend of every month.        {internallinenumbers*} SPRINGFIELD ANTIQUE SHOW AND FLEA MARKET, Clark County Fairgrounds, Route 41, Springfield, Ohio, 45501; 513-325-0053. Year-round, third weekend.        Weekly        {internallinenumbers*} BAKERSFIELD SWAP-O-RAMA, 4501 Wible Road, Bakersfield, Calif. 93313; 805-831-9342. Saturday and Sunday.        {internallinenumbers*} LAMBERTVILLE ANTIQUE MARKET, Route 29, Lambertville, N.J. 08530. Weekend number: 609-397-0456. Weekday: 215-752-4485, between 5 and 7 P.M. Market on Saturday and Sunday.        {internallinenumbers*} ATLANTA FLEA MARKET AND ANTIQUE CENTER, 5360 Peachtree Industrial Boulevard, Chamblee, Ga. 30341; 404-458-0456. Friday, Saturday and Sunday.        {internallinenumbers*} A version of this list appears in print on May 10, 1990, on Page C00006 of the National edition with the headline: A Guide To Markets. Order Reprints| Today’s Paper|Subscribe        Continue reading the main story        What’s Next        Loading...        Go to Home Page »        Site Index The New York Times        Site Index Navigation        News        World        U.S.        Politics        N.Y.        Business        Tech        Science        Health        Sports        Education        Obituaries        Today’s Paper        Corrections        Opinion        Today’s Opinion        Op-Ed Columnists        Editorials        Op-Ed Contributors        Letters        Sunday Review        Video: Opinion        Arts        Today’s Arts        Art & Design        Books        Dance        Movies        Music        N.Y.C. Events Guide        Television        Theater        Video: Arts        Living        Automobiles        Crossword        Food        Education        Fashion & Style        Health        Jobs        Magazine        N.Y.C. Events Guide        Real Estate        T Magazine        Travel        Weddings & Celebrations        Listings & More        Reader Center        Classifieds        Tools & Services        N.Y.C. Events Guide        Multimedia        Photography        Video        NYT Store        Times Journeys        Subscribe        Manage My Account        NYTCo        Subscribe        Subscribe        Home Delivery        Digital Subscriptions        Crossword        Email Newsletters        Gift Subscriptions        Group Subscriptions        Education Rate        Mobile Applications        Replica Edition        Site Information Navigation        © 2019 The New York Times Company        Home        Search        {internallinenumbers*} Accessibility concerns? Email us at accessibility@nytimes.com. We would love to hear from you.        Contact Us        Work With Us        Advertise        Your Ad Choices        Privacy        Terms of Service        Terms of Sale        Site Information Navigation        Site Map        Help        Site Feedback        Subscriptions
J.2.2 Example set 2
        [Trafilatura]              Possible Duplicate:        When should I use an em-dash, an en-dash, and a hyphen?        When do I put a - in a sentence? Is it a more powerful comma? With a bigger pause?        Possible Duplicate:        When should I use an em-dash, an en-dash, and a hyphen?        When do I put a - in a sentence? Is it a more powerful comma? With a bigger pause?        {internallinenumbers*} This question has been asked before and already has an answer. If those answers do not fully address your question, please ask a new question.        {internallinenumbers*} The dashes you described are known respectively as the en-dash and the em-dash. To describe the difference between their origins, Mental Floss writes:        {internallinenumbers*} An en dash (–) is bigger than a hyphen but shorter than an em dash (—). Th e names come from an obscure typographical measurement system, but the dashes have now taken on a life of their own in grammar. The em dash is the spork of English grammar: It ain’t particularly pretty, but you can use it for most anything. Em dashes can replace colons or sets of parentheses, or represent a sudden change in thought or tone.        So when do you use an en-dash? Again from Mental Floss:        {internallinenumbers*} To show numerical ranges, signifying “up to and including”—of dates, ages, pages, etc. (Example: “I read pages 7–22 last night.”)        {internallinenumbers*} The storied “compound adjective hyphen,” an event so rare in the English language that proofreaders shiver with excitement whenever they come across it. Basically “pro-American” gets a regular hyphen because “American” is only one word, whereas “pro–Falkland Islands” gets an en dash because “Falkland Islands” is two words. So, too phrases like “Civil War–era.”        What about an em-dash? From here:        {internallinenumbers*} Similar to an extended hyphen (-), an em dash is used to show a break in thought or a shift of tone.        {internallinenumbers*} If you’d like to read more about the differences between a hyphen (-), en-dash (–), and em-dash (—), see the blog post here which summarizes the above.
        [Resiliparse]              1              Possible Duplicate:        When should I use an em-dash, an en-dash, and a hyphen?              When do I put a - in a sentence? Is it a more powerful comma? With a bigger pause?              {internallinenumbers*} marked as duplicate by waiwai933, MrHen, user2683, Robusto, Thursagen Jul 13 ’11 at 0:32              {internallinenumbers*} This question has been asked before and already has an answer. If those answers do not fully address your question, please ask a new question.              0              {internallinenumbers*} The dashes you described are known respectively as the en-dash and the em-dash. To describe the difference between their origins, Mental Floss writes:              {internallinenumbers*} An en dash (–) is bigger than a hyphen but shorter than an em dash (—). Th e names come from an obscure typographical measurement system, but the dashes have now taken on a life of their own in grammar. The em dash is the spork of English grammar: It ain’t particularly pretty, but you can use it for most anything. Em dashes can replace colons or sets of parentheses, or represent a sudden change in thought or tone.              So when do you use an en-dash? Again from Mental Floss:              {internallinenumbers*} 1. To show numerical ranges, signifying “up to and including”—of dates, ages, pages, etc. (Example: “I read pages 7–22 last night.”)              {internallinenumbers*} 2. The storied “compound adjective hyphen,” an event so rare in the English language that proofreaders shiver with excitement whenever they come across it. Basically “pro-American” gets a regular hyphen because “American” is only one word, whereas “pro–Falkland Islands” gets an en dash because “Falkland Islands” is two words. So, too phrases like “Civil War–era.”              What about an em-dash? From here:              {internallinenumbers*} Similar to an extended hyphen (-), an em dash is used to show a break in thought or a shift of tone.              {internallinenumbers*} If you’d like to read more about the differences between a hyphen (-), en-dash (–), and em-dash (—), see the blog post here which summarizes the above.              {internallinenumbers*} Not the answer you’re looking for? Browse other questions tagged or ask your own question.
        [WET file]              {internallinenumbers*} syntax - What’s the difference between - and -- in a phrase? - English Language & Usage Stack Exchange        Stack Exchange Network        {internallinenumbers*} Stack Exchange network consists of 175 Q&A communities including Stack Overflow, the largest, most trusted online community for developers to learn, share their knowledge, and build their careers.        Visit Stack Exchange        Log In Sign Up        current community        English Language & Usage        help chat        English Language & Usage Meta        your communities        Sign up or log in to customize your list.        more stack exchange communities        company blog        Tour Start here for a quick overview of the site        Help Center Detailed answers to any questions you might have        Meta Discuss the workings and policies of this site        About Us Learn more about Stack Overflow the company        Business Learn more about hiring developers or posting ads with us        {internallinenumbers*} By using our site, you acknowledge that you have read and understand our Cookie Policy, Privacy Policy, and our Terms of Service.        {internallinenumbers*} English Language & Usage Stack Exchange is a question and answer site for linguists, etymologists, and serious English language enthusiasts. Join them; it only takes a minute:        Sign up        Here’s how it works:        Anybody can ask a question        Anybody can answer        The best answers are voted up and rise to the top        Home        Questions        Tags        Users        Unanswered        What’s the difference between - and — in a phrase? [duplicate]        Ask Question        1        Possible Duplicate:        When should I use an em-dash, an en-dash, and a hyphen?        When do I put a - in a sentence? Is it a more powerful comma? With a bigger pause?        syntax dashes symbols        share|improve this question        edited Apr 13 ’17 at 12:38        Community        1        asked Jul 13 ’11 at 0:10        curiouscurious        123115        {internallinenumbers*} marked as duplicate by waiwai933, MrHen, user2683, Robusto, Thursagen Jul 13 ’11 at 0:32        {internallinenumbers*} This question has been asked before and already has an answer. If those answers do not fully address your question, please ask a new question.        add a comment |        1 Answer 1        active oldest votes        0        {internallinenumbers*} The dashes you described are known respectively as the en-dash and the em-dash. To describe the difference between their origins, Mental Floss writes:        {internallinenumbers*} An en dash (–) is bigger than a hyphen but shorter than an em dash (—). Th e names come from an obscure typographical measurement system, but the dashes have now taken on a life of their own in grammar. The em dash is the spork of English grammar: It ain’t particularly pretty, but you can use it for most anything. Em dashes can replace colons or sets of parentheses, or represent a sudden change in thought or tone.        So when do you use an en-dash? Again from Mental Floss:        {internallinenumbers*} To show numerical ranges, signifying “up to and including”—of dates, ages, pages, etc. (Example: “I read pages 7–22 last night.”)        {internallinenumbers*} The storied “compound adjective hyphen,” an event so rare in the English language that proofreaders shiver with excitement whenever they come across it. Basically “pro-American” gets a regular hyphen because “American” is only one word, whereas “pro–Falkland Islands” gets an en dash because “Falkland Islands” is two words. So, too phrases like “Civil War–era.”        What about an em-dash? From here:        {internallinenumbers*} Similar to an extended hyphen (-), an em dash is used to show a break in thought or a shift of tone.        {internallinenumbers*} If you’d like to read more about the differences between a hyphen (-), en-dash (–), and em-dash (—), see the blog post here which summarizes the above.        share|improve this answer        edited Aug 4 ’15 at 17:00        zwol        2,51911424        answered Jul 13 ’11 at 0:16        simchonasimchona        30.9k5112139        add a comment |        {internallinenumbers*} Not the answer you’re looking for? Browse other questions tagged syntax dashes symbols or ask your own question.        asked        7 years, 9 months ago        viewed        2,336 times        active        3 years, 8 months ago        Featured on Meta        Announcing the arrival of Valued Associate #679: Cesar Manara        Planned maintenance scheduled April 17/18, 2019 at 00:00UTC (8:00pm US/Eastern)        Linked        279        When should I use an em-dash, an en-dash, and a hyphen?        Related        279        When should I use an em-dash, an en-dash, and a hyphen?        4        When to use -, – and —?        6        What is the difference between `-` and `--`        5        Does “cost-benefit ratio” use a hyphen or an en-dash?        0        {internallinenumbers*} What kind of dash character should I use at the end of a famous saying to mark of the author?        -1        dash non-restrictive element in the middle of a sentence        1        em dash followed by a comma        0        what’s the difference between a hyphen, a dash and a minus sign?        0        Using comma to delimit the name of a group and its constituents?        1        I tend to overuse the hyphen as a pause, and would appreciate some feedback on this        Hot Network Questions        How could we fake a moon landing now?        Using audio cues to encourage good posture        Chinese Seal on silk painting - what does it mean?        How does the math work when buying airline miles?        What causes the direction of lightning flashes?        Maximum summed subsequences with non-adjacent items        Most bit efficient text communication method?        Significance of Cersei’s obsession with elephants?        Is it a good idea to use CNN to classify 1D signal?        What font is "z" in "z-score"?        How can I use the Python library networkx from Mathematica?        Fundamental Solution of the Pell Equation        Dating a Former Employee        Do I really need recursive chmod to restrict access to a folder?        Generate an RGB colour grid        How to Make a Beautiful Stacked 3D Plot        How to find all the available tools in mac terminal?        {internallinenumbers*} Has negative voting ever been officially implemented in elections, or seriously proposed, or even studied?        What is the meaning of the simile “quick as silk”?        Circuit to "zoom in" on mV fluctuations of a DC signal?        {internallinenumbers*} What are the out-of-universe reasons for the references to Toby Maguire-era Spider-Man in Into the Spider-Verse?        What does this Jacques Hadamard quote mean?        Can a new player join a group only when a new campaign starts?        Why wasn’t DOSKEY integrated with COMMAND.COM?        more hot questions        English Language & Usage        Tour        Help        Chat        Contact        Feedback        Mobile        Company        Stack Overflow        Stack Overflow Business        Developer Jobs        About        Press        Legal        Privacy Policy        Stack Exchange        Network        Technology        Life / Arts        Culture / Recreation        Science        Other        Stack Overflow        Server Fault        Super User        Web Applications        Ask Ubuntu        Webmasters        Game Development        TeX - LaTeX        Software Engineering        Unix & Linux        Ask Different (Apple)        WordPress Development        Geographic Information Systems        Electrical Engineering        Android Enthusiasts        Information Security        Database Administrators        Drupal Answers        SharePoint        User Experience        Mathematica        Salesforce        ExpressionEngine® Answers        Stack Overflow em Português        Blender        Network Engineering        Cryptography        Code Review        Magento        Software Recommendations        Signal Processing        Emacs        Raspberry Pi        Stack Overflow Ha pycckom        Programming Puzzles & Code Golf        Stack Overflow en español        Ethereum        Data Science        Arduino        Bitcoin        more (31)        Photography        Science Fiction & Fantasy        Graphic Design        Movies & TV        Music: Practice & Theory        Worldbuilding        Seasoned Advice (cooking)        Home Improvement        Personal Finance & Money        Academia        Law        more (15)        English Language & Usage        Skeptics        Mi Yodeya (Judaism)        Travel        Christianity        English Language Learners        Japanese Language        Arqade (gaming)        Bicycles        Role-playing Games        Anime & Manga        Puzzling        Motor Vehicle Maintenance & Repair        more (33)        MathOverflow        Mathematics        Cross Validated (stats)        Theoretical Computer Science        Physics        Chemistry        Biology        Computer Science        Philosophy        more (10)        Meta Stack Exchange        Stack Apps        API        Data        Blog        Facebook        Twitter        LinkedIn        {internallinenumbers*} site design / logo © 2019 Stack Exchange Inc; user contributions licensed under cc by-sa 3.0 with attribution required. rev 2019.4.18.33353        English Language & Usage Stack Exchange works best with JavaScript enabled
Appendix K Deduplication
We perform extensive ablations and experimentation on various deduplication pipelines. This section is organized by first describing the deduplication methods considered and then outlining the ablations that lead us to the choice of deduplication pipeline used in generating DCLM-baseline (and other DCLM scales).
Deduplication methods
Prior work such as Lee et al. [86], Penedo et al. [118] use a two-stage deduplication pipeline where near duplicates are first removed at a inter-document level by identifying and removing near-duplicates using the MinHash algorithm, and then at an intra-document level where any substring of a predetermined length that occurs more than once in the entire corpus is removed. Intuitively, this strategy makes sense as the notion of a "duplicate" is poorly defined and can include documents such as: (i) exact copies of entire documents (targeted at the document-level); (ii) documents where the majority of the text is a duplicate, but there are unique differences in just the header or footer (targeted at the document-level); or (iii) documents where there are significant sections of unique text, but also massively repeated boilerplate text (targeted at the intra-document level). Performing multiple resolutions of deduplication can target all such cases, and further, a deduplication pipeline that can target near-duplicates, often referred to as "fuzzy deduplication" can identify documents that humans would intuitively refer to as duplicates.
While we ultimately rely on a Bloom filter based method of deduplication for our datasets, we describe the other pipelines considered:
MinHash.
MinHash is a locality-sensitive hashing technique used to group sets into collections based on their Jaccard similarity [26]. In the context of deduplicating text datasets, MinHash was first employed in Lee et al. [86] and then used in numerous other projects [118, 41]. We point readers to the main text of Lee et al. [86] and Appendix G.3.1 of Penedo et al. [118] for more details. The primary hyperparameters of note are the ngram-size, and the number of permutations used. Following Lee et al. [86], Penedo et al. [118], we use an ngram-size of 5 tokens and target a Jaccard similarity of 0.8. Departing from prior work, however, we modify the number of MinHash permutations used. Both Lee et al. [86] and Penedo et al. [118] use a total of 9,000 permutations, split into 450 buckets of 20 hashes each. We found this to be overly expensive and notice that similar Jaccard similarity plots can be attained with a much smaller number of permutations. For all of our ablations, we instead use a total of 1,395 permutations, split into 93 buckets of size 15. These hyperparameters were chose programmatically to mimic the Jaccard similarity plots as closely as possible, in an $\ell_{2}$ sense, with a fixed hash budget. See Figure˜6 for more details.
Figure 6: Probability of two documents with Jaccard similarity (x-axis) being marked as duplicates (y-axis) with varying (number of buckets, bucket size) parameters. (450, 20) corresponds to Lee et al. [86], Penedo et al. [118], our experiments used (93, 15), chosen to be a cheaper alternative emulating the same performance as (450, 20). The parameters (14,9) were used by Penedo et al. [119].
Suffix arrays.
Suffix arrays, first introduced in Manber & Myers [102], enable efficient identification and removal of substrings of a large corpus of text. This is done by first concatenating all text in the corpus together and then sorting each suffix. By scanning this sorted list, substrings with a common prefix can by identified by scanning the prefices of neighboring elements in the sorted list. This latter step can be done in an embarassingly parallel fashion, but the implementation we employed, borrowed from the codebase provided in Lee et al. [86] is not done in a multi-node fashion and requires loading the entire corpus into RAM. We directly employ the hyperparameters used in Lee et al. [86] and remove all repeated substrings that are at least 50 tokens long.
Bloom filters.
Bloom filters are a data structure that enable space-efficient set membership queries [24]. Explicitly, in sublinear space, a Bloom filter maintains a sketch of a set, that supports an insert operation, and a probabilistic membership_query operation, where the latter will never return any false negatives (i.e., return False for an element in the set), but will occasionally return a false positive (i.e., return True for an element not in the set). These were first used in the context of exact-duplicate removal in Soldaini et al. [147], but have since been extended to perform near-duplicate document and paragraph removal in a tool known as BFF (Big Friendly Filter) [65], and we further modify BFF to perform deduplication at the document and paragraph level simultaneously. We found that this technique is vastly more efficient than a MinhHash and SuffixArray pipeline. However there is one important caveat in that MinHash performs document-level deduplication at a document vs. document level, whereas BFF performs document-level deduplication at a document vs. corpus level.
Paragraph + document BFF.
Here we outline our modified Bloom filter based deduplication algorithm. Upon initialization, we require an estimate of the number of tokens in our entire corpus as well as a desired false-positive rate, to initialize a Bloom filter with a fixed size and number of hashers. The optimal number of hashers, $k$, is given by the formula
$k=-\frac{\ln{\epsilon}}{\ln{2}},$
where $\epsilon$ is the desired false positive rate. The optimal size $m$ for $k$ hashers and $n$ tokens can then be computed by solving for $m$ in the following formula:
$\epsilon=\Big(1-e^{\frac{-kn}{m}}\Big)^{k}.$
While this does not admit an easy analytical solution, it is trivial to solve for $m$ by using a binary search algorithm.
[…]
  • If the paragraph is fewer than min_ngram_size tokens long, it is left as is.
  • If the paragraph is in between min_ngram_size and max_ngram_size (inclusive) then total_ngrams is incremneted and this ngram’s membership is checked in the Bloom filter. If it is present in the Bloom filter, it is removed from the paragraph and contained_ngrams is incremented. Otherwise, it is added to the Bloom filter.
  • If the paragraph is at longer than max_ngram_size tokens, then each ngram of size max_ngram_size increments the total counter and is checked against the Bloom filter. If present, the contained_ngrams counter is incremented. If greater than threshold fraction of the ngrams in this paragraph are contained in the Bloom filter, then the entire paragraph is removed from the document. Otherwise, every non-contained ngram is added to the Bloom filter.
Once all paragraphs have been processed, if the ratio between the counters contained_ngrams and total_ngrams is greater than threshold, then the entire document is removed from the corpus.
To finalize our discussion on the Bloom filter based deduplication, we offer brief explanations on the hyperparameter choices made.
  • False Positive Rate: The two parameters that dictate the memory footprint required by BFF are the number of tokens and the false positive rate. However we only can control the false positive rate, and we notice that the Bloom filter size scales linearly with the negative log of the false positive rate. In particular, for a corpus of 1T tokens, occupying roughly 2TB of disk space, ensuring no false positives, i.e. setting the false positive rate to $1/1T$, would require 6.5TB of RAM. Here we argue analytically that a false positive rate of even as low as 0.01 suffices, which we support with experimentation in the next section.
    In choosing a false positive rate for the ngram-based Bloom filter, it’s important to recognize that removal of a paragraph or document is dictated by having greater than a threshold fraction of the ngrams contained in the set. As an example, suppose we are given a paragraph of $N$ ngrams, where $S$ of them are already contained in the Bloom filter and we set threshold to $T$. Because Bloom filters do not allow false negatives, every one of the $S$ ngrams are marked (correctly) as contained, and NSN-S of them could potentially be marked as a false positive. Indeed, of the NSN-S of these ngrams, at least TNSTN-S of them would need to be marked as a false positive, each of which occurs independently with probability $\epsilon$. This is equivalent to NSN-S Bernoulli random variables with parameter $\epsilon$, and can be bounded by a crude Hoeffding bound. In this particular case, the probability that a document or paragraph is falsely marked as a duplicate is bounded by:
    $\exp\Big(\frac{-2\cdot\big(TN-S-\epsilon\cdot(N-S)\big)^{2}}{N-S}\Big)$
    To put things concretely, in a document with 100 ngrams and a threshold of 0.8 and a false positive rate of 0.01, if 60 of the ngrams have been seen before, the probability of the document being marked as a duplicate is less than $10^{-8}$. Unless otherwise specified, we always use a false positive rate of 0.01.
  • min_ngram_size: In choosing a size for minimum ngrams, we recognize that many documents contain paragraphs that are itemized lists and are quite short; for example, recipes often include bullet-pointed ingredients lists, and MMLU multiple choice questions may often be quite short. While we originally noticed improved Core scores by setting a minimum size to 5 tokens, we noticed that this caused a worse performance on MMLU. After manual inspection, we settled on a min and max ngram size of 13 tokens.
  • threshold: Ablations did not show a noticable difference in deduplication performance.
Deduplication experiments
K.2.1 Deduplication ablations: pipeline at 1B-1x scale
We first perform ablations regarding the full pipeline choice for deduplication at the 1B-1x scale. We start with a pool of 76B tokens subsampled from Common Crawl with the preprocessing steps from Penedo et al. [118] applied. Then we apply a combination of deduplication steps, and subsample the pool further to the 28B tokens required for the 1B-1x scale. Finally we train and evaluate the Core score and the percentage of tokens that were removed by deduplication. The main questions we seek to answer from this round of ablations are:
  • For multi-step deduplication pipelines, how much of a contribution does each step provide?
  • Which deduplication pipeline is worth scaling up to larger pool sizes?
Results are contained in Table˜18. The main conclusions we can arrive at from this table are as follows: i) Suffix Array deduplication seems to help more than MinHash deduplication, thereby giving some signal to the source of the gains procured by a MinHash+SuffixArray pipeline; ii) BFF provides comparable performance to a full Exact+MinHash+SuffixArray pipeline, giving strong evidence that the multiresolution BFF could be an easily scalable alternative to the relatively more expensive MinHash+SuffixArray pipeline of prior works. Interestingly, it appears that a SuffixArray pipeline seems to outperform MinHash alone, though this falls within the range of variance for the Core score due to the nondeterminism in subsampling the dataset and training a model.
Table 18: Deduplication ablations (1B-1x scale). Starting from a pool of 76B tokens acquried from Common Crawl with the RefinedWeb Penedo et al. [118] pipeline applied, we evaluate the removal rate and Core score on different combinations of deduplication methods. Our Bloom filter method performs just as well as a combination of exact deduplication, MinHash and Suffix Array based techniques.
Exact Dedup MinHash Suffix Array Bloom Filter Tokens Removal Rate Core $\Delta$ from Baseline
76B 00% 40.1 +0.0
66B 13% 41.0 +0.9
62B 18% 40.9 +0.8
51B 33% 41.4 +1.3
56B 26% 41.7 +1.6
58B 24% 40.2 +0.1
49B 36% 41.3 +1.3
48B 37% 41.2 +1.2
45B 41% 41.7 +1.6
K.2.2 Deduplication ablations: pipeline at 7B-1x / 7B-2x scale
To further check the effects of BFF versus the more classical MinHash+SuffixArray we ran several experiments at the 7B-1x scale. Here we also introduce another hyperparameter, which we refer to as shards. By "sharding," we mean we break a dataset into chunks of roughly equal size and run the deduplication pipeline on each one of them independently. This is primarily done for engineering purposes, in that sharding is an easy way to further parallelize deduplication and convert single-node algorithms to multi-node algorithms. However, there are the side benefits of sharding for deduplication in that more shards yields a larger token pool: there are fewer documents to compare against and many documents which are repeated only a small number of times can survive such a process. Additionally there is some recent evidence that sharding seems to improve evaluation performance [119]. We also note that RefinedWeb [118] performs their deduplications on a 100-way sharding of the Common Crawl pool.
For this round of ablations, we start with a pool sourced from one tenth of Common Crawl and run the preprocessing steps from Penedo et al. [118] and apply various deduplication pipelines. Then we subsample down to 138B tokens and train and evaluate models at the 7B-1x scale. The main questions we seek to answer from this round of ablations are:
  • Is BFF still competitive with a MinHash+SuffixArray pipeline at larger scales?
  • Which BFF hyperparameters yield the highest Core and MMLU performance at this scale?
Results are contained in Table˜19. The first point to note is that BFF with a min_ngram_size at 13 and 20 yields Core scores and MMLU scores that are comparable to the scores attained by a MinHash+SuffixArray deduplicated pool at the same scale. The second point to note regards the BFF min_ngram_size and sharding: interestingly a lower min_ngram_size yields higher Core scores, but lower MMLU scores. We also see that fewer shards decreases the token yield, but has variable effect on the Core score. We examine the hyperparameters for BFF more fully in the next subsection.
Table 19: Deduplication Ablations (7B-1x scale). Starting with a pool from Common Crawl and the RW-Filter pipeline processing applied, we compared several BFF hyperparameters against the MinHash and Suffix Array pipeline of [86, 118]. Our best BFF run and the prior works are bolded.
Method min-ngram max-ngram Shards MMLU Core Token Yield
Bloom Filter 5 13 32 25.0 52.0 4T
Bloom Filter 5 13 1 27.1 52.7 1.3T
Bloom Filter 13 13 32 28.7 51.7 3.8T
Bloom Filter 20 20 32 27.7 51.5 4T
Bloom Filter 20 20 10 28.4 51.2 3T
MinHash+SA N/A N/A 16 29.1 51.9 3.2T
Encouraged by these results, next we examine the top candidates for a scalable deduplication pipeline at the 7B-2x scale. Again we start with a pool obtained from one tenth of Common Crawl and generate several deduplicated pools. The questions of interest are the same as above and we summarize the results in Table˜20. The key takeaways from this round of ablations is that at the 7B-2x scale, BFF with a min_ngram_size of 13 and 10 shards attains nearly identical performance to a MinHash+SuffixArray pipeline, whereas BFF with a min_ngram_size of 20 and 32 shards starts to lag behind, and that a min_ngram_size of 5 yields competitive Core scores, but falters in MMLU evaluations. While these experiments also vary the sharding choice, we view sharding primarily as a choice made to trade-off scalability with token yield. Larger shards are more expensive and less parallelizable and can decrease the token yield. For this round of ablations, the primary interest is to gain signal about how BFF compares to MinHash and Suffix Arrays at scale, and which are the correct hyperparameters for BFF. On this latter point, we chose to move forward with a min_ngram_size of 13 for generating DCLM-baseline.
Table 20: Deduplication Ablations (7B-2x scale). From the same pools as in Table˜19, we trained and evaluated models at the 7B-2x scale. Notice that a min_ngram_size of 5 yields competitive Core results but drastically reduces MMLU scores.
Method min_ngram max_ngram Shards MMLU Core Token Yield
BFF 20 20 10 43.6 55.7 3T
BFF 13 13 10 44.3 55.5 3T
BFF 5 13 32 32.0 54.2 4T
MinHash+SA N/A N/A 16 44.4 55.4 3.2T
K.2.3 BFF hyperparameter ablations
While the above ablations largely focused on the Core score and MMLU as performance metrics, these are expensive and not suited for large swaths of ablations. Here we instead explore statistics of datasets deduplicated by BFF as we toggle the ngram_size hyperparameters, false positive rate, and input dataset size. We run separate experiments for each hyperparameter and finish each paragraph with the choice of hyperparameter we use for all larger scale runs.
False positive rate
Table 21: False positive rate ablations. Starting with a pool of 75B tokens from the RW-Filter pipeline, we ran BFF with default hyperparameters, varying the false-positive rate to indicate that this does not have a large bearing on output pool size.
False Positive Rate Removal Rate (Bytes)
0.1 20.47%
0.01 20.47%
0.001 20.47%
Here we start with the 75B token data pool as in Section˜K.2.1 and focus on a paragraph-only level BFF. In other words, we run BFF as described above, except omit the full-document removal step. We use the default hyperparameters for ngram sizes as in Groeneveld [65], of 5 and 13 for min_ngram_size and max_ngram_size and a threshold of 0.8. We specifically look at the effect of changing the false positive rate and compute the removal rate (in bytes) of the output. From Table˜21, we can see that a false positive rate of 0.1 suffices for a reasonably small pool such as this one. For larger pools, to be safe, we always set the false positive rate to 0.01.
Min ngram size.
From Table˜19 and Table˜20, we saw that altering the ngram_size hyperparameters can affect both token yield and evaluation metrics. In particular, we seek to examine how surviving documents are altered by deduplication. As a proxy for this, we focus on the document lengths and removal rates. Results for this paragraph and the two following paragraphs are collated in Table˜22. One key observation is that as the min_ngram_size parameter is reduced, the mean and median document lengths become shorter. This indicates that too-low a min_ngram_size parameter can dramatically affect the language statistics of the dataset and should be avoided. This tracks with intuitive sense where many documents include linebreak separated lists where each list element is short and possibly repeated: e.g., many webpages include recipes that might call for "1 stick of butter", which would get removed with a min_ngram_size of 5 but would injuriously damage the source document.
Table 22: BFF hyperparameter ablations. Starting with a pool of 341B tokens taken from Common Crawl with the RW-Filter pipeline applied, we run our Bloom filter deduplication with various hyperparameters noting how the document length and pool size change after deduplication. The input pool statistics are noted in the first row.
Min Max Threshold Avg Tokens/Doc Median Tokens/Doc Total Documents Total Tokens
N/A N/A N/A 883 451 386M 341.0B
1 13 0.8 778 403 246M 191B
5 13 0.8 802 426 250M 201B
13 13 0.8 836 456 246M 205B
13 25 0.8 839 458 248M 208B
13 50 0.8 833 453 253M 211B
13 13 0.75 842 460 241M 203B
13 13 0.8 836 456 246M 205B
13 13 0.9 822 446 256M 211B
13 13 0.99 797 427 275M 218B
Max ngram size.
Next we examine increasing the max_ngram_size hyperparameter. Starting with the chosen min_ngram_size parameter of 13, decided in the previous paragraph, we consider max_ngram_size parameters of 13, 25, and 50. Contrary to the min_ngram_size, we do not see a dramatic alteration of language statistics as this parameter becomes increased. For simplicity, we choose to use a max_ngram_size of 13 for large-scale pools.
Threshold.
The threshold hyperparameter dictates how close a document must be to previously seen ngrams before it is considered a duplicate. We ablate this choice from 0.75 to 0.99, examining how this affects document length statistics and removal rates. Interestingly, as the threshold increases, documents get shorter, mirroring the statistics seen for reducing the min_ngram_size. As expected, higher thresholds yield lower removal rates. Following the Jaccard similarity choice used in MinHash deduplication and noting that 0.8 yields median tokens/doc closest to the baseline, we use a threshold of 0.8 going forward.
Shards.
Finally we simulate how shards affect the statistics of the deduplicated datasets. As above, the key statistics we focus on here are the removal rate and the average and medium document lengths. This is mostly to get a sense for how these features change as the dataset scales, with the prevailing thought that dramatically altering document statistics might adversely effect downstream evaluations. For larger pools, we can always shard them as heavily as desired, so we treat sharding as a hyperparameter that controls removal rate and document statistics. Results are collated in Table˜23 and Figure˜7. The key takeaways here are that removal rates increase monotonically with dataset size as expected, but do so in a concave fashion. This provides some signal for how heavily to shard an input pool if a desired token yield is specified. The next point of interest is to consider the document lengths as the dataset scales. These decrease monotonically as the pool increases in size.
Table 23: Deduplication shard size. We run a single-shard BFF with the ngram_size set to 13, false positive rate 0.01, threshold of 0.80 on pools of varying size. As the pool size scales the deduplication rate increases, documents get shorter and the removal rate increases.
Input Tokens Input Documents Avg Tokens/Doc Median Tokens/Doc Token Removal Rate
114B 129M 466 866 29%
227B 257M 460 848 35%
341B 386M 456 836 40%
455B 516M 453 826 43%
569B 643M 450 918 46%
In building DCLM-baseline, at the point of deduplication, the dataset is approximately 70TB in size. Since Table˜20 shows that BFF had the best performance at a 10-way shard with a roughly 7TB input size, we adhere to a 100-way sharding for DCLM-baseline, where each shard is roughly 700GB in size.
Figure 7: Deduplication shard size. We run a single-shard BFF with the ngram_size set to 13, false positive rate 0.01, threshold of 0.80 on pools of varying size. Larger pools have a larger removal rate, but this scales in a concave fashion. The removal rates for tokens and documents begin to diverge at larger scales.
K.2.4 Global MinHash on Open Datasets
Finally, to get a sense for the duplicates remaining in a dataset after a full processing pipeline has been applied, we run a global (i.e., one shard) MinHash on several open datasets. These results are collated in table table˜24. We evaluate our DCLM-baseline, the official RefinedWeb dataset from HuggingFace, our emulation of the RefinedWeb pipeline, and Dolma V1. MinHash is performed using 14 buckets and a bucket size of 9, corresponding to the green curve in fig.˜6.
We note several observations here. First, we note that pools deduplicated with a Bloom Filter still have large numbers of "fuzzy duplicates" in the MinHash/Jaccard Similarity sense. This indicates that what the Bloom Filter considers a duplicate and what MinHash considers duplicates are not identical concepts. Second, we see that while MinHash is a roughly idempotent procedure, deduplication over shards fails to remove a large portion of the duplicates. Third, we see that our 100-shard Bloom filter deduplication applied to DCLM-baseline still leaves many duplicates in the dataset, yet does not seem to adversely effect downstream performance. This calls into question the general prevailing thought that the presence of any duplicates hinders downstream performance: we instead conjecture that either i) only large amounts of duplicates are detrimental to downstream performance, or ii) aggressive single-sharded deduplication eliminates many high quality documents. We leave such experimentation for future work.
Table 24: Global MinHash on Open Datasets We perform a global MinHash on several open datasets and evaluate the number of duplicates that would be removed. We denote the deduplication applied to generate each pool and the number of shards used (* implies inferred sharding). DolmaV1 contained approximately 600M documents containing only the empty string, so we report numbers with and without the empty strings in the dataset.
Dataset Num Documents Deduplication Applied Shards MinHash Removal Rate
DCLM-baseline 3.2B (Fuzzy) Bloom Filter 100 85%
RefinedWeb (official) 968B MinHash+SA 1* 0%
RefinedWeb (ours) 2.0B MinHash+SA 16 45%
Dolma V1 (w/ empty) 5.2B (Exact) Bloom Filter 1 43%
Dolma V1 (w/o empty) 4.6B (Exact) Bloom Filter 1 36%
Appendix L Mixing sources
In Section˜4.5, we showed that mixing our dataset with the usual sources did not improve its general performance, and hypothesized that this is due to the more stringent filtering performed for our Common Crawl portion. One could argue that improved filtering in the other sources could lead to similar improvements in performance. As such, we perform an experiment where we apply the same fastText classifier for filtering the other sources, as we do for our DCLM-baseline.
We take several source from RedPajama [156], and filter them with the fastText classifier applied in our DCLM-baseline, while keeping only the highest scored ones. We then add the resulting data to our pretraining dataset, and train models at the 1B-1x scale. The results of this can be seen in Table˜25. We see that despite the more uniform handling of mixing across various sources, the additional sources still decrease performance.
We leave further analysis on potential mixtures of our datasets with other sources for future work. We also leverage the mixing track as a way for participants to explore such directions.
Table 25: Mixing with filtered data. We evaluate our models on mixtures of data, where we combine our DCLM-baseline with filtered data from other sources of RedPajama [156]. We find that the case where we use only DCLM-baseline performs the best in our experiments. Evaluation is done at the 1B-1x scale.
Dataset Mixture Core MMLU Extended
DCLM-baseline only 31.7 26.5 16.6
DCLM-baseline + Filtered Wiki 31.0 24.9 16.3
DCLM-baseline + Filtered Books 30.6 25.8 15.5
DCLM-baseline + Filtered Arxiv 31.5 26.0 15.6
DCLM-baseline + Filtered Github 30.0 24.4 15.2
Appendix M Human judgment
Figure 8: Accuracy measurements against ROC-AUC of different quality filters on subsets of our human annotated samples. Top: Majority, bottom: Agreement. Left: Core score, middle: StrategyQA, and right: SQuAD. All models share the same scale (1B-1x) and training hyperparameters and are based on the same pre-filtered pool, using similar filtering-ratios for different classifiers (keeping top $\sim 15\%$ of the pool). The horizontal line marks the baseline score of a model trained on random subset of the unfiltered pool. While it may seem there is some positive correlation for StrategyQA, the opposite is true for SQuAD and in both cases the $R^{2}<0.3$. Similar to what seen in Core score, for almost all other tasks, there is no apparent relationship.
Prior work suggests that using human annotators may introduce undesired bias or noise into the data due to under-training of the annotators for the task, lack of skill or motivation, or unintended leakage of subjective bias [61, 59, 37]. However, human annotators are still widely considered the gold standard for annotating data with a clear task at hand. A natural hypothesis is that if human annotators could manually filter the large pool of raw data, we would end up with a particularly high-quality dataset. To test this, we ask 16 English-speaking AI graduate students and professors to annotate approximately 500 randomly selected documents from a pool of data without a quality filter. We obtain three annotations per document and use the majority vote in each as the gold label. The average inter-annotator agreement is 71%. We further extract the subset of 281 samples where all three annotators are in agreement, naming the full data Majority and the subset Agreement.
We then evaluate various quality filters from Section˜4.4 on this data to search for correlation between dataset quality (as measured by Core accuracy) and filter agreement with human labels. Figure˜8 (a) depicts the Core scores of models trained on datasets filtered with the respective quality filter against ROC-AUC [^5] ROC-AUC measures a classifier’s ability to distinguish between classes by summarizing the trade-off between true positive and false positive rates across all thresholds, making it a robust and common metric for model performance. See https://en.wikipedia.org/wiki/Receiver_operating_characteristic for further details. of our quality-filters on the Majority data. Notably, both the best and worst fastText-based filters score about the same on the Majority data ($\sim 73\%$ ROC-AUC), while the AskLLM filter that is highly correlated with human annotations ($\penalty 10000\ \sim 82\%$ ROC-AUC) performs much worse as a quality filter ($\sim 28.5\%$ Core compared to $31\%$ in several fastText classifiers).
We continue this study by inspecting correlations to specific downstream tasks and comparing them to the ROC-AUC on the Agreement data, where all three annotators agreed on the label. Figure˜8 depicts the scores on a few tasks against the ROC-AUC on the annotated data of the representative set of quality filters. While some positive correlation may be observed for StrategyQA [60], the opposite is true for other QA datasets such as SQuAD [130], and in both cases the $R^{2}<0.3$. In most other downstream tasks, the results are similar to Figure˜8 (a), where no correlation can be observed. This suggests that human intuition may not reliably identify the most useful documents for language model training purposes. We hypothesize that human curators may create datasets that lack sufficient diversity and leave further investigation of these hypotheses to future research.
Figure 9: Histogram of length in words for samples in our human-annotated data (capped at 2,000).
Collecting the data.
We sampled 499 random documents from our pool, after going through the rule-based quality filters and deduplication. Figure˜9 shows a histogram of the length of the documents in words. We asked 16 English speak AI graduate and professors to annotate each example as a good candidate to be included in an LM pretraining corpus (see instructions and some examples given to annotators below). [^6] All annotation efforts were done voluntarily and were not paid for. Out of the 499 samples, in 281 samples there was full agreement between all three annotators. We release both datasets in https://datacomp.ai/dclm.
Instructions given to annotators
        {internallinenumbers*} Your task is to assess the quality of the documents (0-bad, 1-good), sampled randomly from RefinedWebV2, in terms of their usefulness for LLM training.              How do you judge the usefulness of a document?              While this is a subjective task, there are a few things to keep in mind:        {internallinenumbers*} 1. We encourage you to review the details of the evaluation tasks that the LLMs are expected to excel in [link to spreadsheet with the tasks was provided]. Broadly, these tasks include reading comprehension, language understanding, world knowledge, and commonsense reasoning.        {internallinenumbers*} 2. Check whether a particular document will be useful for answering any open-ended imperative tasks or questions. For example: writing a letter to a friend, writing a poem, replying to an email, cooking a dish, etc.        {internallinenumbers*} 3. Check whether the data contains harmful content that you would not want an LLM to generate. This can include personal information, vulgar language, etc.        4. See the examples below to build some intuition.
Examples
Doc1 (Bad)
        Welcome to the Southern California Connection        Prayer Line Ministry        Sponsored by the Central Filipino Church - Women’s Ministry Department        Gwen Shorter - Women’s Ministry Director        Joanne Williams - Prayer Leader        Dial: (712) 432-0075        Enter Participant Code: 624255        To hear the most recent call, please dial (712) 432-1085        Enter Participant Code: 624255        Sunday through Friday - at 7 am and 7 pm        Psalm 91 (King James Version)Prayer Line Moderators        |Sunday, Monday and Friday        |Joanne Williams, Marsha Harold        |Tuesday, Wed., and Thursday        |Gwen Shorter, Gloria Duckett        Note: Moderators schedule may vary from time to time.        Webmasters: Michael Wong, Will Fults        Hard Copies - For CD’s of programs (\$3.00 each) Please call Gloria @ 541-476-0038        Click on the links below for some detailed information:        Prayer Line Speakers - Biographys and Photographs        Prayer Line Recordings by Date        View Prayer Line Testimonials - Submit Prayer Line Testimonial        Download, Print and Share:        |free web hit counter        As of December 9, 2010.
Doc2 (Bad)
        Host: Zander        Program Category: Music        Frequency: Weekly        Length: 2 Hours        Terms: Barter        Delivery Method: Internet        {internallinenumbers*} |“Zander’s knowledge of music and his straight-forward approach has struck a huge interest among our listeners. The Rockin’ 80’s is EXACTLY what we’ve been looking for!"" - Terry West, WQLA        {internallinenumbers*} The Rockin’ 80’s is the only 80’s show with a mix of the best rock from the decade of excess plus “oh wow” tracks that add spice to the weekly line up. The two hour version of the show features rarities from the 80’s “Lost and Found”, an 80’s “Two-Fer,” spotlighting two contrasting songs from one band played back to back and much more. Featured core musical artists include Guns N’ Roses, Motley Crue, Van Halen, Rush and AC/DC.        {internallinenumbers*} |3733 Park East Drive • Room 222 • Cleveland, Ohio 44122 P: 216-831-3761 • F: 216-514-4699 • Email us        |©2014 Envision Networks. All rights reserved.        Site design by Single Source Marketing
Doc3 (Good)
        |Chinese Five-Spice Noodles with Shitake Mushrooms        |Recipes - Chinese Five Spice        Chinese Five-Spice Noodles with Shitake Mushrooms        {internallinenumbers*} Cook the noodles until tender according to the package directions. Drain and rinse under cold running water, and drain well again.        {internallinenumbers*} Bring 2 1/2 cups water to a boil in a small saucepan. Add the dried mushrooms, cover, and simmer for 3 minutes. Strain the liquid through a paper coffee filter into a bowl to remove any grit, then squeeze the mushrooms over the bowl. Roughly chop the mushrooms, then set aside with the liquid.        {internallinenumbers*} Heat the butter and oil in a wok or large saucepan over medium heat. Add the Chinese five-spice powder and shallots and saute, stirring frequently, until tender, about 2 minutes. Add the garlic and saute 2 minutes more. Increase the heat to medium-high and add the fresh and reconstituted-dried shitakes with 1/2 cup of the reserved mushroom water. Cook, stirring frequently, until the mushrooms are tender, about 3 minutes. Add the soy sauce and rice wine. Increase the heat to high and cook, stirring, for 1 minute. Add the remaining mushroom-soaking water.        {internallinenumbers*} Add the noodles to the wok and stir until heated through and coated with the sauce, about 1 minute. Garnish with the green onions and sesame seeds, if using, and serve at once.        Makes 4 servings
Doc4 (Good)
        {internallinenumbers*} "Is it necessary to purchase a travel book or is it realistic that we can get similar information from other resources? Usually, most individuals have a major question on buying a travel book. So here are the pros and cons of purchasing one such book.        Advantages of a Travel Book        {internallinenumbers*} A travel book, which may be a paperback or e-book, comes in handy while traveling. Glancing through a travel book enables you to understand the custom and culture of a particular place in the world. So you can adapt yourself to that particular environment and stay there comfortably for longer periods.        {internallinenumbers*} - They Come In Handy — The travel guide comes in various forms such as, e-books, paperbacks and the file formats. You can have easy access to these books, which would assist you with all details compatible to the region you are traveling to.        {internallinenumbers*} - They Provide Enormous Information — Electronic or traditional travel guides provide you with answers to all types of questions such as how to learn some sayings that can be used in the place where you are traveling to? How to get data on where to reside, what to see and where to eat? How to get a clear knowledge about the history of a specific region or the atmosphere that it has?        {internallinenumbers*} - They Suit To Your Requirements — To access full information about a specific country or a region, both types of general and specific travel books are made available. The e-book may easily fit into your e-book reader whereas the paperback can fit into your backpack.        Disadvantages of Travel Book        {internallinenumbers*} - The Price — The e-book and paperback travel guides are very expensive compared to the information obtained from travel websites or from those who have moved or traveled to that region.        {internallinenumbers*} - Qualitative Images In Travel Books — Most travel books are in black and white. Only a few e-books consist of colored photos. Hence make a thorough revision before purchasing a travel guide or an e-book.        {internallinenumbers*} - Travel Books Make The Trip Less Natural — Traveling can be made more spontaneous by acquiring suggestions from locals than from travel books."
Appendix N Decontamination
In Section˜4.6, we examined the results of the contamination analysis with respect to MMLU. Here, we present some analysis with respect to the other validation sets. Instead of the MMLU-specific decontamination that we performed in Section˜4.6, here we follow a more general approach based on token overlaps.
Overall, a generally applicable decontamination rule is difficult to specify, given the potentially subjective nature of what constitutes as contamination in text data as well as the diversity in formats across tasks. Following Touvron et al. [158], we search for contaminated tokens that exist in overlapping 10-grams (or longer) between DCLM-baseline and our downstream tasks. We measure the percentage of samples in each evaluation set where more than 80% of the tokens are contaminated (such samples are considered “dirty” per Touvron et al. [158]), as well as the percentage where less than 20% of the tokens are contaminated (considered “clean” by the same criterion).
We examine the difference of performance of the same 7B-2x model trained on DCLM-baseline, between the full evaluation set and the evaluation samples that are marked as “not dirty” per the criterion in Touvron et al. [158] (less than 80 % of the tokens in the sample are marked as contaminated), and between the full evaluation set and samples marked as “clean” using the same criterion (less than 20 % of the tokens are marked). Results can be seen in Figure˜10, where we see that the difference in performance over the full dataset and the “not dirty” samples is minimal. In fact, for BoolQ and SQuAD, which are marked as highly contaminated, our model performs slightly better on the “not dirty” subset. Moreover, in Figure˜11 we see that the difference in performance between the full evaluation set and the “clean” subset is similarly small for most datasets. We note here that it’s difficult to identify a correct threshold for what counts as a contaminated sample (as 20 % token overlap might lead to many false positives, but at the same time 80 % might be too high to detect all contaminated samples).
Appendix O Instruction tuning
Figure 10: Analysis of performance on the “not dirty” subset. The x-axis is the percentage of samples from each evaluation task where more than 80% of the tokens are contaminated (such samples are considered “dirty” per Touvron et al. [158]). The y-axis is the performance of our 7B-2x model trained on DCLM-baseline over the full training set, minus the performance on the “not dirty”. Each point is an evaluation task in our Core subset, as well as MMLU. There is no clear correlation with changes in performance over the full and the “not dirty” evaluation subsets and contamination.
Figure 11: Analysis of performance on the “clean” subset. The x-axis is the percentage of samples from each evaluation task where more than 20% of the tokens are contaminated (such samples are considered “not clean” per Touvron et al. [158]). The y-axis is the performance of our 7B-2x model trained on DCLM-baseline over the full training set, minus the performance on the “clean” subset (less than 20% of the tokens contaminated). Each point is an evaluation task in our Core subset, as well as MMLU. Most evaluation tasks (including MMLU) have similar performance in the full eval and in the “clean” subset.
Instruction tuning has emerged as a critical step to allow users to interact with pretrained language models [164, 162, 108, 111, 179]. To investigate whether models trained on DCLM-baseline can have strong instruction-following capabilities, we instruction-tune the DCLM-baseline (7B) with the OpenHermes-2.5 (OH-2.5) dataset [153]. Specifically, we train our model to predict the response given the instruction from the instruction-tuning dataset. We train DCLM-baseline using the Adam [83] optimizer for $10$ epochs with a $10\%$ warmup ratio and a cosine learning rate schedule. We perform a hyperparameter search over the three learning rates $\{1\text{e-}7,5\text{e-}6,2\text{e-}5\}$ and report the best-performing numbers on the evaluation metrics i.e., AlpacaEval 2.0 length-controlled win-rate [48]. Post-training, we generate the responses for the instructions from AlpacaEval using a sampling temperature of $0.3$ and maximum response length of $500$. We benchmark our model with relevant baselines (e.g., LLaMA-2-Chat [158], Zephyr-7B [159]) taken directly from the AlpacaEval leaderboard as well as Mistral-7B [75] finetuned in the same manner as DCLM-baseline.
Table 26: Instruction tuning results on AlpacaEval2.0. We see that DCLM-baseline w/ OH-2.5 performs similarly to Mistral-7B finetuned also on OH-2.5, indicating similar behavior during instruction tuning. Also, with better data, we see DCLM-IT can be even better and can beat many existing models of similar scales.
Model AlpacaEval2.0 LC Win-rate (%)
Our runs
DCLM-IT 16.6
Mistral-7B w/ OH-2.5 15.4
DCLM-baseline w/ OH-2.5 13.8
Reported from the leaderboard
LLaMA-3-Instruct-8B 22.9
Mistral-v0.2-7B 17.1
Mistral-7B w/ OH-2.5 16.2
Zephyr-Beta-7B 13.2
Vicuna-v1.3-13B 10.8
Gemma-Instruct-7B 10.4
Nous-Hermes-13B 9.7
DaVinci001 9.0
LLaMA-2-Chat-13B 8.4
Alpaca-7B 5.9
We present the results in Table˜26. We find that DCLM-baseline finetuned with OH-2.5 outperforms various instruct models such as Zephyr-Beta-7B and Gemma-Instruct-7B. This indicates that we can elicit high-quality responses from the pretrained DCLM-baseline model with instruction-tuning. In addition, we observe that the DCLM-baseline slightly lags behind Mistral-7B-OH-2.5 meaning DCLM-baseline is competitive with other existing models of the same scale for finetuning. The small difference in performance might be attributed to the DCLM-baseline w/ OH-2.5 having longer generations on average than Mistral-7B w/ OH-2.5 or the lesser number of tokens seen during DCLM-baseline pretraining in comparison to Mistral-7B.
A follow-up question is whether DCLM-baseline can be finetuned to be even more competitive with models of similar scale. To further improve the instruction-following capabilities of DCLM-baseline, we curate a custom dataset, DCLM-IT, by combining some of the best instruction-tuning datasets including UltraFeedback [43], Tulu-v2 SFT [74], CodeFeedback [187], OH-2.5, Nectar [190], NoRobots [129], WildChat [185], WebInstruct [180], and StarCoder2-Self-OSS-Instruct [165]. There are roughly $4$ million instances and $8$ billion tokens in this dataset. Subsequently, we perform instruction-tuning and response generation of DCLM-baseline on this dataset with the training recipe mentioned above. We present the results in Table˜26. We find that DCLM-IT outperforms DCLM-baseline w/ OH-2.5 by $2.8$ percentage points. Our results highlight that there is room to enhance the instruction-following capabilities of DCLM-baseline with better datasets such as DCLM-IT. We further clarify that the current instruction-tuned models do not undergo any alignment procedures such as PPO [139], DPO [126] or others [50, 30, 15, 104]. We leave the development of aligned versions of DCLM-baseline for future research.
Appendix P Scaling up
The final run 2.5T shown in Figure˜1 and Table˜9 uses a two stage training procedure as followed in Groeneveld et al. [66], Hu et al. [72] and Team et al. [151]. For stage 1 we use the hyper-parameters from Table˜27.
Table 27: Hyper-parameters for large scale run. Note the LR schedule uses a training length of 4.4T, but we do not train for the full length as we stop early and cooldown.
Hyper-parameter/Config Value
Training Tokens 4,409,222,758,400
Warmup Steps 10,000
Initial Learning Rate $2\times 10^{-3}$
Weight Decay 0.05
Final Learning Rate $3\times 10^{-5}$
Global Batch Size 2048
Accumulation Steps 2
Query-Key Normalization True
Z Loss $5\times 10^{-6}$
Table 28: Results for 2.5T run, first row was run for 2T + 200B (cooldown), second row was run for 2T + 270B (cooldown), third is evaluation of average of weights of first two rows (0.2*CoolDown #1 + 0.8*CoolDown #2)
Model MMLU Core Extended
CoolDown #1 (200B Tokens) 62.7 55.9 43.8
CoolDown #2 (270B Tokens) 63.4 55.9 44.3
Final Model Soup 63.9 56.0 43.7
After 2T tokens, we cooldown on a re-weighted pre-training distribution. For the cooldown distribution we use a mix of 70% DCLM-baseline with a tighter fastText threshold (top 7% rather than top 10%) and 30% ProofPile. We keep all the hyperparameters the same as Table˜27, so we cooldown to the same final learning rate, just over a smaller number of tokens. Before the cool-down MMLU performance was approximately 52%, and the LR was approximately $1\times 10^{-3}$.
We performed 2 independent cooldowns, one for 270B tokens and another for 200B tokens, and created a “model soup" [169] with a weight of 0.8 on the 270B cooldown and a weight of 0.2 on the 200B cooldown. Thus we the total number of tokens seen by this model is 2.5T. We present results of each individual cooldown and the model soup in Table˜28. The model in Figure˜1 and Table˜9 uses the final model soup after long-context training for 100B tokens as described in Section˜P.2.
Instruction tuning the scaled up model
In Appendix˜O we show how instruction tuning the above “model soup" for 80B additional tokens leads to strong performance on instruction tuning benchmarks and out-performs instruction tuned variants of similar 7B models such as Gemma-7B. In addition to the IT benchmarks covered in Appendix˜O, In Table˜29 we show that a small amount of instruction tuning provides large improvements in “Extended" evals at the cost of a small degradation in “Core" and “MMLU" evals. Notably we note that our GSM8k performance goes 2.5% to 52.5% which is comparable to other similar language models that mixed IT data into pretraining such as Gemma-7B.
Table 29: Effect of Instruction Tuning. We compare our final model with its instruction-tuned variant, both trained on a 4k context length. Including instruction tuning maintains performance on language tasks such as MMLU and results in considerable gains on 5-shot GSM8K with chain-of-thought, demonstrating the effectiveness of this training in performing complex reasoning.
Model Params Tokens Core Extended MMLU GSM8K
DCLM-baseline 7B 2.5T 56.0 43.7 63.9 2.1
DCLM-baseline-IT 7B 2.58T 55.0 46.5 62.9 52.5
Continual learning to extend context length
In this section, we present continual learning results for adapting the above DCLM-baseline 7B model (with an original context length of 2048) to a context length of 8192, similar to [175]. We follow the continual learning recipe described in [73], loading the DCLM-baseline 7B checkpoint and warming up to a maximum learning rate of $10^{-4}$ over 2000 steps, then annealing with a cosine schedule to $10^{-5}$. All other hyper-parameters remain the same as original pretraining. The global batch size remains $2^{22}$ tokens per optimization step. We employ a variable sequence length curriculum as in [123], including batches of sequences ranging from 64 to 8192 in length. For this continual learning stage, we train with a total of $\sim 120$B tokens randomly sampled from the main dataset and distributed as follows among different sequence lengths: $64:2^{33},128:2^{33},256:2^{33},512:2^{33},1024:2^{33},2048:2^{33},4096:2^{35},8192:2^{35}$. We use the Grow-Linear curriculum (from short to long sequences) with 4 cycles as described in [123]. As proposed by [122] and similar to [175] for long-context continual learning, we increase the RoPE [149] base frequency from 10,000 to 100,000 during the continual learning stage for long context adaptation. The average context length for 20-Docs and 30-Docs is $\sim 4k$ and $\sim 6k$, respectively. Hence, the original DCLM with context length of 2048 model has poor performance for these benchmarks.
We show that the above strategy results in similar performance on regular evaluations as the starting checkpoint and significantly improves on the multi-document question-answering evaluation. We use the evaluation setup described in [93]: the context is filled with $k$ documents followed by a question. We ensure that one of the $k$ documents includes the answer to the question (a.k.a., golden document). We use $k=1,10,20,30$, and for each case, we run the evaluation multiple times by changing the position of the golden document in the context and report the average. Results are reported in table˜30. We demonstrate that long context adaptation results in a checkpoint (DCLM-8k) that matches the original model on regular evaluations and significantly improves multi-document QA showing its long-context capabilities.
Table 30: Regular and long-context evaluations for DCLM-Baseline 7B model, and DCLM-8k 7B model that is adapted to 8192 context length through continual learning for additional $\sim 120$B tokens.
Model Params Tokens Context length Regular Evaluations Multi-Document Evaluations
Core MMLU Extended 1-Doc 10-Docs 20-Docs 30-Docs
Llama-2 7B 2T 4096 49.2 45.8 34.1 48.5 27.3 25.6 NA
DCLM 7B 2.5T 2048 56.0 63.9 43.7 72.0 43.4 NA NA
DCLM-8k 7B 2.6T 8192 57.1 63.7 45.4 76.9 49.8 46.1 38.8
Appendix Q Account of compute costs
We note the compute cost of training runs for each competition scale in Table˜1. In total, we estimate that our runs for DCLM sum up to approximately 1M H100 hours. Our precise estimate from our experimental test bed is 772K H100 hours for training, but this is likely an underestimate due to additional compute that was not tracked, such as due to training failures.
Appendix R Existing assets used
In this section, we describe the assets we use in our benchmark and their associated licenses.
Evaluation data
Appendix˜G discusses all downstream tasks we use for our evaluation. Below we mention them again, and specify their licenses.
Raw sources
Our main external asset used in constructing DCLM-Pool and its filtered version DCLM-baseline is Common Crawl [40]. In their Terms of Use (https://commoncrawl.org/terms-of-use), they grant a limited, non-transferable license to access and use their service, primarily for innovation, education, and research, with several restrictions on usage. While being relatively permissive, it does not conform to any specific common licenses and emphasize that the usage must comply with local and international laws, and users must respect third-party copyrights. We urge the user’s discretion in verifying their use abide by these terms-of-use.
In addition to the above, as described in Sections˜5, 4.5 and 4.4 and Appendices˜L and O we make use of the following datasets:
  1. 1.
    OpenHermes2.5 [153] for instruction finetuning and to train some of our quality filters. While the authors do not provide a specific license and refer users to determine the license by following the links for the subsets they use [^7] https://huggingface.co/datasets/teknium/OpenHermes-2.5/discussions/9#65f2a0254ab77537428cc000, we note that the dataset is based in part on outputs from OpenAI models, and thus cannot be used for training new models for commercial purposes.
  2. 2.
    StarCoder [88] and StarCoder2 [98] are used for some of our ablations (Section˜5. While constructed from permissive data by extracting datasets that mention permissive licenses (e.g. MIT, Apache 2.0), they involve various licenses, and as described in the Terms of Use [^8] https://huggingface.co/datasets/bigcode/the-stack#terms-of-use-for-the-stack, require the user to follow all terms-of-use and licenses of the different datasets it comprises of.
  3. 3.
    ProofPile2 [13] is used to scale up the dataset to the trillion tokens scale (Section˜5. The authors do not alter the licenses of underlying datasets and ask users to follow guidelines and licenses as described in these datasets.
  4. 4.
    GSM8k [39] was used in some of the ablations in Section˜5 and follow the MIT license.
  5. 5.
    RedPajama [156] is used for ablations in Section˜4.5 and Appendix˜L. Note that we do not release models or datasets that include this data. RedPajama filters the datasets it uses keeping only permissive licenses, and refers the user to adhere to underlying licenses where appropriate, as described in https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T.
  6. 6.
    UltraFeedback [43] is used for instruction tuning and is under the MIT License.
  7. 7.
    Tulu V2 SFT mixture [74] is used for instruction tuning and is under the Open Data Commons License Attribution family.
  8. 8.
    CodeFeedback [187] is used for instruction tuning and is under the Apache 2.0 License.
  9. 9.
    Nectar [190] is used for instruction tuning and is under the Apache 2.0 License.
  10. 10.
    NoRobots [129] is used for instruction tuning and is under the Creative Commons Attribution Non-Commercial 4.0.
  11. 11.
    WildChat [185] is used for instruction tuning and is under the AI2 ImpACT License - Low Risk Artifacts ("LR Agreement").
  12. 12.
    WebInstruct [180] is used for instruction tuning and is under the Apache 2.0 License.
  13. 13.
    StarCoder2-Self-OSS-Instruct [165] is used for instruction tuning and is under the Open Data Commons License Attribution family.
Libraries
The main libraries used in our benchmark pipeline are:
  1. 1.
    transformers uses the Apache 2.0 License. [^9] https://github.com/huggingface/transformers/blob/main/LICENSE
  2. 2.
    PyTorch uses a similar license to the 3-caluse BSD, and is defined in https://github.com/pytorch/pytorch/blob/main/LICENSE.
  3. 3.
    OpenLM [68] which is provided with MIT license. [^10] https://github.com/mlfoundations/open_lm/blob/main/LICENSE
  4. 4.
    llm-foundry uses the Apache 2.0 License. [^11] https://github.com/mosaicml/llm-foundry/blob/main/LICENSE
  5. 5.
    ChatNoir Resiliparse uses the Apache 2.0 License. [^12] https://github.com/chatnoir-eu/chatnoir-resiliparse/blob/develop/LICENSE
  6. 6.
    BFF uses the Apache 2.0 License. [^13] https://github.com/allenai/bff/blob/main/LICENSE
  7. 7.
    Ray uses the Apache 2.0 License. [^14] https://github.com/ray-project/ray/blob/master/LICENSE
  8. 8.
    slurm is accessible under the GPL license. [^15] https://github.com/SchedMD/slurm/tree/master?tab=License-1-ov-file
  9. 9.
  10. 10.
    nltk uses the Apache 2.0 License. [^17] https://github.com/nltk/nltk/blob/develop/LICENSE.txt
  11. 11.
    langdetect uses the Apache 2.0 License. [^18] https://github.com/Mimino666/langdetect/blob/master/LICENSE
In addition, the installation may include common ML and web development packages, and we urge commercial users to verify their endowment to refrain from license violations.
Appendix S Datasheet
Motivation
  1. Q1
    For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
    • The purpose of DCLM and the associated DCLM-Pool and DCLM-baseline datasets is to enable the study of what makes a strong pretraining dataset for large language models. These models are transformative to society and act as the foundation of numerous applications, but they are often associated with steep costs. While prior work explores many curation techniques, it is often coupled with various architectural and training design choices and evaluated in different settings, making controlled comparison nearly impossible. This slows down progress and forces a lot of duplicate work between research teams. Prior work mainly focuses on data curation in the context of supervised datasets and smaller scales (see Section˜2 and Appendix˜B). In our initial release of DCLM, we focus on 53 downstream language understanding tasks that also include reasoning abilities, math, code, and more. For details see Section˜3.5 and Appendix˜G.
  2. Q2
    Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
    • DCLM-Pool and DCLM-baseline were created by a group of researchers with the following affiliations, listed in alphabetical order: Allen Institute for Artificial Intelligence, Apple, Carnegie Mellon University, Columbia University, Contextual AI, Cornell University, DatologyAI, Harvard University, Hebrew University, Juelich Supercomputing Center, Research Center Juelich, SambaNova Systems, Stanford University, SynthLabs, Tel Aviv University, Toyota Research Institute, TU Munich, University of California, Los Angeles, University of California, Santa Barbara, University of Southern California, The University of Texas at Austin, University of Washington.
  3. Q3
    Who funded the creation of the dataset? If there is an associated grant, please provide the name of the grantor and the grant name and number.
    • Funding for this research was generously provided by the University of Washington, the University of Texas (Austin), the Institute for Foundations of Machine Learning (IFML), and Open Philanthropy.
  4. Q4
    Any other comments?
    • We anticipate that DCLM benchmark, tooling and pools will drive data-centric research in ML and AI, fostering the development of the next generation of web-scale datasets, enhancing model abilities, lowering training costs and develop knowledge sharing across research teams.
Composition
  1. Q5
    What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)? Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)? Please provide a description.
    • Each instance represented a web-crawled page (document). It contains the URL and the corresponding HTML content. Each sample is also tagged with metadata about its crawl time and additional information such as the detected language, for processed instances such as those in DCLM-baseline. Additional information can be found in Appendix˜E.
  2. Q6
    How many instances are there in total (of each type, if appropriate)?
    • DCLM-Pool contains $\sim$200B documents, all of which are of the same instance, and comes from hundreds of millions of different sources. The subset DCLM-baseline contains approximately 3B documents.
  3. Q7
    Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? If the dataset is a sample, then what is the larger set? Is the sample representative of the larger set (e.g., geographic coverage)? If so, please describe how this representativeness was validated/verified. If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
    • DCLM-Pool is an unfiltered web-text corpus comprised of all Common Crawl data prior to 2023. As such, it represent the full breadth of possible instances from this source. However, we note that Common Crawl does not cover the entire web data, due to reach and compute limitations for instance. For our DCLM-baseline, we use various filtering and deduplication strategies as described in Section˜4 in the explicit attempt to improve its quality for preatining, thus removing low-quality instances, and in doing so, becoming non-representative of the full set of instances. For a complete treatment and visualization of our data processing funnel, see Sections˜4, 4.3 and 4.2 and Appendix˜E.
  4. Q8
    What data does each instance consist of? “Raw” data (e.g., unprocessed text or images) or features? In either case, please provide a description.
    • Each sample contains an a web-page url for and the extracted HTML content associated with. Additionally, each sample contains metadata fields shown in Table˜10 (e.g., WARC-Type, WARC-date, Content-Type etc.).
  5. Q9
    Is there a label or target associated with each instance? If so, please provide a description.
    • We do not provide any labels associated with the samples, as they are used to pretrain language models by performing self-supervised next-token prediction.
  6. Q10
    Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
    • No, each sample is the full text as extracted from the HTML content, and the respective metadata.
  7. Q11
    Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are made explicit.
    • No, the dataset is released as it is with no explicit attempt to establish relationships between instances. Some links may be drawn based on metadata information such the as the source URL, but we do not deliberately form any such connections.
  8. Q12
    Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
    • No. The evaluation procedure is made of tasks as described in Section˜3.5. We also attempt to prevent test set contamination in as described in Section˜4.6 and Appendix˜N.
  9. Q13
    Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
    • DCLM-Pool is based on Common Crawl, which can be thought of as a snapshot of the internet at a given time. Hence, there can be considerable noise (e.g., placeholder text, broken links, failed extraction of HTML content, duplicate data, etc.)
  10. Q14
    Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? If it links to or relies on external resources, a) are there guarantees that they will exist, and remain constant, over time; b) are there official archival versions of the complete dataset (i.e., including the external resources as they existed at the time the dataset was created); c) are there any restrictions (e.g., licenses, fees) associated with any of the external resources that might apply to a future user? Please provide descriptions of all external resources and any restrictions associated with them, as well as links or other access points, as appropriate.
    • Each sample is associated with a URL that links other external resources on the internet with no guarantee that the resources will exist in perpetuity or that that the resources will not change. However, the dataset itself contains already extracted HTML content and is thus self-contained for the purposes of this benchmark as described in Appendix˜C.
  11. Q15
    Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor–patient confidentiality, data that includes the content of individuals’ non-public communications)? If so, please provide a description.
    • The dataset consists of data that was publicly accessible on the internet at the time of collection. However, it is possible that some of the data may include confidential information, such as private data that is unintentionally or maliciously made public.
  12. Q16
    Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
    • Given the diverse backgrounds of individuals worldwide, it is highly plausible that DCLM-Pool contains content that could be upsetting. Since our dataset consists of text scraped from the internet, it may include hateful, racist, sexist, and other offensive or toxic material. We consider the dataset a research artifact and hope future work will critically examine DCLM-Pool to develop improved safety filters. Our processed dataset, DCLM-baseline does apply a reproduction of the content-filtering from RefinedWeb. This involves url-based filtering using a domain banlist curated from Blacklists UT1191919https://dsi.ut-capitole.fr/blacklists/index_en.php and a set of banned url-substrings curated from the LDNOOBW  [^20] https://github.com/LDNOOBW/List-of-Dirty-Naughty-Obscene-and-Otherwise-Bad-Words/blob/master/en list. While these banlists are extensive, they may still let in content that is harmful.
  13. Q17
    Does the dataset relate to people? If not, you may skip the remaining questions in this section.
    • As a snapshot of the Internet, the dataset may include information about people which they shared intentionally or that was shared about them without permission.
  14. Q18
    Does the dataset identify any subpopulations (e.g., by age, gender)?
    • Our DCLM-Pool does not explicitly identify subpopulations in its metadata, as it is unclear how one can define such division over raw text data from the web.
  15. Q19
    Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset? If so, please describe how.
    • As names and other identifiers are frequent in web data, it is likely that some content can be linked back to specific individuals. However, in most public sites which Common Crawl scrape people publish such information willingly, knowing it will be visible and public.
  16. Q20
    Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals racial or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)? If so, please provide a description.
    • Yes. DCLM-Pool is created from data that is available on the public internet. Since people often debate their political views, sexual preferences, religious beliefs and other such information, it is highly likely such information is contained in the dataset. While such information is often published willingly in the explicit intent that it will be publicly visible (see Q19), we do encourage additional research on filtering such data both to preserve privacy as well as to discard any potentially biased or toxic content from the training data of the models.
  17. Q21
    Any other comments?
    • DCLM-Pool is a research artifact, and we aim for it to be useful for those studying ways to make internet-scale datasets safer.
Collection process
  1. Q22
    How was the data associated with each instance acquired? Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)? If data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified? If so, please describe how.
    • Data is directly scraped from the public internet by Common Crawl.
  2. Q23
    What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)? How were these mechanisms or procedures validated?
    • We begin by downloading the entire Common Crawl data prior to 2023. We ran Python-based processing scripts to parse these archives, filtering low-quality or irrelevant content, deduplicate samples and in some cases decontaminate against downstream tests sets, and compute various model-based features. We ran processes on hundreds of AWS CPU nodes for Common Crawl parsing and data downloading. Model-based features were run on GPU clusters. For software links see Q37 and Appendix˜R or refer to https://datacomp.ai/dclm.
  3. Q24
    If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
    • DCLM-Pool is not a probabilistic sample. As described in Q7, DCLM-Pool contains all data from Common Crawl before 2023. Common Crawl is a sample of the Web, and we refer to Common Crawl documentation for details of their sampling process.
  4. Q25
    Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
    • The authors participated in the data collection as part of an open-source effort. No researchers received specific compensation for their contributions to this project.
  5. Q26
    Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created.
    • The data was downloaded between January 2023 and May 2023. The urls are collected from Common Crawl archives up to 2023. Common Crawl archives may include URLs from the early days of the internet. Hence, the download / collection timeframe does not match the creation timeframe. Additionally, future users of DCLM-Pool and its subsets will have to download data themselves using our tooling, though the snapshot should not be altered in any way.
  6. Q27
    Were any ethical review processes conducted (e.g., by an institutional review board)? If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
    • A formal ethics review / IRB has not been conducted to date because DCLM-Pool contains only data that is already publicly available as part of Common Crawl.
  7. Q28
    Does the dataset relate to people? If not, you may skip the remaining questions in this section.
    • Yes. As described in Q17, people’s data may appear as part of the data scraped.
  8. Q29
    Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
    • The data was gathered from data scattered across the web.
  9. Q30
    Were the individuals in question notified about the data collection? If so, please describe (or show with screenshots or other information) how notice was provided, and provide a link or other access point to, or otherwise reproduce, the exact language of the notification itself.
    • Individuals were not notified about the data collection.
  10. Q31
    Did the individuals in question consent to the collection and use of their data? If so, please describe (or show with screenshots or other information) how consent was requested and provided, and provide a link or other access point to, or otherwise reproduce, the exact language to which the individuals consented.
    • Following our usage of Common Crawl, we respect robots.txt files, which specify parts of websites that a crawler may access. It is, however, possible that some private content of people such as personal notes, medical information or private correspondence were uploaded to the internet without a person’s consent or under the assumption the host site is private. To mitigate against such safety concerns we make an effort to exclude some malicious domains and filter such content as low quality.
  11. Q32
    If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses? If so, please provide a description, as well as a link or other access point to the mechanism (if appropriate).
    • While we have no control over the raw data scraped and hosted by Common Crawl, we will make an effort to provide user a mechanism to request exclusion of specific URLs, which can be filtered out of our DCLM-Pool and its derived datasets.
  12. Q33
    Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted? If so, please provide a description of this analysis, including the outcomes, as well as a link or other access point to any supporting documentation.
    • Luccioni & Viviano [99], Bender et al. [18] conducted such research that web-based datasets still contain substantial amounts of hate speech and sexually explicit content, even after filtering. Such content can propagate biases and harmful stereotypes when used to train language models, resulting in outputs that may be inappropriate or offensive in various contexts.
  13. Q34
    Any other comments?
    • We anticipate and hope that future studies will leverage DCLM-Pool and DCLM-baseline to investigate techniques for building better web-scale datasets.
Preprocessing, cleaning, and/or labeling
  1. Q35
    Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? If so, please provide a description. If not, you may skip the remainder of the questions in this section.
  2. Q36
    Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
    • The raw data is stored and accessible through Common Crawl. DCLM-Pool contains raw text data after HTML extraction using resiliparse.
  3. Q37
    Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point.
  4. Q38
    Any other comments?
    • The creation of DCLM-Pool, DCLM-baseline, the DCLM tooling and our trained models relies heavily on tools developed by the open-source community and would not have been possible without it.
Uses
  1. Q39
    Has the dataset been used for any tasks already? If so, please provide a description.
    • The full dataset (and subsets) have been used to train hundreds of language models at various scales and compute budgets as presented in our main paper. We evaluate these models on our testbed of 53 zero- and few-shot downstream tasks. See Sections˜4 and 3.5.
  2. Q40
    Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
    • No. There is, however, a leaderboard connected to DCLM. Those interested can review the submissions and examine publications that utilize our data. Refer to: https://datacomp.ai/dclm/leaderboard.
  3. Q41
    What (other) tasks could the dataset be used for?
    • Large language models are now widespread and used for an incredibly large spectrum of tasks, ranging from spell-checking and translation to interactive agents. The dataset could provide the necessary data to pretrain such models. DCLM-Pool could also be used for sociological studies, such as examining biases and trends in human communication, as well as studying human behavior on the public internet.
  4. Q42
    Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? For example, is there anything that a future user might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other undesirable harms (e.g., financial harms, legal risks) If so, please provide a description. Is there anything a future user could do to mitigate these undesirable harms?
    • DCLM-Pool and its related datasets and models are not designed for use in production systems, particularly those involving sensitive areas such as race, gender identity or expression, ethnicity, sexual orientation, age, socioeconomic status, disability, religion, national origin, or creed. DCLM-Pool is unsuitable for applications that involve decision-making about individuals. Since DCLM-Pool is sourced from the internet, it inherently contains biases, unfairness, and stereotypes prevalent in society. It is intended solely as a research tool to examine language-modeling dataset curation on a large scale and to study the impact of various data curation methods on downstream models.
  5. Q43
    Are there tasks for which the dataset should not be used? If so, please provide a description.
    • As mentioned in Q42, neither DCLM-Pool in its current state nor the subsets included in this paper should be used in decision-making software involving individuals. It is intended solely as a research tool for academic study.
  6. Q44
    Any other comments?
    • Our aim with DCLM-Pool and DCLM was to establish a benchmark for the community to measure dataset progress across various dimensions (e.g., model performance on diverse tasks). We consider this essential for creating more effective and safer datasets, minimizing redundant efforts, promoting knowledge sharing, and making large language model research more accessible.
Distribution
  1. Q45
    Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
    • Yes. We use HuggingFace datasets for public release.
  2. Q46
    How will the dataset be distributed (e.g., tarball on website, API, GitHub)? Does the dataset have a digital object identifier (DOI)?
    • The dataset will be distributed via HuggingFace.
  3. Q47
    When will the dataset be distributed?
    • DCLM-Pool and DCLM-baseline will be available starting June 2024.
  4. Q48
    Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
    • We distribute our datasets in full, including extracted page content and associated metadata under a standard CC-BY-4.0 licence (see Appendix˜E). The code associated with DCLM is released under the MIT license. We also note that the use of this dataset is also subject to CommonCrawl’s Terms of Use as described in https://commoncrawl.org/terms-of-use.
  5. Q49
    Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
    • We do not copyright samples in the dataset.
  6. Q50
    Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
    • No, the dataset is provided as individual samples with extracted content and associated metadata based on the content in Common Crawl hosted data.
  7. Q51
    Any other comments?
    • We provide several subsets of DCLM-Pool in different sizes, along with extensive tooling to sample from it which makes it easy for any research entity to download and experiment with the data at scale suited for them.
      We release our code and dataset as open-source with permissive licenses as described in Q48.
      We, the authors, bear all responsibility for any violation of rights associated with this dataset. While we have made maximal efforts to respect all licenses of used assets and to mitigate any risks of causing harm, the responsibility for any misuse of the dataset by others does not rest with us. This dataset is intended solely for scientific research and not for use in production systems. We strongly encourage all users to adhere to local and national laws, respect privacy, and make every effort to avoid harming anyone when using this dataset.
Maintenance
  1. Q52
    Who will be supporting/hosting/maintaining the dataset?
    • HuggingFace currently hosts the datasets. The DCLM team will be responsible for maintaining the dataset.
  2. Q53
    How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
  3. Q54
    Is there an erratum? If so, please provide a link or other access point.
    • There are no errata at this time. If any issues arise, we will inform the public through our website at https://datacomp.ai/dclm.
  4. Q55
    Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to users (e.g., mailing list, GitHub)?
    • Currently, there are no plans to update DCLM-Pool to maintain scientific integrity and comparability among participants in the DCLM competition. However, we will address user takedown requests (see Q56). DCLM-Pool is inherently noisy, and its release aims to encourage researchers to study dataset cleaning in the context of image-text samples.
  5. Q56
    If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)? If so, please describe these limits and explain how they will be enforced.
    • Until we establish an automated method for takedown requests, users can contact us through contact@datacomp.ai with takedown requests and specify the offending URL.
  6. Q57
    Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to users.
    • This is the first version of DCLM-Pool and derivative DCLM-baseline dataset. We do not intend to maintain deprecated version of DCLM-Pool. Any deprecation or modification will be announced on our website at https://datacomp.ai/dclm.
  7. Q58
    If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to other users? If so, please provide a description.
    • Each proposed modification to the dataset will be addressed individually.
  8. Q59
    Any other comments?
    • We encourage community members to contact us at contact@datacomp.ai with any suggestion or questions about dataset maintenance.

### v2

@@
 As a baseline for DCLM, we conduct extensive experiments and find that model-based filtering is key to assembling a high-quality training set.
-The resulting dataset, DCLM-baseline, enables training a 7B parameter language model from scratch to 6464% 5-shot accuracy on MMLU with 2.6T training tokens.
+The resulting dataset, DCLM-baseline, enables training a 7B parameter language model from scratch to $64$% 5-shot accuracy on MMLU with 2.6T training tokens.
 Compared to MAP-Neo, the previous state-of-the-art in open-data language models, DCLM-baseline represents a 6.6 percentage point improvement on MMLU while being trained with 40% less compute.
-Our baseline model is also comparable to Mistral-7B-v0.3 and Llama 3 8B on MMLU (63% & 66%), and performs similarly on an average of 53 natural language understanding tasks while being trained with $6.6\times$ less compute than Llama 3 8B.
+Our baseline model is also comparable to Mistral-7B-v0.3 and Llama 3 8B on MMLU (63% & 66%), and performs similarly on an average of 53 natural language understanding tasks while being trained with $\times6$ less compute than Llama 3 8B.
 Our results highlight the importance of dataset design for training language models and offer a starting point for further research on data curation.
+We release the DCLM benchmark, framework, models, and datasets at https://datacomp.ai/dclm.
 Introduction
@@
 Specifically DCLM-baseline (orange) shows favorable performance relative to both close-source models (crosses) and other open-source datasets and models (circles).
+Models in this figure are from [161, 162, 4, 97, 22, 189, 154, 121, 43, 160, 68, 150, 130, 100, 10, 156].
 A key challenge in this emerging research area is a lack of controlled comparisons.
@@
 We find that models as small as 400M parameters can still provide signal on which training sets perform better at larger scales.
-Based on our experiments, we organize DCLM into five compute scales spanning a range of about $600\times$ in compute from 400M parameter models to over-trained 7B models.
+Based on our experiments, we organize DCLM into five compute scales spanning a range of about 600×600$\times$600 × in compute from 400M parameter models to over-trained 7B models.
 This multi-scale design makes DCLM accessible to researchers with varying compute budgets.
@@
 Finally, we combine our results into DCLM-baseline, a new state-of-the-art public training set for language models.
-When training a 7B parameter language model on 2.6 trillion tokens using DCLM-baseline, the resulting model achieves 6464% on MMLU, which is state-of-the-art among open-data models and close to models such as Mistral-7B-v0.3 (63%) or Llama 3 8B (66%) that are trained with up to $6.6\times$ more compute (Llama 3 8B).
-Compared to Llama 2 7B, training a 7B parameter model on 280B tokens from DCLM-baseline achieves 5 pp higher MMLU while being trained with $7\times$ less compute.
+When training a 7B parameter language model on 2.6 trillion tokens using DCLM-baseline, the resulting model achieves $64$% on MMLU, which is state-of-the-art among open-data models and close to models such as Mistral-7B-v0.3 (63%) or Llama 3 8B (66%) that are trained with up to $\times6$ more compute (Llama 3 8B).
+Compared to Llama 2 7B, training a 7B parameter model on 280B tokens from DCLM-baseline achieves 5 pp higher MMLU while being trained with 7×7$\times$7 × less compute.
 As our 7B model uses a standard decoder-only Transformer [165, 127, 161], our results also highlight that a systematic approach to data curation is key to training performant language models.
-We publicly release our DCLM infrastructure, models, and training sets at https://datacomp.ai/dclm to enable other researchers to participate in DCLM and to strengthen the empirical foundations for data-centric research on language models.
+We publicly release our DCLM framework, models, and training sets at https://datacomp.ai/dclm to enable other researchers to participate in DCLM and to strengthen the empirical foundations for data-centric research on language models.
 Related work
-We summarize closely related work in this section and provide additional related work in Appendix˜B.
+We summarize closely related work in this section and provide additional related work in Appendix B.
 Data curation for language models.
@@
 There are also large domain-specific datasets, such as the code-focused StackV2 with 900B tokens [101], as well as high-quality filtered subsets such as FineWeb-Edu [100] with 1.3T tokens.
-We include performance comparisons with various datasets in Figure˜1 and examine FineWeb’s LightEval evaluation framework more closely in Appendix˜G.
+We include performance comparisons with various datasets in Figure 1 and examine FineWeb’s LightEval evaluation framework more closely in Appendix G.
 We release the largest pool of raw text data to date with 240T web-crawled tokens.
@@
 DCLM contains five competition scales, enabling research in varying compute regimes.
-Each scale specifies the model size (‘Model parameters’, NN), the number of tokens seen during training (‘Train tokens’, $D$), and the size of the original pool that can be used for filtering (‘Pool size’).
-We provide an estimate of the compute required for training (‘Train FLOPs’=6ND=6ND) and GPU hours (‘Train H100 hours’) using the OpenLM training framework [68].
+Each scale specifies the model size (‘Model parameters’, $N$), the number of tokens seen during training (‘Train tokens’, $D$), and the size of the original pool that can be used for filtering (‘Pool size’).
+We provide an estimate of the compute required for training (‘Train FLOPs’$=6ND$) and GPU hours (‘Train H100 hours’) using the OpenLM training framework [70].
 Scale Model parameters Train tokens Train FLOPs Train H100 hours Pool size
@@
 This section describes the main components of DCLM.
-We start with DCLM-Pool, the raw text corpus underlying our benchmark (Section˜3.1).
-We then develop the DCLM workflow, visualized in Figure˜2: selecting a competition scale (Section˜3.2), curating a dataset by filtering DCLM-Pool and potentially mixing in other sources (Section˜3.3), training a model with fixed hyperparameters (Section˜3.4), and evaluating the model to score the dataset (Section˜3.5).
+We start with DCLM-Pool, the raw text corpus underlying our benchmark (Section 3.1).
+We then develop the DCLM workflow, visualized in Figure 2: selecting a competition scale (Section 3.2), curating a dataset by filtering DCLM-Pool and potentially mixing in other sources (Section 3.3), training a model with fixed hyperparameters (Section 3.4), and evaluating the model to score the dataset (Section 3.5).
 DCLM-Pool
 DCLM-Pool is an unfiltered web-text corpus comprised of all Common Crawl [42] data prior to 2023.
-Based on Section˜4.2, we re-extract text from HTML using resiliparse [19, 20] instead of using Common Crawl’s pre-extracted text.
+Based on Section 4.2, we re-extract text from HTML using resiliparse [20, 21] instead of using Common Crawl’s pre-extracted text.
 DCLM-Pool contains 200B documents (370TB after gzip compression), resulting in 240T GPT-NeoX [24] tokens.
-See Appendix˜E for additional details.
+See Appendix E for additional details.
 Decontamination.
@@
 For the highest scoring submissions, we plan to specifically evaluate them for contamination.
-In Section˜4.6, we apply our tools to DCLM-Pool and evaluate whether contamination affects our models.
+In Section 4.6, we apply our tools to DCLM-Pool and evaluate whether contamination affects our models.
 Competition scales: Supporting participants with different compute constraints
 Figure 3: Datasets rank consistently across competition scales in DCLM. This makes it possible to iterate on data curation at small scales.
-To ensure DCLM is accessible to researchers with different compute constraints and to facilitate the study of scaling trends, we create different competition scales spanning three orders of compute magnitude (Table˜1).Each scale (i.e., 400M-1x, 1B-1x, 1B-5x, 7B-1x, and 7B-2x) specifies the number of model parameters (e.g., 7B) and a Chinchilla multiplier (e.g., 1x).
-The number of training tokens for each scale is 2020 $\times$ number of parameters $\times$ Chinchilla multiplier so that a multiplier of 1x corresponds to a compute allocation that Hoffmann et al.
+To ensure DCLM is accessible to researchers with different compute constraints and to facilitate the study of scaling trends, we create different competition scales spanning three orders of compute magnitude (Table 1).Each scale (i.e., 400M-1x, 1B-1x, 1B-5x, 7B-1x, and 7B-2x) specifies the number of model parameters (e.g., 7B) and a Chinchilla multiplier (e.g., 1x).
+The number of training tokens for each scale is $20$$\times$ number of parameters $\times$ Chinchilla multiplier so that a multiplier of 1x corresponds to a compute allocation that Hoffmann et al.
 [73] found near-optimal.
 A potential pitfall in our multi-scale design is that the ranking of data curation methods may change when increasing the compute scale.
-To better understand this concern, in Figure˜3, we plot the performance of 10 methods at the 7B-1x scale as a function of their 400M-1x and 1B-1x performance.
+To better understand this concern, in Figure 3, we plot the performance of 10 methods at the 7B-1x scale as a function of their 400M-1x and 1B-1x performance.
 We find high rank correlation between the smaller 400M-1x, 1B-1x results and larger 7B-1x results (Pearson’s $r=0.885$ and $r=0.919$, respectively), suggesting better curation strategies at smaller scales transfer to larger scales.
-For more competition scale ablations, including experiments suggesting dataset improvements are largely orthogonal to training hyperparameters, see Appendix˜H.
+For more competition scale ablations, including experiments suggesting dataset improvements are largely orthogonal to training hyperparameters, see Appendix H.
 Benchmark tracks: Filtering and mixing
@@
 (i) In the filtering track, participants propose algorithms to select training data from a candidate pool.
-We start with five pools, one for each scale in Table˜1, which are random document subsets of DCLM-Pool.
+We start with five pools, one for each scale in Table 1, which are random document subsets of DCLM-Pool.
 We restrict initial pool sizes by scale to encourage scalable filtering strategies and reflect realistic data download and storage constraints.
@@
 For instance, participants can synthesize documents from DCLM-Pool, a custom crawl, Stack Overflow, and Wikipedia.
-Appendix˜C provides detailed rules for each track, and Appendix˜D describes our open-source, extensible tooling for executing filtering and mixing operations.
+Appendix C provides detailed rules for each track, and Appendix D describes our open-source, extensible tooling for executing filtering and mixing operations.
 Training
@@
 We also provide unified data processing utilities.
-Appendix˜F contains additional training details.
+Appendix F contains additional training details.
 Evaluation
@@
 Finally, we report Extended centered accuracy, which averages the centered performance for all of our 53 tasks.
-For more metric details, see Appendix˜G.
+For more metric details, see Appendix G.
 Building high-quality training datasets with DCLM
 We now show how the DCLM workflow can lead to high-quality datasets and quantify the effect of data curation methods.
-This section describes the process of converting Common Crawl into our dataset, DCLM-baseline, as shown in Figure˜4.
+This section describes the process of converting Common Crawl into our dataset, DCLM-baseline, as shown in Figure 4.
 We provide ablation experiments for each step along the way.
-We first evaluate open-source datasets as a starting point (Section˜4.1).
-Next, we experiment with alternatives for several key phases of dataset construction: text extraction (Section˜4.2), deduplication (Section˜4.3), and model-based filtering (Section˜4.4).
-We then experiment with mixing in high-quality sources (Section˜4.5) and provide a contamination analysis (Section˜4.6).
-In Section˜5, we scale up this approach to train a 7B model for 2T tokens.
+We first evaluate open-source datasets as a starting point (Section 4.1).
+Next, we experiment with alternatives for several key phases of dataset construction: text extraction (Section 4.2), deduplication (Section 4.3), and model-based filtering (Section 4.4).
+We then experiment with mixing in high-quality sources (Section 4.5) and provide a contamination analysis (Section 4.6).
+In Section 5, we scale up this approach to train a 7B model for 2T tokens.
 Figure 4: Construction of DCLM-baseline from DCLM-Pool.
@@
 Percentages are based on the total number of original documents.
-We begin by evaluating several well-known open-source datasets (C4 [127, 46], RefinedWeb [118], RedPajama [156], and Dolma-V1 [147]) in Table˜2.
+Evaluating existing training datasets
+We begin by evaluating several well-known open-source datasets (C4 [130, 48], RefinedWeb [121], RedPajama [160], and Dolma-V1 [150]) in Table 2.
 While all four datasets use various heuristic filters and data cleaning steps, we find that RefinedWeb performs the best on our Core and Extended metrics at the 7B-1x scale.
@@
 RefinedWeb 36.9 19.8
-Evaluating existing training datasets
 Takeaway: For DCLM-baseline and other experiments, we adopt RefinedWeb’s heuristic filters.
@@
 We then apply RefinedWeb’s heuristic quality filters to each of the text extractions.
-In Table˜3, we find both resiliparse and trafilatura improve Core by at least 2.5 points over the WET extraction.
-This is significant because most open source datasets, including C4, RedPajama, and Dolma-V1, use the WET extraction, which could partially explain their worse performance in Table˜2.
-While resiliparse and trafilatura have similar downstream performance, resiliparse is $8\times$ faster to run and hence more practical for large-scale processing.
-For more analysis, see Appendix˜J.
+In Table 3, we find both resiliparse and trafilatura improve Core by at least 2.5 points over the WET extraction.
+This is significant because most open source datasets, including C4, RedPajama, and Dolma-V1, use the WET extraction, which could partially explain their worse performance in Table 2.
+While resiliparse and trafilatura have similar downstream performance, resiliparse is 8×8$\times$8 × faster to run and hence more practical for large-scale processing.
+For more analysis, see Appendix J.
 Takeaway: For DCLM-Pool and the remaining experiments, we use resiliparse to extract text.
@@
 However, our modified Bloom filter approach scales more easily to datasets surpassing 10TB.
-We provide additional analysis in Appendix˜K.
+We provide additional analysis in Appendix K.
 Takeaway: We use a Bloom filter for DCLM-baseline and MinHash for other experiments.
@@
 For the perplexity filtering and the top-k average logits strategies, we utilize a 154M parameter causal Transformer trained on a mix of English Wikipedia, the books subset of RedPajama v1, and peS2o  [149, 160].
-We compare the aforementioned approaches in Table˜4 and find that fastText-based filtering outperforms all other approaches.
+We compare the aforementioned approaches in Table 4 and find that fastText-based filtering outperforms all other approaches.
 We next aim to understand how fastText training recipes affect its effectiveness as a data filtering network [55].
 Text classifier ablations.
-To better understand the limits of fastText, we train several variants, exploring different choices for the reference data (i.e., the examples given positive labels), feature space, and filtering threshold, as shown in Table˜5.
+To better understand the limits of fastText, we train several variants, exploring different choices for the reference data (i.e., the examples given positive labels), feature space, and filtering threshold, as shown in Table 5.
 For reference positive data, we considered commonly used sources like Wikipedia [59], OpenWebText2 [59], and RedPajama-books [160], following the reference data used for GPT-3 [30].
@@
 It is natural to ask whether using OH-2.5 data for filtering could preclude additional gains from instruction-tuning.
-In Appendix˜P, we show this is not the case, further suggesting the strength and compatibility of this approach with modern fine-tuning paradigms.
+In Appendix P, we show this is not the case, further suggesting the strength and compatibility of this approach with modern fine-tuning paradigms.
 Finally, we observe that using a fairly strict threshold, which keeps the top-10% of examples, helps over more permissive top-15% and top-20% thresholds.
-We further study the unintuitive behavior of dataset filtering and its connection to human judgment in Appendix˜M.
+We further study the unintuitive behavior of dataset filtering and its connection to human judgment in Appendix M.
 Table 5: fastText ablations (7B-1x scale).
@@
 For the CC component, we consider different variants: a subset of our DCLM-baseline, RedPajama’s CC portion, RefinedWeb, and C4.
-The results in Table˜6 show that mixing improves performance for the lower-performing CC subsets (C4, RedPajama-CC, and RefinedWeb).
+The results in Table 6 show that mixing improves performance for the lower-performing CC subsets (C4, RedPajama-CC, and RefinedWeb).
 In the case of DCLM-baseline however, mixing actually hurts performance on average, which suggests it can be counterproductive given performant filtering.
-For additional mixing results, see Appendix˜L.
+For additional mixing results, see Appendix L.
 Table 6: Mixing high-quality sources with subsets of CommonCrawl (1B-1x scale).
@@
 Dataset Base w/ RPJ extras Base w/ RPJ extras
-C4 23.7 25.9 (+2.2+2.2) 12.5 13.3 (+0.8+0.8)
-RPJ CC only 24.0 25.7 (+1.7+1.7) 12.1 13.5 (+1.4+1.4)
-RefinedWeb 25.1 26.5 (+1.4+1.4) 12.9 13.1 (+0.2+0.2)
-DCLM-baseline  [^1] The DCLM-baseline subset used here relied on an earlier version of our classifier and our pool. 31.1 29.5 (1.6-1.6) 16.0 15.3 (0.7-0.7)
+C4 23.7 25.9 ($+2.2$) 12.5 13.3 ($+0.8$)
+RPJ CC only 24.0 25.7 ($+1.7$) 12.1 13.5 ($+1.4$)
+RefinedWeb 25.1 26.5 ($+1.4$) 12.9 13.1 ($+0.2$)
+DCLM-baseline  [^1] The DCLM-baseline subset used here relied on an earlier version of our classifier and our pool. 31.1 29.5 ($-1.6$) 16.0 15.3 ($-0.7$)
 Decontamination
@@
 We then train a 7B-2x model with our DCLM-baseline without the detected MMLU overlap.
-The results of this analysis can be seen in Table˜7.
+The results of this analysis can be seen in Table 7.
 We see that this removal of contaminated samples does not lead to a decrease in performance for our model.
@@
 We also apply the above removal strategy on Dolma-V1.7 [150] and FineWeb-Edu [100], to measure what the contamination differences are between DCLM-baseline and those datasets.
-The results can be seen in Table˜8.
+The results can be seen in Table 8.
 We see that our DCLM-baseline has roughly similar contamination stats as other high performing datasets according to this analysis.
-We provide further contamination analysis that extends to our entire benchmark suite in Appendix˜N.
+We provide further contamination analysis that extends to our entire benchmark suite in Appendix N.
 Table 8: MMLU overlap removal comparison.
@@
 To ensure our trained model is broadly useful, including for math and coding tasks, we combine our 3.8T DCLM-baseline with the StarCoder [90] and ProofPile2 [14] data to arrive at a 4.1T token dataset.
-We train a 7B model for 2.5T tokens on this dataset with the same hyperparameters as our largest competition scale except for an annealing phase for the 200B tokens on a modified distributed that was 70% DCLM-baseline with a tighter fastText threshold, and 30% math datasets (see Appendix˜P).
+We train a 7B model for 2.5T tokens on this dataset with the same hyperparameters as our largest competition scale except for two separate cool-downs phase for the 200B and 270B tokens on a modified distribution that was 70% DCLM-baseline with a tighter fastText threshold, and 30% math datasets (see Appendix P).
+We then take a “model soup" of these two separate cool-downs[173].
 We then adopt the continual pre-training methodology from Pouransari et al.
-[123] for 100B tokens on the same distribution to increase the context length from 2048 to 8192, we provide more details on this procedure in Section˜P.2.
-In Table˜9, we show that our model outperforms all 7B models trained on public training sets and approaches closed-data models trained for more tokens such as Llama-8B, Mistral-7B, and Gemma-7B.
-Additionally, in Appendix˜O, we show that our model achieves strong instruction-tuning performance.
+[126] for 100B tokens on the same distribution to increase the context length from 2048 to 8192, we provide more details on this procedure in Section P.2.
+In Table 9, we show that our model outperforms all 7B models trained on public training sets and approaches closed-data models trained for more tokens such as Llama-8B, Mistral-7B, and Gemma-7B.
+Additionally, in Appendix O, we show that our model achieves strong instruction-tuning performance.
 After instruction tuning on publicly available IT datasets, our model maintains most of its benchmark performance and achieves an AlpacaEval2.0 LC Win-rate of 16.6, which outperforms Gemma-Instruct (10.4), while approaching the strong performance of Mistral-v0.2-7B (17.1) and Llama3-Instruct (22.9).
@@
 Appendix
+[…]
 Appendix A Contributions
@@
 Paper writing
-Alon Albalak, Hritik Bansal, Yair Carmon, Achal Dave, Alex Dimakis, Alex Fang, Samir Gadre, Etash Guha, Reinhard Heckel, Maor Ivgi, Sedrick Keh, Jeffrey Li, Niklas Muennighoff, Sarah Pratt, Ludwig Schmidt, Vaishaal Shankar, Georgios Smyrnis
+Alon Albalak, Hritik Bansal, Yair Carmon, Achal Dave, Alexandros G.
+Dimakis, Alex Fang, Samir Gadre, Etash Guha, Reinhard Heckel, Maor Ivgi, Sedrick Keh, Jeffrey Li, Niklas Muennighoff, Sarah Pratt, Ludwig Schmidt, Vaishaal Shankar, Georgios Smyrnis
 Leadership and advising
@@
 Filtering track
-The defining characteristic of entries in the filtering track is that they form the dataset by applying a processing pipeline on the subset of DCLM-Pool corresponding to the chosen compute scale (see Table˜1) without including any external data.
+The defining characteristic of entries in the filtering track is that they form the dataset by applying a processing pipeline on the subset of DCLM-Pool corresponding to the chosen compute scale (see Table 1) without including any external data.
 The rationale behind this requirement is twofold.
@@
 For the dataset released to participants, we release various sizes of DCLM-Pool, that we make available for download.
-For details on the data, see Appendix˜E.
+For details on the data, see Appendix E.
 Processing.
@@
 We also provide scripts that train models using each configuration, and produce json files that describe a trained model in detail.
-For further training details, see Appendix˜F.
+For further training details, see Appendix F.
 Evaluation.
@@
 All of our results, including data processing, model training, evaluations, and plots included in this paper, are reproducible using our open-source framework and the recipes in https://datacomp.ai/dclm.
-We list compute requirements for our code in Appendix˜Q.
+We list compute requirements for our code in Appendix Q.
 Appendix E DCLM-Pool
@@
 We release DCLM-Pool as a set of .jsonl files similar to Dolma-V1 and RedPajama.
-We provide the fields that are in the .jsonl in Table˜10.
+We provide the fields that are in the .jsonl in Table 10.
 The entire pool is 5.1M gzip compressed .jsonl\ files, and 340TB compressed on disk.
@@
 While some of our pools, such as DCLM-baseline, underwent some filtering of malicious URLs, none have had any special treatment for PII and sensitive content to preserve representativeness of the raw data.
-For a more complete discussion on PII and consent regarding our pools, see Appendix˜S.
+For a more complete discussion on PII and consent regarding our pools, see Appendix S.
 Table 10: Metadata provided in DCLM-Pool data.
@@
 Hyperparameters.
-We detail the hyperparameters for our models in Table˜11.
-For the 400M-1x and 1B-1x, we follow hyperparameters from [56], which were tuned to optimize perplexity on a validation set containing tokens from recent arXiv papers, the OpenLM codebase itself, and news articles.
-For the 1B-1x scale, we also investigated alternative hyperparameters in Table˜13, and find the hyperparameters from [56] perform best.
-For the 7B-1x and 7B-2x, we used a higher learning rate, and a lower weight decay, guided by the hyperparameter sweep in Table˜12.
-We use a cooldown of 3ee-5 for all experiments.
-For Table˜2, we trained with a lower learning rate following [56] as these experiments were performed before our sweep.
-Specifically, we used a learning rate of 3ee-4 and weight decay of 0.33.
+[…]
 Table 11: Main models and hyperparameters used in our investigation.
@@
 Scale $n_{\text{layers}}$ $n_{\text{heads}}$ $d_{\text{model}}$ $d_{\text{head}}$ Warmup Learning rate Weight decay z-loss Batch size
-400M-1x 24 8 1,024 128 2,000 3ee-3 0.033 1ee-4 512
-1B-1x 24 16 2,048 128 5,000 3ee-3 0.033 1ee-4 256
-7B-1x, 7B-2x 32 32 4,096 128 5,000 2ee-3 0.05 5e-6 2,048
+[…]
+[…]
+[…]
 Table 12: Learning rate and weight decay sweep (7B-1x scale).
 We evaluated the impact of learning rate and weight decay on an earlier iteration of DCLM-baseline.
-Based on this sweep, we specify the settings for Table˜11 for the 7B-1x and 7B-2x scales.
+Based on this sweep, we specify the settings for Table 11 for the 7B-1x and 7B-2x scales.
 LR WD Core
-1ee-03 0.1 44.1
-2ee-03 0.05 44.8
-3ee-03 0.033 44.7
-1ee-02 0.01 43.8
+[…]
+[…]
+[…]
+[…]
 Appendix G Evaluation details
@@
 The 1B Hugging Face model trained on FineWeb-Edu [100] has shown to work well on this, so we wanted to more closely examine how LightEval evaluation scores correlate with evaluation scores from LLM Foundry.
-We present our findings in Figure˜5.
+We present our findings in Figure 5.
 Figure 5: Comparisons Between LightEval MMLU scores (x-axis) and LLM Foundry MMLU scores (y-axis).
@@
 The key difference between LightEval and LLM Foundry for multiple choice tasks like MMLU is that LightEval considers the log probabilities of entire answer sequences, whereas LLM Foundry only considers log probabilities of single letters.
-Nonetheless, Figure˜5 shows a positive correlation between the two evaluation frameworks on MMLU 0-shot accuracy.
-In Figure˜5 we were able to reproduce the MMLU scores reported in the FineWeb-Edu blog [97].
+Nonetheless, Figure 5 shows a positive correlation between the two evaluation frameworks on MMLU 0-shot accuracy.
+In Figure 5 we were able to reproduce the MMLU scores reported in the FineWeb-Edu blog [100].
 Notably, we found that LightEval indeed gave MMLU scores above random for the 1B scales, whereas in LLM Foundry, all the 1B models have accuracies around 0.25.
@@
 A potential concern is that the training recipe can change conclusions about which dataset is optimal for training, due to interaction between training hyperparameters and dataset distributions.
-To address this confounder, we show in Table˜13 that orderings between datasets are preserved for various combinations of weight decay and learning rate.
+To address this confounder, we show in Table 13 that orderings between datasets are preserved for various combinations of weight decay and learning rate.
 Moreover we find that performance gains from optimal hyper-parameter choice and dataset design tend to be orthogonal and complement each other.
-We illustrate this effect in Table˜14.
+We illustrate this effect in Table 14.
 Table 13: Rankings are stable across hyperparameters (1B-1x scale).
@@
 Appendix I Model-based quality filters
-We presented results for several different model-based quality filters in Section˜4.4.
+We presented results for several different model-based quality filters in Section 4.4.
 In this section, we describe their implementations in further detail, focusing especially on fastText classifiers which were our method of choice for DCLM-baseline.
@@
 In terms of training hyperparameters, we mostly used the default choices from the fastText package; the only hyperparameter change that we tried was to expand the feature space from unigrams only to both unigrams and bigrams (via setting the wordNgrams argument to be 2 instead of the default 1).
-This helped improve the quality of downstream filtered datasets, as shown in Table˜15 (which extends Table˜5 from Section˜4.4).
+This helped improve the quality of downstream filtered datasets, as shown in Table 15 (which extends Table 5 from Section 4.4).
 Table 15: fastText feature-space ablation (7B-1x scale). Adding bigrams to the feature space helps over the default setting of unigrams only.
@@
 The negatively labeled examples were sampled randomly from a set of documents that came from an earlier (smaller) version of our RefinedWeb reproduction.
-This version used trafilatura as the extractor instead of resiliparse, which we hypothesize might actually help for training the filtering model; as shown in Appendix˜J, trafilatura more aggressively removes boilerplate content that may appear in many pages (especiall from the same website).
+This version used trafilatura as the extractor instead of resiliparse, which we hypothesize might actually help for training the filtering model; as shown in Appendix J, trafilatura more aggressively removes boilerplate content that may appear in many pages (especiall from the same website).
 This type of content, if left in, may lead to the fastText models over-relying on these “spurious” features instead of the main contents of the page.
@@
 Other quality filtering baselines
-We also examined other quality filters, though found none as effective as the fastText methods described above, as shown in Table˜4. We now provide further details for some of these baselines.
+We also examined other quality filters, though found none as effective as the fastText methods described above, as shown in Table 4. We now provide further details for some of these baselines.
 PageRank.
@@
 Next we partitioned our RefinedWeb reproduction into quintiles based on their PageRank score and trained several models at the 1B-1x scale.
-These results are collated in Table˜16, but unfortunately no quintile performed better than a pool sampled from the union of all quintiles.
+These results are collated in Table 16, but unfortunately no quintile performed better than a pool sampled from the union of all quintiles.
 Table 16: PageRank-based filtering (1B-1x scale).
@@
 - Does it offer insights, useful facts or relevant information?
-{internallinenumbers*} - Does it teach how to comply with open-ended tasks such as writing letters, poems, emails etc.?
+- Does it teach how to comply with open-ended tasks such as writing letters, poems, emails etc.?
 - Is it free from harmful content?
-{internallinenumbers*} If most criteria are met based on the content below, indicate ’yes’ for suitable.
+If most criteria are met based on the content below, indicate ’yes’ for suitable.
 Otherwise, indicate ’no’ for unsuitable.
@@
 where <input> is replaced with the document tokens, clipped if too long, and appended with ‘‘…[The rest of the paragraph is omitted]’’ in such cases.
-While this method worked slightly better than random sampling from our pool (see Table˜4), it significantly underperformed compared to our fastText experiments.
+While this method worked slightly better than random sampling from our pool (see Table 4), it significantly underperformed compared to our fastText experiments.
 Considering the high costs associated with applying it at scale, we did not perform this experiment on a larger scale.
 Semantic deduplication.
-Following the success of the different deduplication methods we used (Appendix˜K), we studied the effect of Semantic deduplication as proposed by Abbas et al.
+Following the success of the different deduplication methods we used (Appendix K), we studied the effect of Semantic deduplication as proposed by Abbas et al.
 [1].
@@
 We sampled documents after discarding 25% of the data.
-As seen in Table˜4, this intervention only negatively impacted the trained model.
+As seen in Table 4, this intervention only negatively impacted the trained model.
 We hypothesize that the model used for embedding has a significant impact on the outcomes of this method.
-However, due to the large computational overhead when scaled, making it infeasible, we opted to rely on the deduplication methods outlined in Appendix˜K and leave this line of research for future work.
+However, due to the large computational overhead when scaled, making it infeasible, we opted to rely on the deduplication methods outlined in Appendix K and leave this line of research for future work.
 Appendix J Text extraction comparison
@@
 Profiling
-We compute basic summary statistics for each extractor based on a sample of 10 WARC files (corresponding to 900K individual pages), presenting the results in Table˜17.
+We compute basic summary statistics for each extractor based on a sample of 10 WARC files (corresponding to 900K individual pages), presenting the results in Table 17.
 Notably, both resiliparse and trafilatura result at least 2x shorter documents on average compared to WET files.
-As shown in the examples in Section˜J.2, WET files indeed contain many additional lines with seemingly little value for pre-training (e.g.
+As shown in the examples in Section J.2, WET files indeed contain many additional lines with seemingly little value for pre-training (e.g.
 navigation bars, boilerplate notices, copyright statements).
@@
 J.2.1 Example set 1
-        [Trafilatura]              {internallinenumbers*} HERE is a sampling of some of the better antiques and flea markets around the United States.
-Two or Three Times a Year        {internallinenumbers*} BRIMFIELD Route 20, Brimfield, Mass.
+     [Trafilatura]    HERE is a sampling of some of the better antiques and flea markets around the United States.
+Two or Three Times a Year  BRIMFIELD Route 20, Brimfield, Mass.
 01010; 413-245-3436.
 Second weekend of May and July, and the second weekend after Labor Day.
-{internallinenumbers*} RENNINGER’S OUTDOOR EXTRAVAGANZA Noble Street, Kutztown, Pa.; 717-385-0104.
+RENNINGER’S OUTDOOR EXTRAVAGANZA Noble Street, Kutztown, Pa.; 717-385-0104.
 Thursday, Friday and Saturday of the last weekend of April, June, September.
-{internallinenumbers*} FARMINGTON ANTIQUES WEEKEND Farmington Polo Grounds, Town Farm Road, Farmington, Conn.
+FARMINGTON ANTIQUES WEEKEND Farmington Polo Grounds, Town Farm Road, Farmington, Conn.
 06032; 508-839-9735.
@@
 1-2.
-Monthly        {internallinenumbers*} ANN ARBOR ANTIQUES MARKET, P.O.
+Monthly  ANN ARBOR ANTIQUES MARKET, P.O.
 Box 1512, Ann Arbor, Mich.
@@
 May through October, third Sunday.
-{internallinenumbers*} Continue reading the main storyKANE COUNTY FLEA MARKET, Kane County Fairgrounds, P.O.
+Continue reading the main storyKANE COUNTY FLEA MARKET, Kane County Fairgrounds, P.O.
 Box 549, St.
@@
 Year-round, first weekend.
-{internallinenumbers*} THE METROLINA EXPO, 7100 Statesville Road, Charlotte, N.C.
+THE METROLINA EXPO, 7100 Statesville Road, Charlotte, N.C.
 28213; 704-596-4643.
 Year-round, first weekend of every month.
-{internallinenumbers*} SPRINGFIELD ANTIQUE SHOW AND FLEA MARKET, Clark County Fairgrounds, Route 41, Springfield, Ohio, 45501; 513-325-0053.
+SPRINGFIELD ANTIQUE SHOW AND FLEA MARKET, Clark County Fairgrounds, Route 41, Springfield, Ohio, 45501; 513-325-0053.
 Year-round, third weekend.
-Weekly        {internallinenumbers*} BAKERSFIELD SWAP-O-RAMA, 4501 Wible Road, Bakersfield, Calif.
+Weekly  BAKERSFIELD SWAP-O-RAMA, 4501 Wible Road, Bakersfield, Calif.
 93313; 805-831-9342.
 Saturday and Sunday.
-{internallinenumbers*} LAMBERTVILLE ANTIQUE MARKET, Route 29, Lambertville, N.J.
+LAMBERTVILLE ANTIQUE MARKET, Route 29, Lambertville, N.J.
 08530.
@@
 Market on Saturday and Sunday.
-{internallinenumbers*} ATLANTA FLEA MARKET AND ANTIQUE CENTER, 5360 Peachtree Industrial Boulevard, Chamblee, Ga.
+ATLANTA FLEA MARKET AND ANTIQUE CENTER, 5360 Peachtree Industrial Boulevard, Chamblee, Ga.
 30341; 404-458-0456.
@@
 Continue reading the main story
-        [Resiliparse]              {internallinenumbers*} This is a digitized version of an article from The Times’s print archive, before the start of online publication in 1996.
+     [Resiliparse]    This is a digitized version of an article from The Times’s print archive, before the start of online publication in 1996.
 To preserve these articles as they originally appeared, The Times does not alter, edit or update them.
-{internallinenumbers*} Occasionally the digitization process introduces transcription errors or other problems.
+Occasionally the digitization process introduces transcription errors or other problems.
 Please send reports of such problems to archive_feedback@nytimes.com.
-May 10, 1990, Page 00006 The New York Times Archives              {internallinenumbers*} HERE is a sampling of some of the better antiques and flea markets around the United States.
-Two or Three Times a Year              {internallinenumbers*} BRIMFIELD Route 20, Brimfield, Mass.
+May 10, 1990, Page 00006 The New York Times Archives    HERE is a sampling of some of the better antiques and flea markets around the United States.
+Two or Three Times a Year    BRIMFIELD Route 20, Brimfield, Mass.
 01010; 413-245-3436.
 Second weekend of May and July, and the second weekend after Labor Day.
-{internallinenumbers*} RENNINGER’S OUTDOOR EXTRAVAGANZA Noble Street, Kutztown, Pa.; 717-385-0104.
+RENNINGER’S OUTDOOR EXTRAVAGANZA Noble Street, Kutztown, Pa.; 717-385-0104.
 Thursday, Friday and Saturday of the last weekend of April, June, September.
-{internallinenumbers*} FARMINGTON ANTIQUES WEEKEND Farmington Polo Grounds, Town Farm Road, Farmington, Conn.
+FARMINGTON ANTIQUES WEEKEND Farmington Polo Grounds, Town Farm Road, Farmington, Conn.
 06032; 508-839-9735.
@@
 1-2.
-Monthly              {internallinenumbers*} ANN ARBOR ANTIQUES MARKET, P.O.
+Monthly    ANN ARBOR ANTIQUES MARKET, P.O.
 Box 1512, Ann Arbor, Mich.
@@
 May through October, third Sunday.
-Continue reading the main story              {internallinenumbers*} KANE COUNTY FLEA MARKET, Kane County Fairgrounds, P.O.
+Continue reading the main story    KANE COUNTY FLEA MARKET, Kane County Fairgrounds, P.O.
 Box 549, St.
@@
 Year-round, first weekend.
-{internallinenumbers*} THE METROLINA EXPO, 7100 Statesville Road, Charlotte, N.C.
+THE METROLINA EXPO, 7100 Statesville Road, Charlotte, N.C.
 28213; 704-596-4643.
 Year-round, first weekend of every month.
-{internallinenumbers*} SPRINGFIELD ANTIQUE SHOW AND FLEA MARKET, Clark County Fairgrounds, Route 41, Springfield, Ohio, 45501; 513-325-0053.
+SPRINGFIELD ANTIQUE SHOW AND FLEA MARKET, Clark County Fairgrounds, Route 41, Springfield, Ohio, 45501; 513-325-0053.
 Year-round, third weekend.
-Weekly              {internallinenumbers*} BAKERSFIELD SWAP-O-RAMA, 4501 Wible Road, Bakersfield, Calif.
+Weekly    BAKERSFIELD SWAP-O-RAMA, 4501 Wible Road, Bakersfield, Calif.
 93313; 805-831-9342.
 Saturday and Sunday.
-{internallinenumbers*} LAMBERTVILLE ANTIQUE MARKET, Route 29, Lambertville, N.J.
+LAMBERTVILLE ANTIQUE MARKET, Route 29, Lambertville, N.J.
 08530.
@@
 Market on Saturday and Sunday.
-{internallinenumbers*} ATLANTA FLEA MARKET AND ANTIQUE CENTER, 5360 Peachtree Industrial Boulevard, Chamblee, Ga.
+ATLANTA FLEA MARKET AND ANTIQUE CENTER, 5360 Peachtree Industrial Boulevard, Chamblee, Ga.
 30341; 404-458-0456.
@@
 Continue reading the main story
-        [WET file]              A Guide To Markets - The New York Times        {internallinenumbers*} NYTimes.com no longer supports Internet Explorer 9 or earlier.
+     [WET file]    A Guide To Markets - The New York Times  NYTimes.com no longer supports Internet Explorer 9 or earlier.
 Please upgrade your browser.
 LEARN MORE »  Sections  Home  Search  Skip to content Skip to navigation View mobile version  The New York Times  Archives|A Guide To Markets  Search  Subscribe Now  Log In  0  Settings  Close search  Site Search Navigation  Search NYTimes.com  Clear this text input  Go  https://nyti.ms/29nVV3Q  Loading...
-See next articles        See previous articles        Site Navigation        Site Mobile Navigation        Advertisement        Archives | 1990        A Guide To Markets        MAY 10, 1990        Continue reading the main story Share This Page        Continue reading the main story        About the Archive        {internallinenumbers*} This is a digitized version of an article from The Times’s print archive, before the start of online publication in 1996.
+See next articles  See previous articles  Site Navigation  Site Mobile Navigation  Advertisement  Archives | 1990  A Guide To Markets  MAY 10, 1990  Continue reading the main story Share This Page  Continue reading the main story  About the Archive  This is a digitized version of an article from The Times’s print archive, before the start of online publication in 1996.
 To preserve these articles as they originally appeared, The Times does not alter, edit or update them.
-{internallinenumbers*} Occasionally the digitization process introduces transcription errors or other problems.
+Occasionally the digitization process introduces transcription errors or other problems.
 Please send reports of such problems to archive_feedback@nytimes.com.
-May 10, 1990, Page 00006 The New York Times Archives        {internallinenumbers*} HERE is a sampling of some of the better antiques and flea markets around the United States.
-Two or Three Times a Year        {internallinenumbers*} BRIMFIELD Route 20, Brimfield, Mass.
+May 10, 1990, Page 00006 The New York Times Archives  HERE is a sampling of some of the better antiques and flea markets around the United States.
+Two or Three Times a Year  BRIMFIELD Route 20, Brimfield, Mass.
 01010; 413-245-3436.
 Second weekend of May and July, and the second weekend after Labor Day.
-{internallinenumbers*} RENNINGER’S OUTDOOR EXTRAVAGANZA Noble Street, Kutztown, Pa.; 717-385-0104.
+RENNINGER’S OUTDOOR EXTRAVAGANZA Noble Street, Kutztown, Pa.; 717-385-0104.
 Thursday, Friday and Saturday of the last weekend of April, June, September.
-{internallinenumbers*} FARMINGTON ANTIQUES WEEKEND Farmington Polo Grounds, Town Farm Road, Farmington, Conn.
+FARMINGTON ANTIQUES WEEKEND Farmington Polo Grounds, Town Farm Road, Farmington, Conn.
 06032; 508-839-9735.
@@
 1-2.
-Advertisement        Continue reading the main story        Monthly        {internallinenumbers*} ANN ARBOR ANTIQUES MARKET, P.O.
+Advertisement  Continue reading the main story  Monthly  ANN ARBOR ANTIQUES MARKET, P.O.
 Box 1512, Ann Arbor, Mich.
@@
 May through October, third Sunday.
-Continue reading the main story        Advertisement        Continue reading the main story        {internallinenumbers*} KANE COUNTY FLEA MARKET, Kane County Fairgrounds, P.O.
+Continue reading the main story  Advertisement  Continue reading the main story  KANE COUNTY FLEA MARKET, Kane County Fairgrounds, P.O.
 Box 549, St.
@@
 Year-round, first weekend.
-{internallinenumbers*} THE METROLINA EXPO, 7100 Statesville Road, Charlotte, N.C.
+THE METROLINA EXPO, 7100 Statesville Road, Charlotte, N.C.
 28213; 704-596-4643.
 Year-round, first weekend of every month.
-{internallinenumbers*} SPRINGFIELD ANTIQUE SHOW AND FLEA MARKET, Clark County Fairgrounds, Route 41, Springfield, Ohio, 45501; 513-325-0053.
+SPRINGFIELD ANTIQUE SHOW AND FLEA MARKET, Clark County Fairgrounds, Route 41, Springfield, Ohio, 45501; 513-325-0053.
 Year-round, third weekend.
-Weekly        {internallinenumbers*} BAKERSFIELD SWAP-O-RAMA, 4501 Wible Road, Bakersfield, Calif.
+Weekly  BAKERSFIELD SWAP-O-RAMA, 4501 Wible Road, Bakersfield, Calif.
 93313; 805-831-9342.
 Saturday and Sunday.
-{internallinenumbers*} LAMBERTVILLE ANTIQUE MARKET, Route 29, Lambertville, N.J.
+LAMBERTVILLE ANTIQUE MARKET, Route 29, Lambertville, N.J.
 08530.
@@
 Market on Saturday and Sunday.
-{internallinenumbers*} ATLANTA FLEA MARKET AND ANTIQUE CENTER, 5360 Peachtree Industrial Boulevard, Chamblee, Ga.
+ATLANTA FLEA MARKET AND ANTIQUE CENTER, 5360 Peachtree Industrial Boulevard, Chamblee, Ga.
 30341; 404-458-0456.
 Friday, Saturday and Sunday.
-{internallinenumbers*} A version of this list appears in print on May 10, 1990, on Page C00006 of the National edition with the headline: A Guide To Markets.
+A version of this list appears in print on May 10, 1990, on Page C00006 of the National edition with the headline: A Guide To Markets.
 Order Reprints| Today’s Paper|Subscribe  Continue reading the main story  What’s Next  Loading...
@@
 Events Guide  Real Estate  T Magazine  Travel  Weddings & Celebrations  Listings & More  Reader Center  Classifieds  Tools & Services  N.Y.C.
-Events Guide        Multimedia        Photography        Video        NYT Store        Times Journeys        Subscribe        Manage My Account        NYTCo        Subscribe        Subscribe        Home Delivery        Digital Subscriptions        Crossword        Email Newsletters        Gift Subscriptions        Group Subscriptions        Education Rate        Mobile Applications        Replica Edition        Site Information Navigation        © 2019 The New York Times Company        Home        Search        {internallinenumbers*} Accessibility concerns?
+Events Guide  Multimedia  Photography  Video  NYT Store  Times Journeys  Subscribe  Manage My Account  NYTCo  Subscribe  Subscribe  Home Delivery  Digital Subscriptions  Crossword  Email Newsletters  Gift Subscriptions  Group Subscriptions  Education Rate  Mobile Applications  Replica Edition  Site Information Navigation  © 2019 The New York Times Company  Home  Search  Accessibility concerns?
 Email us at accessibility@nytimes.com.
@@
 With a bigger pause?
-{internallinenumbers*} This question has been asked before and already has an answer.
+This question has been asked before and already has an answer.
 If those answers do not fully address your question, please ask a new question.
-{internallinenumbers*} The dashes you described are known respectively as the en-dash and the em-dash.
-To describe the difference between their origins, Mental Floss writes:        {internallinenumbers*} An en dash (–) is bigger than a hyphen but shorter than an em dash (—).
+The dashes you described are known respectively as the en-dash and the em-dash.
+To describe the difference between their origins, Mental Floss writes:  An en dash (–) is bigger than a hyphen but shorter than an em dash (—).
 Th e names come from an obscure typographical measurement system, but the dashes have now taken on a life of their own in grammar.
@@
 So when do you use an en-dash?
-Again from Mental Floss:        {internallinenumbers*} To show numerical ranges, signifying “up to and including”—of dates, ages, pages, etc.
-(Example: “I read pages 7–22 last night.”)        {internallinenumbers*} The storied “compound adjective hyphen,” an event so rare in the English language that proofreaders shiver with excitement whenever they come across it.
+Again from Mental Floss:  To show numerical ranges, signifying “up to and including”—of dates, ages, pages, etc.
+(Example: “I read pages 7–22 last night.”)  The storied “compound adjective hyphen,” an event so rare in the English language that proofreaders shiver with excitement whenever they come across it.
 Basically “pro-American” gets a regular hyphen because “American” is only one word, whereas “pro–Falkland Islands” gets an en dash because “Falkland Islands” is two words.
 So, too phrases like “Civil War–era.”  What about an em-dash?
-From here:        {internallinenumbers*} Similar to an extended hyphen (-), an em dash is used to show a break in thought or a shift of tone.
-{internallinenumbers*} If you’d like to read more about the differences between a hyphen (-), en-dash (–), and em-dash (—), see the blog post here which summarizes the above.
+From here:  Similar to an extended hyphen (-), an em dash is used to show a break in thought or a shift of tone.
+If you’d like to read more about the differences between a hyphen (-), en-dash (–), and em-dash (—), see the blog post here which summarizes the above.
      [Resiliparse]    1    Possible Duplicate:  When should I use an em-dash, an en-dash, and a hyphen?
@@
 With a bigger pause?
-{internallinenumbers*} marked as duplicate by waiwai933, MrHen, user2683, Robusto, Thursagen Jul 13 ’11 at 0:32              {internallinenumbers*} This question has been asked before and already has an answer.
+marked as duplicate by waiwai933, MrHen, user2683, Robusto, Thursagen Jul 13 ’11 at 0:32    This question has been asked before and already has an answer.
 If those answers do not fully address your question, please ask a new question.
-0              {internallinenumbers*} The dashes you described are known respectively as the en-dash and the em-dash.
-To describe the difference between their origins, Mental Floss writes:              {internallinenumbers*} An en dash (–) is bigger than a hyphen but shorter than an em dash (—).
+0    The dashes you described are known respectively as the en-dash and the em-dash.
+To describe the difference between their origins, Mental Floss writes:    An en dash (–) is bigger than a hyphen but shorter than an em dash (—).
 Th e names come from an obscure typographical measurement system, but the dashes have now taken on a life of their own in grammar.
@@
 So when do you use an en-dash?
-Again from Mental Floss:              {internallinenumbers*} 1.
+Again from Mental Floss:      1.
 To show numerical ranges, signifying “up to and including”—of dates, ages, pages, etc.
-(Example: “I read pages 7–22 last night.”)              {internallinenumbers*} 2.
+(Example: “I read pages 7–22 last night.”)      2.
 The storied “compound adjective hyphen,” an event so rare in the English language that proofreaders shiver with excitement whenever they come across it.
@@
 So, too phrases like “Civil War–era.”    What about an em-dash?
-From here:              {internallinenumbers*} Similar to an extended hyphen (-), an em dash is used to show a break in thought or a shift of tone.
-{internallinenumbers*} If you’d like to read more about the differences between a hyphen (-), en-dash (–), and em-dash (—), see the blog post here which summarizes the above.
-{internallinenumbers*} Not the answer you’re looking for?
+From here:    Similar to an extended hyphen (-), an em dash is used to show a break in thought or a shift of tone.
+If you’d like to read more about the differences between a hyphen (-), en-dash (–), and em-dash (—), see the blog post here which summarizes the above.
+Not the answer you’re looking for?
 Browse other questions tagged or ask your own question.
-        [WET file]              {internallinenumbers*} syntax - What’s the difference between - and -- in a phrase?
-- English Language & Usage Stack Exchange        Stack Exchange Network        {internallinenumbers*} Stack Exchange network consists of 175 Q&A communities including Stack Overflow, the largest, most trusted online community for developers to learn, share their knowledge, and build their careers.
+     [WET file]    syntax - What’s the difference between - and -- in a phrase?
+- English Language & Usage Stack Exchange  Stack Exchange Network  Stack Exchange network consists of 175 Q&A communities including Stack Overflow, the largest, most trusted online community for developers to learn, share their knowledge, and build their careers.
 Visit Stack Exchange  Log In Sign Up  current community  English Language & Usage  help chat  English Language & Usage Meta  your communities  Sign up or log in to customize your list.
-more stack exchange communities        company blog        Tour Start here for a quick overview of the site        Help Center Detailed answers to any questions you might have        Meta Discuss the workings and policies of this site        About Us Learn more about Stack Overflow the company        Business Learn more about hiring developers or posting ads with us        {internallinenumbers*} By using our site, you acknowledge that you have read and understand our Cookie Policy, Privacy Policy, and our Terms of Service.
-{internallinenumbers*} English Language & Usage Stack Exchange is a question and answer site for linguists, etymologists, and serious English language enthusiasts.
+more stack exchange communities  company blog  Tour Start here for a quick overview of the site  Help Center Detailed answers to any questions you might have  Meta Discuss the workings and policies of this site  About Us Learn more about Stack Overflow the company  Business Learn more about hiring developers or posting ads with us  By using our site, you acknowledge that you have read and understand our Cookie Policy, Privacy Policy, and our Terms of Service.
+English Language & Usage Stack Exchange is a question and answer site for linguists, etymologists, and serious English language enthusiasts.
 Join them; it only takes a minute:  Sign up  Here’s how it works:  Anybody can ask a question  Anybody can answer  The best answers are voted up and rise to the top  Home  Questions  Tags  Users  Unanswered  What’s the difference between - and — in a phrase?
@@
 With a bigger pause?
-syntax dashes symbols        share|improve this question        edited Apr 13 ’17 at 12:38        Community        1        asked Jul 13 ’11 at 0:10        curiouscurious        123115        {internallinenumbers*} marked as duplicate by waiwai933, MrHen, user2683, Robusto, Thursagen Jul 13 ’11 at 0:32        {internallinenumbers*} This question has been asked before and already has an answer.
+syntax dashes symbols  share|improve this question  edited Apr 13 ’17 at 12:38  Community  1  asked Jul 13 ’11 at 0:10  curiouscurious  123115  marked as duplicate by waiwai933, MrHen, user2683, Robusto, Thursagen Jul 13 ’11 at 0:32  This question has been asked before and already has an answer.
 If those answers do not fully address your question, please ask a new question.
-add a comment |        1 Answer 1        active oldest votes        0        {internallinenumbers*} The dashes you described are known respectively as the en-dash and the em-dash.
-To describe the difference between their origins, Mental Floss writes:        {internallinenumbers*} An en dash (–) is bigger than a hyphen but shorter than an em dash (—).
+add a comment |  1 Answer 1  active oldest votes  0  The dashes you described are known respectively as the en-dash and the em-dash.
+To describe the difference between their origins, Mental Floss writes:  An en dash (–) is bigger than a hyphen but shorter than an em dash (—).
 Th e names come from an obscure typographical measurement system, but the dashes have now taken on a life of their own in grammar.
@@
 So when do you use an en-dash?
-Again from Mental Floss:        {internallinenumbers*} To show numerical ranges, signifying “up to and including”—of dates, ages, pages, etc.
-(Example: “I read pages 7–22 last night.”)        {internallinenumbers*} The storied “compound adjective hyphen,” an event so rare in the English language that proofreaders shiver with excitement whenever they come across it.
+Again from Mental Floss:  To show numerical ranges, signifying “up to and including”—of dates, ages, pages, etc.
+(Example: “I read pages 7–22 last night.”)  The storied “compound adjective hyphen,” an event so rare in the English language that proofreaders shiver with excitement whenever they come across it.
 Basically “pro-American” gets a regular hyphen because “American” is only one word, whereas “pro–Falkland Islands” gets an en dash because “Falkland Islands” is two words.
 So, too phrases like “Civil War–era.”  What about an em-dash?
-From here:        {internallinenumbers*} Similar to an extended hyphen (-), an em dash is used to show a break in thought or a shift of tone.
-{internallinenumbers*} If you’d like to read more about the differences between a hyphen (-), en-dash (–), and em-dash (—), see the blog post here which summarizes the above.
-share|improve this answer        edited Aug 4 ’15 at 17:00        zwol        2,51911424        answered Jul 13 ’11 at 0:16        simchonasimchona        30.9k5112139        add a comment |        {internallinenumbers*} Not the answer you’re looking for?
+From here:  Similar to an extended hyphen (-), an em dash is used to show a break in thought or a shift of tone.
+If you’d like to read more about the differences between a hyphen (-), en-dash (–), and em-dash (—), see the blog post here which summarizes the above.
+share|improve this answer  edited Aug 4 ’15 at 17:00  zwol  2,51911424  answered Jul 13 ’11 at 0:16  simchonasimchona  30.9k5112139  add a comment |  Not the answer you’re looking for?
 Browse other questions tagged syntax dashes symbols or ask your own question.
@@
 6  What is the difference between `-` and `--`  5  Does “cost-benefit ratio” use a hyphen or an en-dash?
-0        {internallinenumbers*} What kind of dash character should I use at the end of a famous saying to mark of the author?
+0  What kind of dash character should I use at the end of a famous saying to mark of the author?
 -1  dash non-restrictive element in the middle of a sentence  1  em dash followed by a comma  0  what’s the difference between a hyphen, a dash and a minus sign?
@@
 Generate an RGB colour grid  How to Make a Beautiful Stacked 3D Plot  How to find all the available tools in mac terminal?
-{internallinenumbers*} Has negative voting ever been officially implemented in elections, or seriously proposed, or even studied?
+Has negative voting ever been officially implemented in elections, or seriously proposed, or even studied?
 What is the meaning of the simile “quick as silk”?
 Circuit to "zoom in" on mV fluctuations of a DC signal?
-{internallinenumbers*} What are the out-of-universe reasons for the references to Toby Maguire-era Spider-Man in Into the Spider-Verse?
+What are the out-of-universe reasons for the references to Toby Maguire-era Spider-Man in Into the Spider-Verse?
 What does this Jacques Hadamard quote mean?
@@
 Why wasn’t DOSKEY integrated with COMMAND.COM?
-more hot questions        English Language & Usage        Tour        Help        Chat        Contact        Feedback        Mobile        Company        Stack Overflow        Stack Overflow Business        Developer Jobs        About        Press        Legal        Privacy Policy        Stack Exchange        Network        Technology        Life / Arts        Culture / Recreation        Science        Other        Stack Overflow        Server Fault        Super User        Web Applications        Ask Ubuntu        Webmasters        Game Development        TeX - LaTeX        Software Engineering        Unix & Linux        Ask Different (Apple)        WordPress Development        Geographic Information Systems        Electrical Engineering        Android Enthusiasts        Information Security        Database Administrators        Drupal Answers        SharePoint        User Experience        Mathematica        Salesforce        ExpressionEngine® Answers        Stack Overflow em Português        Blender        Network Engineering        Cryptography        Code Review        Magento        Software Recommendations        Signal Processing        Emacs        Raspberry Pi        Stack Overflow Ha pycckom        Programming Puzzles & Code Golf        Stack Overflow en español        Ethereum        Data Science        Arduino        Bitcoin        more (31)        Photography        Science Fiction & Fantasy        Graphic Design        Movies & TV        Music: Practice & Theory        Worldbuilding        Seasoned Advice (cooking)        Home Improvement        Personal Finance & Money        Academia        Law        more (15)        English Language & Usage        Skeptics        Mi Yodeya (Judaism)        Travel        Christianity        English Language Learners        Japanese Language        Arqade (gaming)        Bicycles        Role-playing Games        Anime & Manga        Puzzling        Motor Vehicle Maintenance & Repair        more (33)        MathOverflow        Mathematics        Cross Validated (stats)        Theoretical Computer Science        Physics        Chemistry        Biology        Computer Science        Philosophy        more (10)        Meta Stack Exchange        Stack Apps        API        Data        Blog        Facebook        Twitter        LinkedIn        {internallinenumbers*} site design / logo © 2019 Stack Exchange Inc; user contributions licensed under cc by-sa 3.0 with attribution required.
+more hot questions  English Language & Usage  Tour  Help  Chat  Contact  Feedback  Mobile  Company  Stack Overflow  Stack Overflow Business  Developer Jobs  About  Press  Legal  Privacy Policy  Stack Exchange  Network  Technology  Life / Arts  Culture / Recreation  Science  Other  Stack Overflow  Server Fault  Super User  Web Applications  Ask Ubuntu  Webmasters  Game Development  TeX - LaTeX  Software Engineering  Unix & Linux  Ask Different (Apple)  WordPress Development  Geographic Information Systems  Electrical Engineering  Android Enthusiasts  Information Security  Database Administrators  Drupal Answers  SharePoint  User Experience  Mathematica  Salesforce  ExpressionEngine® Answers  Stack Overflow em Português  Blender  Network Engineering  Cryptography  Code Review  Magento  Software Recommendations  Signal Processing  Emacs  Raspberry Pi  Stack Overflow Ha pycckom  Programming Puzzles & Code Golf  Stack Overflow en español  Ethereum  Data Science  Arduino  Bitcoin  more (31)  Photography  Science Fiction & Fantasy  Graphic Design  Movies & TV  Music: Practice & Theory  Worldbuilding  Seasoned Advice (cooking)  Home Improvement  Personal Finance & Money  Academia  Law  more (15)  English Language & Usage  Skeptics  Mi Yodeya (Judaism)  Travel  Christianity  English Language Learners  Japanese Language  Arqade (gaming)  Bicycles  Role-playing Games  Anime & Manga  Puzzling  Motor Vehicle Maintenance & Repair  more (33)  MathOverflow  Mathematics  Cross Validated (stats)  Theoretical Computer Science  Physics  Chemistry  Biology  Computer Science  Philosophy  more (10)  Meta Stack Exchange  Stack Apps  API  Data  Blog  Facebook  Twitter  LinkedIn  site design / logo © 2019 Stack Exchange Inc; user contributions licensed under cc by-sa 3.0 with attribution required.
 rev 2019.4.18.33353  English Language & Usage Stack Exchange works best with JavaScript enabled
@@
 These hyperparameters were chose programmatically to mimic the Jaccard similarity plots as closely as possible, in an $\ell_{2}$ sense, with a fixed hash budget.
-See Figure˜6 for more details.
+See Figure 6 for more details.
 Figure 6: Probability of two documents with Jaccard similarity (x-axis) being marked as duplicates (y-axis) with varying (number of buckets, bucket size) parameters.
@@
 As an example, suppose we are given a paragraph of $N$ ngrams, where $S$ of them are already contained in the Bloom filter and we set threshold to $T$.
-Because Bloom filters do not allow false negatives, every one of the $S$ ngrams are marked (correctly) as contained, and NSN-S of them could potentially be marked as a false positive.
-Indeed, of the NSN-S of these ngrams, at least TNSTN-S of them would need to be marked as a false positive, each of which occurs independently with probability $\epsilon$.
-This is equivalent to NSN-S Bernoulli random variables with parameter $\epsilon$, and can be bounded by a crude Hoeffding bound.
+Because Bloom filters do not allow false negatives, every one of the $S$ ngrams are marked (correctly) as contained, and $N-S$ of them could potentially be marked as a false positive.
+Indeed, of the $N-S$ of these ngrams, at least $TN-S$ of them would need to be marked as a false positive, each of which occurs independently with probability $\epsilon$.
+This is equivalent to $N-S$ Bernoulli random variables with parameter $\epsilon$, and can be bounded by a crude Hoeffding bound.
 In this particular case, the probability that a document or paragraph is falsely marked as a duplicate is bounded by:
@@
   • Which deduplication pipeline is worth scaling up to larger pool sizes?
-Results are contained in Table˜18.
+Results are contained in Table 18.
 The main conclusions we can arrive at from this table are as follows: i) Suffix Array deduplication seems to help more than MinHash deduplication, thereby giving some signal to the source of the gains procured by a MinHash+SuffixArray pipeline; ii) BFF provides comparable performance to a full Exact+MinHash+SuffixArray pipeline, giving strong evidence that the multiresolution BFF could be an easily scalable alternative to the relatively more expensive MinHash+SuffixArray pipeline of prior works.
@@
   • Which BFF hyperparameters yield the highest Core and MMLU performance at this scale?
-Results are contained in Table˜19.
+Results are contained in Table 19.
 The first point to note is that BFF with a min_ngram_size at 13 and 20 yields Core scores and MMLU scores that are comparable to the scores attained by a MinHash+SuffixArray deduplicated pool at the same scale.
@@
 Again we start with a pool obtained from one tenth of Common Crawl and generate several deduplicated pools.
-The questions of interest are the same as above and we summarize the results in Table˜20.
+The questions of interest are the same as above and we summarize the results in Table 20.
 The key takeaways from this round of ablations is that at the 7B-2x scale, BFF with a min_ngram_size of 13 and 10 shards attains nearly identical performance to a MinHash+SuffixArray pipeline, whereas BFF with a min_ngram_size of 20 and 32 shards starts to lag behind, and that a min_ngram_size of 5 yields competitive Core scores, but falters in MMLU evaluations.
@@
 Table 20: Deduplication Ablations (7B-2x scale).
-From the same pools as in Table˜19, we trained and evaluated models at the 7B-2x scale.
+From the same pools as in Table 19, we trained and evaluated models at the 7B-2x scale.
 Notice that a min_ngram_size of 5 yields competitive Core results but drastically reduces MMLU scores.
@@
 0.001 20.47%
-Here we start with the 75B token data pool as in Section˜K.2.1 and focus on a paragraph-only level BFF.
+Here we start with the 75B token data pool as in Section K.2.1 and focus on a paragraph-only level BFF.
 In other words, we run BFF as described above, except omit the full-document removal step.
@@
 We specifically look at the effect of changing the false positive rate and compute the removal rate (in bytes) of the output.
-From Table˜21, we can see that a false positive rate of 0.1 suffices for a reasonably small pool such as this one.
+From Table 21, we can see that a false positive rate of 0.1 suffices for a reasonably small pool such as this one.
 For larger pools, to be safe, we always set the false positive rate to 0.01.
 Min ngram size.
-From Table˜19 and Table˜20, we saw that altering the ngram_size hyperparameters can affect both token yield and evaluation metrics.
+From Table 19 and Table 20, we saw that altering the ngram_size hyperparameters can affect both token yield and evaluation metrics.
 In particular, we seek to examine how surviving documents are altered by deduplication.
 As a proxy for this, we focus on the document lengths and removal rates.
-Results for this paragraph and the two following paragraphs are collated in Table˜22.
+Results for this paragraph and the two following paragraphs are collated in Table 22.
 One key observation is that as the min_ngram_size parameter is reduced, the mean and median document lengths become shorter.
@@
 For larger pools, we can always shard them as heavily as desired, so we treat sharding as a hyperparameter that controls removal rate and document statistics.
-Results are collated in Table˜23 and Figure˜7.
+Results are collated in Table 23 and Figure 7.
 The key takeaways here are that removal rates increase monotonically with dataset size as expected, but do so in a concave fashion.
@@
 In building DCLM-baseline, at the point of deduplication, the dataset is approximately 70TB in size.
-Since Table˜20 shows that BFF had the best performance at a 10-way shard with a roughly 7TB input size, we adhere to a 100-way sharding for DCLM-baseline, where each shard is roughly 700GB in size.
+Since Table 20 shows that BFF had the best performance at a 10-way shard with a roughly 7TB input size, we adhere to a 100-way sharding for DCLM-baseline, where each shard is roughly 700GB in size.
 Figure 7: Deduplication shard size.
@@
 Finally, to get a sense for the duplicates remaining in a dataset after a full processing pipeline has been applied, we run a global (i.e., one shard) MinHash on several open datasets.
-These results are collated in table table˜24.
+These results are collated in table table 24.
 We evaluate our DCLM-baseline, the official RefinedWeb dataset from HuggingFace, our emulation of the RefinedWeb pipeline, and Dolma V1.
-MinHash is performed using 14 buckets and a bucket size of 9, corresponding to the green curve in fig.˜6.
+MinHash is performed using 14 buckets and a bucket size of 9, corresponding to the green curve in fig.
+6.
 We note several observations here.
@@
 DCLM-baseline 3.2B (Fuzzy) Bloom Filter 100 85%
-RefinedWeb (official) 968B MinHash+SA 1* 0%
+RefinedWeb (official) 968M MinHash+SA 1* 0%
 RefinedWeb (ours) 2.0B MinHash+SA 16 45%
@@
 Appendix L Mixing sources
-In Section˜4.5, we showed that mixing our dataset with the usual sources did not improve its general performance, and hypothesized that this is due to the more stringent filtering performed for our Common Crawl portion.
+In Section 4.5, we showed that mixing our dataset with the usual sources did not improve its general performance, and hypothesized that this is due to the more stringent filtering performed for our Common Crawl portion.
 One could argue that improved filtering in the other sources could lead to similar improvements in performance.
@@
 We then add the resulting data to our pretraining dataset, and train models at the 1B-1x scale.
-The results of this can be seen in Table˜25.
+The results of this can be seen in Table 25.
 We see that despite the more uniform handling of mixing across various sources, the additional sources still decrease performance.
@@
 We further extract the subset of 281 samples where all three annotators are in agreement, naming the full data Majority and the subset Agreement.
-We then evaluate various quality filters from Section˜4.4 on this data to search for correlation between dataset quality (as measured by Core accuracy) and filter agreement with human labels.
-Figure˜8 (a) depicts the Core scores of models trained on datasets filtered with the respective quality filter against ROC-AUC [^5] ROC-AUC measures a classifier’s ability to distinguish between classes by summarizing the trade-off between true positive and false positive rates across all thresholds, making it a robust and common metric for model performance.
+We then evaluate various quality filters from Section 4.4 on this data to search for correlation between dataset quality (as measured by Core accuracy) and filter agreement with human labels.
+Figure 8 (a) depicts the Core scores of models trained on datasets filtered with the respective quality filter against ROC-AUC [^5] ROC-AUC measures a classifier’s ability to distinguish between classes by summarizing the trade-off between true positive and false positive rates across all thresholds, making it a robust and common metric for model performance.
 See https://en.wikipedia.org/wiki/Receiver_operating_characteristic for further details.
 of our quality-filters on the Majority data.
-Notably, both the best and worst fastText-based filters score about the same on the Majority data ($\sim 73\%$ ROC-AUC), while the AskLLM filter that is highly correlated with human annotations ($\penalty 10000\ \sim 82\%$ ROC-AUC) performs much worse as a quality filter ($\sim 28.5\%$ Core compared to $31\%$ in several fastText classifiers).
+Notably, both the best and worst fastText-based filters score about the same on the Majority data ($\sim 73\%$ ROC-AUC), while the AskLLM filter that is highly correlated with human annotations ($\leavevmode\nobreak\ \sim 82\%$ ROC-AUC) performs much worse as a quality filter ($\sim 28.5\%$ Core compared to $>31\%$ in several fastText classifiers).
 We continue this study by inspecting correlations to specific downstream tasks and comparing them to the ROC-AUC on the Agreement data, where all three annotators agreed on the label.
-Figure˜8 depicts the scores on a few tasks against the ROC-AUC on the annotated data of the representative set of quality filters.
+Figure 8 depicts the scores on a few tasks against the ROC-AUC on the annotated data of the representative set of quality filters.
 While some positive correlation may be observed for StrategyQA [62], the opposite is true for other QA datasets such as SQuAD [133], and in both cases the $R^{2}<0.3$.
-In most other downstream tasks, the results are similar to Figure˜8 (a), where no correlation can be observed.
+In most other downstream tasks, the results are similar to Figure 8 (a), where no correlation can be observed.
 This suggests that human intuition may not reliably identify the most useful documents for language model training purposes.
@@
 We sampled 499 random documents from our pool, after going through the rule-based quality filters and deduplication.
-Figure˜9 shows a histogram of the length of the documents in words.
+Figure 9 shows a histogram of the length of the documents in words.
 We asked 16 English speak AI graduate and professors to annotate each example as a good candidate to be included in an LM pretraining corpus (see instructions and some examples given to annotators below).
@@
 Instructions given to annotators
-        {internallinenumbers*} Your task is to assess the quality of the documents (0-bad, 1-good), sampled randomly from RefinedWebV2, in terms of their usefulness for LLM training.
+     Your task is to assess the quality of the documents (0-bad, 1-good), sampled randomly from RefinedWebV2, in terms of their usefulness for LLM training.
 How do you judge the usefulness of a document?
-While this is a subjective task, there are a few things to keep in mind:        {internallinenumbers*} 1.
+While this is a subjective task, there are a few things to keep in mind:      1.
 We encourage you to review the details of the evaluation tasks that the LLMs are expected to excel in [link to spreadsheet with the tasks was provided].
 Broadly, these tasks include reading comprehension, language understanding, world knowledge, and commonsense reasoning.
-{internallinenumbers*} 2.
+2.
 Check whether a particular document will be useful for answering any open-ended imperative tasks or questions.
 For example: writing a letter to a friend, writing a poem, replying to an email, cooking a dish, etc.
-{internallinenumbers*} 3.
+3.
 Check whether the data contains harmful content that you would not want an LLM to generate.
@@
 Doc2 (Bad)
-        Host: Zander        Program Category: Music        Frequency: Weekly        Length: 2 Hours        Terms: Barter        Delivery Method: Internet        {internallinenumbers*} |“Zander’s knowledge of music and his straight-forward approach has struck a huge interest among our listeners.
-The Rockin’ 80’s is EXACTLY what we’ve been looking for!"" - Terry West, WQLA        {internallinenumbers*} The Rockin’ 80’s is the only 80’s show with a mix of the best rock from the decade of excess plus “oh wow” tracks that add spice to the weekly line up.
+     Host: Zander  Program Category: Music  Frequency: Weekly  Length: 2 Hours  Terms: Barter  Delivery Method: Internet  |“Zander’s knowledge of music and his straight-forward approach has struck a huge interest among our listeners.
+The Rockin’ 80’s is EXACTLY what we’ve been looking for!"" - Terry West, WQLA  The Rockin’ 80’s is the only 80’s show with a mix of the best rock from the decade of excess plus “oh wow” tracks that add spice to the weekly line up.
 The two hour version of the show features rarities from the 80’s “Lost and Found”, an 80’s “Two-Fer,” spotlighting two contrasting songs from one band played back to back and much more.
 Featured core musical artists include Guns N’ Roses, Motley Crue, Van Halen, Rush and AC/DC.
-{internallinenumbers*} |3733 Park East Drive • Room 222 • Cleveland, Ohio 44122 P: 216-831-3761 • F: 216-514-4699 • Email us        |©2014 Envision Networks.
+|3733 Park East Drive • Room 222 • Cleveland, Ohio 44122 P: 216-831-3761 • F: 216-514-4699 • Email us  |©2014 Envision Networks.
 All rights reserved.
@@
 Doc3 (Good)
-        |Chinese Five-Spice Noodles with Shitake Mushrooms        |Recipes - Chinese Five Spice        Chinese Five-Spice Noodles with Shitake Mushrooms        {internallinenumbers*} Cook the noodles until tender according to the package directions.
+     |Chinese Five-Spice Noodles with Shitake Mushrooms  |Recipes - Chinese Five Spice  Chinese Five-Spice Noodles with Shitake Mushrooms  Cook the noodles until tender according to the package directions.
 Drain and rinse under cold running water, and drain well again.
-{internallinenumbers*} Bring 2 1/2 cups water to a boil in a small saucepan.
+Bring 2 1/2 cups water to a boil in a small saucepan.
 Add the dried mushrooms, cover, and simmer for 3 minutes.
@@
 Roughly chop the mushrooms, then set aside with the liquid.
-{internallinenumbers*} Heat the butter and oil in a wok or large saucepan over medium heat.
+Heat the butter and oil in a wok or large saucepan over medium heat.
 Add the Chinese five-spice powder and shallots and saute, stirring frequently, until tender, about 2 minutes.
@@
 Add the remaining mushroom-soaking water.
-{internallinenumbers*} Add the noodles to the wok and stir until heated through and coated with the sauce, about 1 minute.
+Add the noodles to the wok and stir until heated through and coated with the sauce, about 1 minute.
 Garnish with the green onions and sesame seeds, if using, and serve at once.
@@
 Doc4 (Good)
-        {internallinenumbers*} "Is it necessary to purchase a travel book or is it realistic that we can get similar information from other resources?
+     "Is it necessary to purchase a travel book or is it realistic that we can get similar information from other resources?
 Usually, most individuals have a major question on buying a travel book.
 So here are the pros and cons of purchasing one such book.
-Advantages of a Travel Book        {internallinenumbers*} A travel book, which may be a paperback or e-book, comes in handy while traveling.
+Advantages of a Travel Book  A travel book, which may be a paperback or e-book, comes in handy while traveling.
 Glancing through a travel book enables you to understand the custom and culture of a particular place in the world.
 So you can adapt yourself to that particular environment and stay there comfortably for longer periods.
-{internallinenumbers*} - They Come In Handy — The travel guide comes in various forms such as, e-books, paperbacks and the file formats.
+- They Come In Handy — The travel guide comes in various forms such as, e-books, paperbacks and the file formats.
 You can have easy access to these books, which would assist you with all details compatible to the region you are traveling to.
-{internallinenumbers*} - They Provide Enormous Information — Electronic or traditional travel guides provide you with answers to all types of questions such as how to learn some sayings that can be used in the place where you are traveling to?
+- They Provide Enormous Information — Electronic or traditional travel guides provide you with answers to all types of questions such as how to learn some sayings that can be used in the place where you are traveling to?
 How to get data on where to reside, what to see and where to eat?
 How to get a clear knowledge about the history of a specific region or the atmosphere that it has?
-{internallinenumbers*} - They Suit To Your Requirements — To access full information about a specific country or a region, both types of general and specific travel books are made available.
+- They Suit To Your Requirements — To access full information about a specific country or a region, both types of general and specific travel books are made available.
 The e-book may easily fit into your e-book reader whereas the paperback can fit into your backpack.
-Disadvantages of Travel Book        {internallinenumbers*} - The Price — The e-book and paperback travel guides are very expensive compared to the information obtained from travel websites or from those who have moved or traveled to that region.
-{internallinenumbers*} - Qualitative Images In Travel Books — Most travel books are in black and white.
+Disadvantages of Travel Book  - The Price — The e-book and paperback travel guides are very expensive compared to the information obtained from travel websites or from those who have moved or traveled to that region.
+- Qualitative Images In Travel Books — Most travel books are in black and white.
 Only a few e-books consist of colored photos.
 Hence make a thorough revision before purchasing a travel guide or an e-book.
-{internallinenumbers*} - Travel Books Make The Trip Less Natural — Traveling can be made more spontaneous by acquiring suggestions from locals than from travel books."
+- Travel Books Make The Trip Less Natural — Traveling can be made more spontaneous by acquiring suggestions from locals than from travel books."
 Appendix N Decontamination
-In Section˜4.6, we examined the results of the contamination analysis with respect to MMLU.
+In Section 4.6, we examined the results of the contamination analysis with respect to MMLU.
 Here, we present some analysis with respect to the other validation sets.
-Instead of the MMLU-specific decontamination that we performed in Section˜4.6, here we follow a more general approach based on token overlaps.
+Instead of the MMLU-specific decontamination that we performed in Section 4.6, here we follow a more general approach based on token overlaps.
 Overall, a generally applicable decontamination rule is difficult to specify, given the potentially subjective nature of what constitutes as contamination in text data as well as the diversity in formats across tasks.
@@
 [162] (less than 80 % of the tokens in the sample are marked as contaminated), and between the full evaluation set and samples marked as “clean” using the same criterion (less than 20 % of the tokens are marked).
-Results can be seen in Figure˜10, where we see that the difference in performance over the full dataset and the “not dirty” samples is minimal.
+Results can be seen in Figure 10, where we see that the difference in performance over the full dataset and the “not dirty” samples is minimal.
 In fact, for BoolQ and SQuAD, which are marked as highly contaminated, our model performs slightly better on the “not dirty” subset.
-Moreover, in Figure˜11 we see that the difference in performance between the full evaluation set and the “clean” subset is similarly small for most datasets.
+Moreover, in Figure 11 we see that the difference in performance between the full evaluation set and the “clean” subset is similarly small for most datasets.
 We note here that it’s difficult to identify a correct threshold for what counts as a contaminated sample (as 20 % token overlap might lead to many false positives, but at the same time 80 % might be too high to detect all contaminated samples).
@@
 Most evaluation tasks (including MMLU) have similar performance in the full eval and in the “clean” subset.
-Instruction tuning has emerged as a critical step to allow users to interact with pretrained language models [164, 162, 108, 111, 179].
-To investigate whether models trained on DCLM-baseline can have strong instruction-following capabilities, we instruction-tune the DCLM-baseline (7B) with the OpenHermes-2.5 (OH-2.5) dataset [153].
-Specifically, we train our model to predict the response given the instruction from the instruction-tuning dataset.
-We train DCLM-baseline using the Adam [83] optimizer for $10$ epochs with a $10\%$ warmup ratio and a cosine learning rate schedule.
-We perform a hyperparameter search over the three learning rates $\{1\text{e-}7,5\text{e-}6,2\text{e-}5\}$ and report the best-performing numbers on the evaluation metrics i.e., AlpacaEval 2.0 length-controlled win-rate [48].
-Post-training, we generate the responses for the instructions from AlpacaEval using a sampling temperature of $0.3$ and maximum response length of $500$.
-We benchmark our model with relevant baselines (e.g., LLaMA-2-Chat [158], Zephyr-7B [159]) taken directly from the AlpacaEval leaderboard as well as Mistral-7B [75] finetuned in the same manner as DCLM-baseline.
+[…]
 Table 26: Instruction tuning results on AlpacaEval2.0.
@@
 Alpaca-7B 5.9
-We present the results in Table˜26.
+We present the results in Table 26.
 We find that DCLM-baseline finetuned with OH-2.5 outperforms various instruct models such as Zephyr-Beta-7B and Gemma-Instruct-7B.
@@
 Subsequently, we perform instruction-tuning and response generation of DCLM-baseline on this dataset with the training recipe mentioned above.
-We present the results in Table˜26.
+We present the results in Table 26.
 We find that DCLM-IT outperforms DCLM-baseline w/ OH-2.5 by $2.8$ percentage points.
@@
 Appendix P Scaling up
-The final run 2.5T shown in Figure˜1 and Table˜9 uses a two stage training procedure as followed in Groeneveld et al.
+The final run 2.5T shown in Figure 1 and Table 9 uses a two stage training procedure as followed in Groeneveld et al.
 [68], Hu et al.
@@
 [154].
-For stage 1 we use the hyper-parameters from Table˜27.
+For stage 1 we use the hyper-parameters from Table 27.
 Table 27: Hyper-parameters for large scale run. Note the LR schedule uses a training length of 4.4T, but we do not train for the full length as we stop early and cooldown.
@@
 For the cooldown distribution we use a mix of 70% DCLM-baseline with a tighter fastText threshold (top 7% rather than top 10%) and 30% ProofPile.
-We keep all the hyperparameters the same as Table˜27, so we cooldown to the same final learning rate, just over a smaller number of tokens.
+We keep all the hyperparameters the same as Table 27, so we cooldown to the same final learning rate, just over a smaller number of tokens.
 Before the cool-down MMLU performance was approximately 52%, and the LR was approximately $1\times 10^{-3}$.
@@
 Thus we the total number of tokens seen by this model is 2.5T.
-We present results of each individual cooldown and the model soup in Table˜28.
-The model in Figure˜1 and Table˜9 uses the final model soup after long-context training for 100B tokens as described in Section˜P.2.
+We present results of each individual cooldown and the model soup in Table 28.
+The model in Figure 1 and Table 9 uses the final model soup after long-context training for 100B tokens as described in Section P.2.
 Instruction tuning the scaled up model
-In Appendix˜O we show how instruction tuning the above “model soup" for 80B additional tokens leads to strong performance on instruction tuning benchmarks and out-performs instruction tuned variants of similar 7B models such as Gemma-7B.
-In addition to the IT benchmarks covered in Appendix˜O, In Table˜29 we show that a small amount of instruction tuning provides large improvements in “Extended" evals at the cost of a small degradation in “Core" and “MMLU" evals.
+In Appendix O we show how instruction tuning the above “model soup" for 80B additional tokens leads to strong performance on instruction tuning benchmarks and out-performs instruction tuned variants of similar 7B models such as Gemma-7B.
+In addition to the IT benchmarks covered in Appendix O, In Table 29 we show that a small amount of instruction tuning provides large improvements in “Extended" evals at the cost of a small degradation in “Core" and “MMLU" evals.
 Notably we note that our GSM8k performance goes 2.5% to 52.5% which is comparable to other similar language models that mixed IT data into pretraining such as Gemma-7B.
@@
 We use $k=1,10,20,30$, and for each case, we run the evaluation multiple times by changing the position of the golden document in the context and report the average.
-Results are reported in table˜30.
+Results are reported in table 30.
 We demonstrate that long context adaptation results in a checkpoint (DCLM-8k) that matches the original model on regular evaluations and significantly improves multi-document QA showing its long-context capabilities.
@@
 Appendix Q Account of compute costs
-We note the compute cost of training runs for each competition scale in Table˜1.
+We note the compute cost of training runs for each competition scale in Table 1.
 In total, we estimate that our runs for DCLM sum up to approximately 1M H100 hours.
@@
 Evaluation data
-Appendix˜G discusses all downstream tasks we use for our evaluation. Below we mention them again, and specify their licenses.
+Appendix G discusses all downstream tasks we use for our evaluation. Below we mention them again, and specify their licenses.
 Raw sources
@@
 We urge the user’s discretion in verifying their use abide by these terms-of-use.
-In addition to the above, as described in Sections˜5, 4.5 and 4.4 and Appendices˜L and O we make use of the following datasets:
+In addition to the above, as described in Sections 5, 4.5 and 4.4 and Appendices L and O we make use of the following datasets:
   1. 1.
@@
   2. 2.
-    StarCoder [88] and StarCoder2 [98] are used for some of our ablations (Section˜5.
+    StarCoder [90] and StarCoder2 [101] are used for some of our ablations (Section 4).
 While constructed from permissive data by extracting datasets that mention permissive licenses (e.g.
@@
   3. 3.
-    ProofPile2 [13] is used to scale up the dataset to the trillion tokens scale (Section˜5.
+    ProofPile2 [14] is used to scale up the dataset to the trillion tokens scale (Section 5).
 The authors do not alter the licenses of underlying datasets and ask users to follow guidelines and licenses as described in these datasets.
   4. 4.
-    GSM8k [39] was used in some of the ablations in Section˜5 and follow the MIT license.
+    GSM8k [41] was used in some of the ablations in Section 4 and follows the MIT license.
   5. 5.
-    RedPajama [156] is used for ablations in Section˜4.5 and Appendix˜L.
+    RedPajama [160] is used for ablations in Section 4.5 and Appendix L.
 Note that we do not release models or datasets that include this data.
@@
   1. 1.
-    transformers uses the Apache 2.0 License. [^9] https://github.com/huggingface/transformers/blob/main/LICENSE
+    transformers uses the Apache 2.0 License.999 https://github.com/huggingface/transformers/blob/main/LICENSE
   2. 2.
@@
 This slows down progress and forces a lot of duplicate work between research teams.
-Prior work mainly focuses on data curation in the context of supervised datasets and smaller scales (see Section˜2 and Appendix˜B).
+Prior work mainly focuses on data curation in the context of supervised datasets and smaller scales (see Section 2 and Appendix B).
 In our initial release of DCLM, we focus on 53 downstream language understanding tasks that also include reasoning abilities, math, code, and more.
-For details see Section˜3.5 and Appendix˜G.
+For details see Section 3.5 and Appendix G.
   2. Q2
@@
 Each sample is also tagged with metadata about its crawl time and additional information such as the detected language, for processed instances such as those in DCLM-baseline.
-Additional information can be found in Appendix˜E.
+Additional information can be found in Appendix E.
   2. Q6
@@
 However, we note that Common Crawl does not cover the entire web data, due to reach and compute limitations for instance.
-For our DCLM-baseline, we use various filtering and deduplication strategies as described in Section˜4 in the explicit attempt to improve its quality for preatining, thus removing low-quality instances, and in doing so, becoming non-representative of the full set of instances.
-For a complete treatment and visualization of our data processing funnel, see Sections˜4, 4.3 and 4.2 and Appendix˜E.
+For our DCLM-baseline, we use various filtering and deduplication strategies as described in Section 4 in the explicit attempt to improve its quality for preatining, thus removing low-quality instances, and in doing so, becoming non-representative of the full set of instances.
+For a complete treatment and visualization of our data processing funnel, see Sections 4, 4.3 and 4.2 and Appendix E.
   4. Q8
     What data does each instance consist of? “Raw” data (e.g., unprocessed text or images) or features? In either case, please provide a description.
-    • Each sample contains an a web-page url for and the extracted HTML content associated with.
-Additionally, each sample contains metadata fields shown in Table˜10 (e.g., WARC-Type, WARC-date, Content-Type etc.).
+    • Each sample contains a web-page url for and the extracted HTML content associated with.
+Additionally, each sample contains metadata fields shown in Table 10 (e.g., WARC-Type, WARC-date, Content-Type etc.).
   5. Q9
@@
     Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
-    • No. The evaluation procedure is made of tasks as described in Section˜3.5. We also attempt to prevent test set contamination in as described in Section˜4.6 and Appendix˜N.
+    • No. The evaluation procedure is made of tasks as described in Section 3.5. We also attempt to prevent test set contamination in as described in Section 4.6 and Appendix N.
   9. Q13
@@
     • Each sample is associated with a URL that links other external resources on the internet with no guarantee that the resources will exist in perpetuity or that that the resources will not change.
-However, the dataset itself contains already extracted HTML content and is thus self-contained for the purposes of this benchmark as described in Appendix˜C.
+However, the dataset itself contains already extracted HTML content and is thus self-contained for the purposes of this benchmark as described in Appendix C.
   11. Q15
@@
 Model-based features were run on GPU clusters.
-For software links see Q37 and Appendix˜R or refer to https://datacomp.ai/dclm.
+For software links see Q37 and Appendix R or refer to https://datacomp.ai/dclm.
   3. Q24
@@
 We evaluate these models on our testbed of 53 zero- and few-shot downstream tasks.
-See Sections˜4 and 3.5.
+See Sections 4 and 3.5.
   2. Q40
@@
 If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
-    • We distribute our datasets in full, including extracted page content and associated metadata under a standard CC-BY-4.0 licence (see Appendix˜E).
+    • We distribute our datasets in full, including extracted page content and associated metadata under a standard CC-BY-4.0 licence (see Appendix E).
 The code associated with DCLM is released under the MIT license.
@@
 However, we will address user takedown requests (see Q56).
-DCLM-Pool is inherently noisy, and its release aims to encourage researchers to study dataset cleaning in the context of image-text samples.
+DCLM-Pool is inherently noisy, and its release aims to encourage researchers to study dataset cleaning in the context of raw, web-crawled text samples.
   5. Q56